A Complete Bibliography of Publications in International Studies in the Philosophy of Science

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 April 2025 Version 1.13

Title word cross-reference

$12.99 [Orr24]. $45.00 [Bus19]. _{2} [Nee02]. ΔEΔt ≥ h [Ant97]. E = mc^{2} [Flo05, Flo06, Kra06].

//doi.org/10.4324/9781003441809 [Hue25].

0 [Ano02a]. 0-521-65027-5 [Ano02a]. 0-521-65976-0 [Ano02a]. 0-7923-6305-5 [Rad02].

1 [Rad02]. 1-4020-0247-5 [Rad02]. 18th [Rue02]. 18th-century [Rue02]. 19 [Kra06]. 1925 [Bor13].

20 [Flo06]. 2023 [Par24]. 20th [Ara20].

4th [Uga15].

5 [Rad02].

6 [Bus19].

8 [Ala24].

9 [Orr24]. 978 [Ala24, Bus19, Orr24]. 978-0-19-871853-6 [Duw16]. 978-0-19-921323-8 [Pla10]. 978-0-19-965498-7 [Bor13]. 978-0-262-03841-6 [Bus19]. 978-0-7456-5062-3 [Dud17]. 978-0-7456-5063-0 [Dud17]. 978-0-7456-5374-7 [Zha15]. 978-0-7456-5375-4 [Zha15]. 978-1-009-20188-9 [Orr24]. 978-1-032-58279-5 [Hue25].

978-1-107-00397-2 [Pau12]. 978-3-031-54226-8 [Ala24]. 9781003441809 [Hue25].

Abandon [Sha07]. Abduction [Paa06, Bla96, Nee99]. Absolute [Bel07]. Abstinence [Rup05]. Abstract [Kuo09, RA16, Ye10, Bro89]. Academia [Kre17]. Academic [Jac11, Kre17, Lan11, Rad02]. Accelerating [Kre17]. acceptance [Jon00]. accidental [Kos00]. accorded [Feh90]. Account [Asa13, Bae21, Bus19, Dim16, Gij09, Kow14, Max09, Web09, Dow92, Sed96, Sza03]. Accounts [BTB16, Gin12, Gur22]. Acknowledgement [Ano21, Ano22, Ano23]. Action [Bul09, Ori24, Par15, Alb93, Ber90, Feh88, Fuc93, Til99]. Activism [CS24]. Activities [Mac04]. acts [Til99]. actual [Mil91a]. Adaptationist [Vil23]. Addendum [Ano88a]. Adding [CT19, Til98]. Addition [Isl17]. Address [dMM23]. Adjusting [Rec17]. advancement [Elg02]. Advances [Col20, Gao15]. Aesthetic [RB02, Rue02, Tod08, Eng01, Eng05]. aesthetics [Dav98, Mam87, McA02]. affordances [Har90]. African [OM16]. After [Bor12, Jae17, Lew16]. Against [Fuc93, Gin05, Kha05, Liu16, Lyo13, PZ20, Rei01, Row10, Sol14, Tym21, DD02, Feh98a, Pay95, Sua03]. Age [Abr16]. Agency [Buz14, Pop16, Til99]. Agenda [Sch11]. agent [Jac93]. Agnes [Fla03]. Agnostic [Psi00]. Agnotological [dMM23]. Agnotology [Gir23]. AIC [HGS24]. AIDS [Bab15]. Aim [Gao17]. Aim-oriented [Gao17]. aimless [D’A90]. aims [BV24]. Akaike [HGS24]. Alan [Lon97]. alchemy [New01]. algebra [Lar94]. Algebraic [EV14]. Almagest} [Rec17]. also [Gir94]. Ampliative [Bla96]. Analogies [MR88, Har88]. Analogy [VG15, Feh88, Wyl88]. Analyses [Bae10].

Analysing [Bog04]. Analysis [Car10b, CMF16, Gon12, Rus10, Feh87, PZ14]. analysis-synthesis [Feh87]. Analytic [Bre15, Pla10, vdS14]. Animal [LS94]. Animals [McF92, McF91]. anomaly [Fri03, Lew03]. Answer [Ped18, How11]. Ante [SM94]. anthropological [Kut87]. Anthropomorphism [Buz14]. Anti [Wra08, Len05]. anti-Platonism [Len05]. Anti-realism [Wra08, Wra21]. antinomy [Sl91]. Antiquity [Zem13]. Antireductionism [Roh04]. any [Nee04]. anything [Mil91a]. Apology [Ano92a]. aporia [Sal93]. Applicability [Rus13]. Applicable [HGS24]. Application [Sve14, Bra94]. applications [Liu98]. Applied [Boo06, Mor06, Yag17]. Applying [Bod06, Hei06, Mor06, BBB06]. appreciation [Rue02]. Approach [Bar20, BB21, Che23b, Mal07, Wag13, YTH23, ZG16, Bra94, FL99, Ler01, Sce98, dCF93]. Approaches [Gon12, Rus10]. Approximate [Hun11]. Approximation [CT19]. apriorism [Mis01]. Arber [Fla03]. archaeological [Wyl88]. Archaeologies [Bab15]. Argument [All17, And15, Bel07, Bus11, Ken11, Miz13, Reu13, Saa07, Sko10, WĆ06, Wra08, YKMM22, DD02, Goł99, Gre94, Mat98, Nad93, Sik92]. argumentation [Feh98b]. Arguments [Che23b, Mat21, Feh98a, Hoe00]. Aristotelian [MR88]. Aristotle [Len01, Mam16, New01, Uga15, Ano02a]. Armstrong [Pag02]. Arrow [CGL05, Reu13]. Art [Ban12, Elg02, Law93]. Artefact [JB15]. Arthur [Zha15]. Article [Ci03, Lel88, Rad02]. Articulating [Rou11]. Articulation [Lyo13]. Artifacts [Fra09]. Artificial [Kee12, YH24, Jac93]. Artifictional [Tym21]. Ashgate [Kno11]. Aspects [Rea20, Cul95, Got00]. Assess [CL19]. Assessing [Kra21]. Assessment [CG12, Mäk13, Pay15]. assisted [Aga92]. assumption [Wil91]. assumptions [Wyl88].

Astronomer [De 15b]. Astrophysical [WĆ06]. Astrophysics [Suá13]. Asymmetric [CGL05]. Asymmetry [Pal16]. Atom [Bor13]. Atomic [Bor13]. atoms [Mac01, Nee04]. attack [Feh98a]. Attempt [PV20, Sch09, Mat87]. Auction [Bol12]. Author [Ano10, Ano11, Ano12b, Ano13b, Ano16]. Autonomous [Obe90]. Autonomy [Kuu12]. Axiomatic [And15].

B [Duc05, Hue25]. B/W [Hue25]. Bachelard [Cut88]. Background [Kow17, Mit89]. Backtracking [Kut11]. BACON [Wat92]. Bad [Feh98a, Kha10, Sch22, Sha16, Nad93]. Bailer [Mac07]. Bailer-Jones [Mac07]. Bakar [Nie08]. Balancing [Til99]. Bananaworld [Duw16]. Bandyopadhyay [How16]. Barrier [Bre15]. barriers [Pea02]. Based [Bae10, Chi19, Kou15, McF92, The14]. Basics [Kee12]. Basis [Per14, Vah94]. Bayesian [BTB16, Bra17, HGS24, How16, Réd92]. Bayesianism [BB10, DS12]. BC [Uga15]. Be [Bou19, Ci06, Reu13, Tod08, Bus03, Cha97, Hon91, Lel89, Mis96, Nii90, Réd92, Tuc90, Ueb01]. Bearers [DF94]. beautiful [Eng02]. Beauty [De 02, Kos02, Tod08]. Before [HS09]. begin [Nee04]. Beginning [Zin08]. Beginnings [Bab10]. Behaviour [VG15]. Behaviourism [Sle91]. Belief [Nor07, Oka00a, Psi00, Vah94]. Beliefs [Sch22, Lin93]. believe [Per13b, Vai98]. Believing [Alc10]. Bell [Fer11]. Beneath [Tod08]. Benefit [Dou00]. Bentley [Pau12, Mac01]. Berkeley [Pap88b]. Bernard [LS94]. best [Jae02, Mar89]. best-system [Jae02]. Between [Bre15, Del22, Kin17, Okr98, CW09, Dor07, Kut10, Mac92, Ses11, Yan08]. Beyond [Ban12, Del22, PS23, Sle09, Sve14, Whi15]. BIC [HGS24]. Big [Row08]. bin [Nie08].

Bio [Tam25]. Bio-Inspired [Tam25]. Biodiversity [Ese09]. Biography [Bab15]. biological [Tho92, Tis89, dS89]. Biology [Ano02a, Bae25, Hue25, Ken11, Mar15, Yan08, Jan89]. Biomedical [Leo12, YTH23]. Biorobotics [Tam25]. Bird [Row10]. Birth [Mel23, Bol87, Feh88, Til87b]. bivalence [Mis90]. Black [Bok11]. blend [Lel89]. Bloch [Woo07]. block [Ear08]. Blocking [Pil23]. Bloomsbury [vdS14]. Bloor [Sle91]. blue [Law93]. bluff [Sle91]. Board [Ano12a, Ano13a, Ano14b, Ano87a, Ano87b, Ano88b, Ano88c, Ano89a, Ano89b, Ano90a, Ano90b, Ano90c, Ano91a, Ano91b, Ano91c, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano94a, Ano94b, Ano94c, Ano95a, Ano95b, Ano95c, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano98c, Ano99a, Ano99b, Ano99c]. Bodies [Chi10a]. Body [Mor11]. Bohm [Acu16, Myr03]. Bohmian [Sko10]. Bohr [Bor13, How18, MA13, Rio17, vS24]. Book [Abe91, Ala24, AKLV98, Ano00a, Ano00b, Ano00c, Ano01a, Ano01b, Ano02a, Ano02b, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano05a, Ano05b, Ano05c, Ano06a, Ano06b, BSU97, BBD09, Bor13, Bus19, CHLL07, CIL96, CBG06, Cla13, CJD^{+}08, DLVW08, Dud17, Duw16, HF92, HMO^{+}99, Hue25, Ili91, JSJ^{+}94, JBM^{+}95, Kno11, KSA07, LB98, LPR99, LLOA93, Man90, MV90, MWRB98, OKHT96, Okr98, Orr24, Pau12, PLO94, Pla10, PS92, STL^{+}96, Sia19, Str17, SM91, VHO^{+}07, WR93, WMB97, YTP08, Zha15, vdS14, Mam10]. Bose [Das12]. Boskovic [Cul95, Cul93, Mar90]. Boudon [Aga93, BR93, Fuc93, Kli93, Lin93, Mon93a]. Boundaries [Pao17, Lon97]. boundedness [Kuk93]. Boxes [Kas04]. Boyle [Mac92, Mac01]. Bradford [Tre21]. Bradley [Man91]. Braess [Irv93, Mar96]. Brain [Abr16, Swa11]. Branching [DL23]. brand [Bro88]. Break [Hop13]. breakdown [Pea98]. breaking [Ear04]. Bridge [CE11].

Bridging [Per12]. Brittan [How16]. Broglie [Acu16]. Bruno [Dud17]. Bub [Duw16, Pea02]. Bugbear [Kow17]. Bugbears [Bab10]. Butterfield [Pau12]. butterflies [Til90b].

calculus [dS91]. called [Fri03]. Calling [Bab15]. Cambridge [Ano02a, Bus19, Duc05, Dud17, Orr24, Pau12, Zha15]. Can [Bou19, Bul09, Car23b, Cha97, Ci06, Hon91, Lel89, Pet23, Reu13, Sha16, Tuc90, Ueb01, VG15, Web09, Bro89, Nor00, Nug91, Piz93, Réd92]. Canada [Irv97]. Cancer [Mal07, Mar05, Tha02]. Capacities [Hoh03]. Capital [Des18]. Care [And20a, BV14]. career [Tem96]. Carl [Ano98d]. Carlo [BN12]. Carnap [Ueb01]. Cartwright [Okr98, Orr24, Spu01]. Case [Cru24, Gon15, Leo12, Pie12, Ruz11, Feh98a, Mar89, Rue02, Tem96]. Cases [CACA19, Pay95, Sik92]. Cat [Okr98, CS24]. Categories [Hul14]. Category [JWDD12]. Catherine [Bai98]. Causal [Ala24, Bau09, Ben20, BG12, CG12, CMF14, CMF16, Kuo09, Lan14, Par15, Pin10, Pop16, RA16, Rus10, Via99, Vic06, Way15, Wil19, Woo04, Dow92, Feh87, Mil91a, Mit89]. Causalism [dR06, Tuc90]. Causality [Bog04, But07, Buz14, EV14, Ort24, RW07, Tei13]. Causation [Bae25, Cam14, HK14, Mac04, Par16b, Per13a, Per14, Web09]. Causes [CE11, Kat13, Mel09, Pap90]. Cavendish [Par16a]. Century [Leo11, Uga15, Zem13, Car87, PZ14, Rio17, Rue02, Ara20]. Ceteris [Kow14, Kow17]. Chagas [Joy13]. Challenge [Gir23, Kin11, Par15, Why25, vS24, San96]. Challenges [And20b, CACA19]. Cham [Ala24]. Chance [Bae13, Pal16, Woo07]. chances [Mel90]. Chang [BV24]. Change [Che23a, Dim16, McA08c, Roe17, She21, Ven21, Mar99, Mel23]. changes [PZ14]. Changing [Kre17, PS01, Ste11]. Character [Ken22]. Characterise [Yag17].

Characteristic [GY11]. Characterization [Gon15, Pin06]. Charades [Hun11]. Chemical [Thy22, BV24]. Chemistry [Sei16, Sim16, Car87, Nee99, Nee00, Nee04, Sce98]. Chemists [Thy22]. Cherries [Miz15]. Cherry [Miz15]. Cherry-picking [Miz15]. China [Sl91]. Chinese [Qin16]. Chirality [Hoe00]. Christian [Phi14]. circa [De 02]. Circularity [Ada04]. Circulation [SS15]. Citation [YTH23]. Citation-Exemplar [YTH23]. Claim [Bul09]. Claims [Hop13, Wil19]. Clarke [Mac01]. Classical [All17, Bae10, Dar07, How18, Kaz14, Lar06, Pie12, Cha97, Hug99, Lar01, Tem91, dS91]. Classification [Brz16, Hul14, dP15]. Classifications [Tso07]. Classificatory [Leo12]. Claude [LS94]. clear [Law93]. Clifton [Bub03]. Climate [CL19, RS10]. Climategate [RS10]. Closer [Bru13, Miz17]. coexistence [Pea98]. Cognition [Käs15, RR95, LM95, McF91, RB02, Via99]. Cognitive [Kha05, Kou15, Mar87, PZ20, Roh04, The12, Was08, Mil91b]. Cohen [Duc05]. coherent [Lel89, Mat87]. Cold [De 92]. collapse [Fri03, Lew03]. Collection [Thy22]. College [Nar01]. collision [Lar01]. Color [Chi10b, Dan20]. Colour [JWDD12]. Combining [Bri10]. Comment [Hes92, Aga93, Kli93, Duc05]. commentary [Tra92]. commentators [JL94b]. Comments [Aga92, BR93, Gil92, Hun94, Lin93, Luc97, Mon93a, JLL92]. committal [Pin10]. Commodification [Jac11]. commonsense [Mis01]. Communities [Roe17]. commutative [Réd92]. Companion [Kno11, vdS14]. comparative [Pec02, Til98, Wil91]. Compatibility [Iva11]. Compatible [Rus10]. competitor [Lyo09]. Complementarity [Bok11, dR06]. Completeness [Vic06, Spu99]. Complexity [The14]. Complexity-based [The14]. Component [Moo12]. Componential [Kuo09]. composition [Spu01]. Computer [BK14, Dur17, Jeb17, Kro08, Par08, Aga92]. computer-assisted [Aga92]. computerized [Aga92]. Computers [JM16, Tym21]. Concept [Des18, Tho14, Ven21, Gon95, Tet89, Bus19]. Conception [And15, Liu15, Ci03]. Conceptions [PB14, PS01]. Concepts [Bae10, Kuo09, Wag09]. Conceptual [All13, Bab03, Bel07, Mel23, Rou11, Ven21, ZG16, Hac95, PZ14]. Conceptualisation [Jod24]. Concerning [SS14]. Conclusive [Mac00]. Concrete [Leo11]. Condition [DS21]. Conditionals [Suá13]. Conditions [Bul09, Lyo09]. Confirmation [Bra17, Bri10, Ped18]. conflict [BR93]. Connection [Col22, Kut10, Mil91b]. connectionist [dS91]. Consciousness [Nes22]. consequence [Hug99]. Consequences [Ral07, Wol15]. Conservation [Des18]. conservatism [Res94]. considerations [Eng05, Ram93]. Consistency [Bok11]. constant [Cha97]. Constituting [Pea10]. Constitution [Har10]. Constraint [Lan23]. Constraints [JWDD12, The14]. Constructing [Mar01]. Construction [Mar08, RS10, Lep88, Ram93]. constructive [BO91]. Constructivism [Rea20]. Contemporary [Kno11, WĆ06]. Content [NL16, Pri00]. Contents [Ano10, Ano11, Ano12b, Ano13b, Ano14a, Ano16]. Contested [Bus19]. Context [Mam16, Run11, RH05, O’H94]. Continent [Joy13]. Continental [Bab10, Nel20]. Contingency [Kid13]. Continuity [ZG16, Mis01]. Contours [McA08a]. Contradictions [Kaz14, Mam10]. Contributors [Ano91f, Ano95d, Ano87c, Ano87d, Ano88d, Ano88e, Ano89c, Ano89d, Ano90d, Ano90e, Ano90f, Ano91e, Ano91g, Ano92e, Ano92f, Ano93d, Ano93e, Ano94d, Ano94e]. Controlled [Kra21]. Controversial [Pan17]. Convention [Ban12]. Conventional [Sed96]. Conventionalism [Kle01, Pet93a]. Convergence [Bae10]. Conversations [May23]. cooperation [BR93]. Copenhagen [MA13]. Copernican [And91]. Copernicus [De 15b]. Corpuscular [New01]. correction [Bla95]. Correspondence [Ano03d, Ano05d, CS24, Lod01]. Corroboration [Row08]. Cosmic [De 15b]. Cosmopolitanism [Rol15]. cost [McF92]. cost-based [McF92]. could [Bus03]. count [Har88]. Countered [BTB16]. Counterexample [BTB16]. Counterfactual [Bog04, Was08]. Counterfactuals [Ful22, Woo04]. counterparts [Bri99]. Counting [Fri17, Fri03, Lew03]. countries [Car98b]. cover [Ala24]. Creating [Das12]. Creation [Elg02]. Creations [Fra09]. creativity [Par90]. Credence [P´ 18]. Credibility [Wil23]. Crisis [And20b]. Criterion [BG12, HGS24]. Critical [Alm07, Ano01c, Auf12, Bab10, Bai98, Des18, Eme17, Pay15, vdB14, Dav98]. Criticisms [The14]. critics [Bou93a, Sim92b]. Critique [Alb06, Lyo13, OM16, BO91, Mis01]. Cube [Bog20]. Cultural [Bre15]. Cure [And20a]. Curie [Ear04]. Curing [Tha02]. Curious [Ben22, Lan11]. Cutting [Han09].

Dadaism [Ken22]. Daniel [New01]. Daniela [Mac07]. Dartmouth [Nar01]. Darwinian [Oft10, SS15, She21]. Darwinian-Selectionist [She21]. Data [Leo12, Oka00b, Zib94]. Data-intensive [Leo12]. Davidsonian [DD02]. day [Mam16]. dealing [Lew03]. Debate [Che23b, Lyo09, NL16, Ort24, Rea20, Rup05, Wra21, Dav99, Lod01]. Debates [Den15, Ken11, Run11]. Debating [Hop13]. Debias [Pet23]. Decades [She07]. December [Par24]. Decision [Ede10, Bra94]. Decisive [Sim19].

Deduction [PB14, Mos94]. deductive [New94]. Default [Kha10]. Defeat [How16]. Defence [Bae21, DS12, Ful22, Wra08, Dow92]. Defending [Pin24]. Defense [Che23b, Car98a, Mis01]. Defining [McF91]. Definition [Mäk13, Ray09]. Deleuze [Wil11, vdT14]. Deliver [Per13a]. Demarcation [BB21, Fas18, Han09, Let19, Yan08]. Demetra [Bro94]. Democracy [Ci11, Sl91]. democratic [Ci03]. Demon [All13]. Demonstrative [Mag08]. DENDRAL [Til90a]. Denialism [Bab15]. Denkstil [Bab03]. dependence [Ada04]. Depth [Gij09]. Descartes [Kva16]. description [DF94, McA97a]. Descriptions [Lau15]. Design [Ram93, WĆ06]. Designation [Ina14]. Despite [Roh04]. Determination [Car10a, SC13]. Determining [HGS24]. Determinism [Bel95, But88, Gom91]. Developing [Bar94]. Development [Jod24, Käs15, Svo15, vdB14, Bol87]. Developmental [CW13]. Deviant [Han20]. diagonal [Sik92]. Diagrams [Car10b, Mey08]. Dialectical [Kaz14, Ken22]. Dialogue [BB21, The12]. Diatribe [Del22]. Dichotomy [She21]. Dictator [JB15]. Did [Zin08, Nee04, Nug96]. Difference [Mor06]. Difficulties [Mäk13]. difficulty [Les96]. Digital [Kro08]. Dilemma [FHS21]. dimension [Til98]. Disambiguating [Ill11]. Disbelief [Nor07]. Disciplines [DL23, Han20, Tam25, Yan08]. Discorsi [Feh98b]. Discovering [SH92]. Discovery [Auf17, Bri10, Gil92, Mar08, Mar10a, Paa06, Roe17, Sha07, Vil23, De 92, Hes92, Los92, Nee02, NS92, Pet92, Pet93a, Sim92a, Sim92b, Tra92]. Discovery/Justification [Auf17]. Discriminant [Fas18]. Discrimination [Pet23]. discussion [How11]. Disease [Joy13]. Disfluency [Pet23]. disparity [Eng01]. Dispositional [Tug13]. Dispositions [Kat13, Mel09]. disputes [Ueb01]. dissolve [Wil95]. Distance [Par15]. Distances [Car10a]. Distant [HS09]. Distinction [Auf17]. distinctive [Sep25]. disunities [Dav96]. Disunity [CS24, Hoh03, Len01]. Divergence [DL23]. Diversity [Ci11]. Do [Cas17, Bro89, Mil91a, Nee04, Piz93]. Doctrinal [Let22]. Doctrines [Han20]. Does [BV14, Jan89, Lyr08, Per13a, dS91]. dogma [Rei01]. Dogmas [BB10]. doi.org [Hue25]. doing [Das12]. Dominance [JB15]. domination [Bak91]. Dordrecht [Rad02]. Double [Fri17]. Doubt [Dou00]. Doubts [Mar92]. Downward [Par16b]. dream [Til88]. dreams [Til90b]. Dretske [Pag02]. Droplets [VG15]. Dual [Nor07]. duality [Har97c]. Dubrovnik [Bro99a]. due [Zib94]. Duhem [Sha20, Wei95]. Dynamical [Dor07, Rea20]. Dynamical/Geometrical [Rea20]. Dynamics [Hei06, PZ14, Rue03, Tem91].

Early [Bai98, Mar16, Lar94, Myr03].

Earman [Hoe00]. Earth

[De 15b, KBdR05, vdT14]. EBM [Tre21]. ecology [Gil91]. Econocracy [Bag17]. Economic [CACA19, CMF14, CMF16, BO91, Bol12]. Economics [Bag17, JB15, Rol15, Mon93b]. Ecuador [SS15]. edge [Pea98]. editor [Ano03d, Ano97d, Ano02c, Ano03f, Ano05e, Ano06c, De 15a, FD10, HS15, McA07, McA08b, McA09, McA10, McA11, McA12, McA13, McA14, McA15, McA16, McA17]. Editorial [Ano87a, Ano87b, Ano88b, Ano88c, Ano89a, Ano89b, Ano90a, Ano90b, Ano90c, Ano91a, Ano91b, Ano91c, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano94a, Ano94b, Ano94c, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano98c, Ano99a, Ano99b, Ano99c, Ano03e, Ano03f, Ano04c, Ano12a, Ano13a, Ano14b, Bro01, Kin18, Kin19, Kin20, NS95, Wil92, McA01]. Eds [Pla10]. Educators [Thy22]. Effectiveness [Fil16, Tre21]. Effects [Tso07]. efficacious [Til99]. Effluvia [Par15]. Ehrenhaft [Col22]. eighteenth [Car87]. eighteenth-century [Car87]. Einstein [Kra06, Bro88, Das12, DiS92, Eng02, Eng05, Flo99, Flo05, Gha17, MHB22]. Electrodynamics [Pie12]. Electron [Cla13]. Element [Thy22]. Elena [Par24]. Elimination [Bla95, Bai89]. Elongations [Rec17]. Elusive [DS21]. emancipation [Bak91]. Embodiment [Pin11b]. Emergence [HT00, Pin11b, The14]. Emergent [Vic13]. Emerging [Sch11]. Emotions [TK10]. Empirical [Ale09, Miz20, Pao17, Sar00, DiS92, Sep25, Vah94]. empiricality [Ada04]. Empiricism [Gao17, Uga15, Psi00, Rei01]. empiricist [BO91, Dow92]. Enantiomorphy [Rom05]. Ends [Let19]. Engineering [Sch11, Zhu17, Mar99]. English [Pet23, Wol15]. Enough [CT19, Miz17]. Enquiry [God14]. Entanglement [Lam13]. Entities [Ye10, Bro89]. Entrenched [YTH23]. Entropy [Reu13]. Environmental [OM17, Vaz16]. Epigenesis [vdB14]. Epiphenomenalism [Per13a]. Epistemic [Bar20, DL23, Fol87, Fre06, Käs15, Kuk93, PV20, P´ 18, Rol15, Row10, Sve14]. Epistemological [Gon12, Rat05]. Epistemology [Ben17, Bre15, Gao15, Leo11, Mac04, Par08, Ren13, Sha20, TK10, Tam25, Bla95, Mac92, Res94]. Epochal [Hop13]. Equation [Flo05, Sko10, Flo06, Kra06]. equilibrium [Bac92, BO91]. equivalence [Sar00]. Ernest [Che23a]. Errata [Ano91d]. error [Bai89]. Errour} [Hon91]. Erwin [Mal12]. Essay [Chi10b, Mar92]. Essays [CL11, Pin11a, Pin11b, Pla10, Sei16, Thy22]. Essential [Hoc13]. essentialism [Fuc93]. Essentialist [Bus19]. Establishing [Wil19].

Ethics [OM17, Wal13, Zhu17]. Euler

[PZ14]. EUR [Ala24, Dud17, Rad02, Zha15]. Europe [Bjö12]. European [Gon15, McA08a, Rad15, SS15]. evaluation [Got00]. Event [Bri17]. events [Har91]. Everett [Les96]. Evidence [BTB16, CG12, Fri17, How16, Ill11, Tei13, Vaz16, Web09]. Evidential [Bog20]. Evolution [Boo16, Bou19, Hue25]. Evolutionary [Gon12, Lel89, Ren13, Bla95, O’H94, Rat05]. Evolving [Yan08]. Exactly [Gir23]. Exception [Koc09]. Exceptional [Nel20]. Exclusion [Bau09]. Exemplar [YTH23]. exile [Ueb01]. Exist [Lyr08]. Existence [Sch17, Ghi03]. Existentialism [The12]. existents [Gol93]. Experience [Dan20, Uga15, Vah94, Gil91, Got00]. experiences [Vai98]. Experiencing [Vit16]. Experiment [Jeb17, Par16b, Per21, And91, Lep88, Tis89]. Experimental [Ale09, Bae13, Boo11, Che23b, Col20, Gao15, JB15, Rin12, Rou11, Goo90]. experimentation [Boo04, LS94]. Experiments [CACA19, Col22, Yan13, Art99, BN12, Bon97, Bro87, Gra18, Kuj98, Mis92, Rue02]. Experts [Bag17]. Explain [Kro08]. Explaining [Iva14, JK15, Alb93]. Explanation [Ber90, Bri10, Dan20, Dur17, For96, Gij09, Jan15, KBdR05, Oka00a, Par16b, Ral07, Rat89, She21, Tei13, Woo04, Bro89, Cor89, Cul95, Det89, Dow92, Feh87, Flo99, Hoh03, Hun01, Jon00, Mar89, Mit89, Qiu89, Sea87]. Explanationist [Saa07]. Explanations [Dor07, Har10, Lan23, Pin10, RA16, Rös89]. explanative [McA89]. Explanatory [Bul09, CF06, Dav96, Dur17, Ful22, Kha10, Mak23, Way15, Nee04]. Exploring [HS09]. Expose [Sha16]. Extended [Ci06, Lod01]. Extensionalism [Cha11b]. extent [Cra00]. External [Ano24, Yan13]. Extra [Dan20, Why25]. Extra-Mathematical [Dan20]. Extra-Terrestrial [Why25]. extragalactic [Roc98]. Extrapolation [Bae21]. eye [Fla03, dS91].

F [Flo06]. F. [Kra06, Man91]. Faces [Cru24]. Facilitates [Fas18]. Fact [Bab15, Nor00]. Facticity [Ral07]. Factivist [Gur22]. Facto} [Gol20]. Facts [RS10]. Factual [Bog04]. fails [Mat98]. Failures [Bae25]. Fake [Car23a, Car23b]. fall [Bak91]. Falls [Auf17]. Falsification [Tra20]. Far [VG15]. Faradayan [Dar07]. Fear [Per13a]. Features [YH24]. Feeding [TP23]. Festschrift [Har97a]. few [Cor92]. Feyerabend [CS24, Col22, Cru24, Del22, Far00, Ken22, Kid16, KF22, SS22, SS24]. Feyerabendian [Ori24]. Feynman [Mey08]. Fiction [Per13b]. Fictionalism [Liu16, Pap88a]. Fictions [Pin11a, Suá13]. Field [Dar07, EV14, Lam13, Sim16, Ven21, Eng01, Goł99]. fifth [Mar89]. Findings [Car23a]. Fine [Fri17, WĆ06]. Fine-tuning [Fri17]. Finite [And15]. first [Rio17]. Fit [HGS24]. Five [Chi10a]. Fleck [Okr98, Bab03, Bab15]. flight [Bis03]. Flores [Kra06]. Fluid [Hei06, PZ14]. Focus [Ara20]. Force [Moo12, Wil11, Mar89, Nug96]. Forces [Moo12, Liu03]. Fore [Gin12]. Fore-structure [Gin12]. forerunners [Ler01]. Forking [Bic22]. Form [Kuo09, Cha04, Fla03]. Formal [Bac11, Bul09, Kaz14, Sza03, Ste10]. Formalisation [DS21]. Formation [JWDD12]. Forms [Sci16, Mos94, Til88]. forward [Far00]. Foucauldian [Mar16]. foundation [Jan87]. Foundations [All13, Del22, Fel14, Fer11, Ori24, Sia19, DiS92]. Four [Lew03, Wra15]. Fraassen [Gro91, Kin11, Mar89]. Fractal [Tem96]. fragmented [Mis96]. framework [Hac95]. Franca} [Wol15]. Franklin [Par08]. Fraternity [Yan08]. Free [Sch12]. Frege [Ant12]. Fregean [Det89]. Freud [Gom91].

Friedman [Dim16, Iva11, Kin11]. frigorifick [Mac01]. frontier [Lin93]. Frontiers [PS23]. Function [All17, Laz17, Mat21]. Functional [ES07, Ful22, Kha05, Mit89]. Fundamental [Kos00, Per14, Spu99, Nug91]. Fundamentality [Fil16]. Further [Rea20]. fusion [De 92].

G [Ano02a, Cla13]. Gadamerian [Alb06]. Galilean [All17, Kut10]. Galileo [Feh98b, Kva16, Mat87, Sea87]. Game [JB15]. Gap [Per12]. Garden [Bic22]. Garg [Cru24]. Gaston [Cut88]. Gauge [Fil16, Liu03]. GDR [Sch98]. Gender [Kou23]. Gene [Bae10]. Genealogy [Sim16]. General [Bod06, Kat13, Pri00, Bro97]. General-purpose [Pri00]. Generalisations [CMF14, CMF16]. genesis [Nug88]. Genetic [CF06]. Genetics [Pio14]. Geometric [Lam12]. Geometrical [Rea20]. Geometry [Gau05, Sia19, DiS92, Kle01, Tem96]. Gerard [Dud17]. German [Car98b]. German-speaking [Car98b]. Getting [Miz17, Pih96]. Ghins [Ala24]. Global [Lar06, Ruh13, RS10]. Globalized [Wol15]. Glock [Pla10]. goals [LM95]. God [Mac01, Pau12]. goes [Loc97]. Goethe [Hue25]. Gold [Wei16]. Gone [HS09]. Good [BV24, Sch22, Sha20, Bou93a, Cre93, Feh98a, LR93, Nad93]. Gordian [Han09, Let19]. Grammar [Cha11a]. Graph [Tug13]. Graph-theoretic [Tug13]. Graveyard [Miz16]. gravitation [Lar01]. Gravitational [Roc98]. gravity [Liu03, Nug88]. Greek [Zem13]. Gregory [Hue25]. Grounds [Wra08]. Group [All17]. growing [Ear08]. Growth [Mam10]. Guidance [Sko10]. Guide [Cam14, Run11].

H [BM22, Man91, Nee02]. H. [NS92].

Haack [Vah94]. Hacker [Pla10]. Hacking [Tso07, Mar16, Roc98]. Hackingian [Che23b]. Handbook [Chi19, Ruz11]. hands [Hoe00]. Hans [Pla10, Kle01, Phi14]. Hans-Johan [Pla10]. Hansonian [Paa06]. Happen [Sch22]. hard [Ala24]. hardback [Ano02a, Bor13, Bus19, Dud17, Duw16, Pla10, Rad02, Zha15]. Harmanian [Paa06]. Harmful [Car23a, Wil23]. Harmony [Saa07]. Harms [dMM23]. Harré [Pay95]. Hasok [BV24]. Healey [Rob92]. Health [CG12, RW07]. Heaviside [Bol87]. Heidegger [Sch17]. Heisenberg [Cam05]. Helge [Bor13]. Hell [Bus11]. Helmholtz [Hac95]. Help [Pet23]. Hempel [Ano98d, FHS21]. Herbert [Gil92, Pau12, Hes92, Los92, Tra92]. Herder [Hue25]. Here [CE11]. Hermann [Gau05]. Hermeneutic [Gin12]. heroes [Til90b]. Heuristic [Aga92, Vil23, JLL92, Mam87, Pet92]. Hidden [Pie12]. hierarchies [Det89]. Higgs [Lyr08]. Higher [HK14]. Higher-level [HK14]. Hill [Tre21]. Historian [Ara17]. Historians [Thy22]. Historical [Bre15, Kid13, Miz15, Rös89, Car87]. Historically [Lyo13, Wag13]. historicism [Rey99]. Histories [Leo11]. Historiography [NL16, Kut87]. History [Abr16, Ara17, Cla13, Joy13, Lar94, Lew16, Miz16, Par20, Pau12, Qin16, Sch09, Val12, YTH23, Zem13, McA97a, Sch98]. Hole [Bok11, YKMM22]. Holism [CS24]. holistic [Rob92]. Holmes [Par08]. Homeopathy [Bab15]. Homer [Til90a]. Horizon [Mar10a]. Howell [Mar01]. Howson [BTB16, Koc09, Ste11]. https [Hue25]. Human [Bus19, Hul14, Sci16, Alb93, Par90, Til90b]. Humanities [Bor12]. Humanity [Tym21]. Hume [How11, Sho18, Ste10]. Humean [Jae02]. Hunting [CE11]. Hybrid [Liu16]. Hyman [Pla10]. Hypotheses [dP15, Got00, How16]. Hypothesis [DS21, Miz20, Vil23].

Ian [Mar16]. Ibn [Ray09]. Iconic [Amb14]. Ideal [Kra97]. Idealization [Pin11a]. Idealizations [JK15]. Ideas [PG06, SS15, Car87, De 92]. Identification [Mäk13]. Identities [Ina14]. identity [Che02, vBG88]. If [Ste10]. Ignorance [Kou23]. II [Liu98, SS24]. Illustrations [Hue25]. Image [Per12, Sim19]. Imagery [Mil91b]. images [Til87a]. Imagination [Per13b]. Immature [Hib16]. Impact [Kid13, Car87]. imperatives [Nii90]. Imperialism [CW09, CW13, Kid13, Mäk13, Rol15]. Implications [Tra20, Wyl88]. Importance [Fey22]. Imprints [Mar16]. Incident [Ben22]. Inclusive [Rad15]. Incognita [KBdR05]. incoherence [Kuk95]. Incommensurability [Alb06, Kin17, MS09, Pol17, Zhe88, Bab03, DD02, Hun01, Mil91a, San98, Mil91a]. Incongruent [Bri99]. Incorrigible [Let22]. Independent [Che23b]. Indeterminism [Lar01]. Index [Ano10, Ano11, Ano12b, Ano13b, Ano16]. Indexicality [Ruy21]. Indifferent [Tso07]. Indigenous [Wag09]. Indispensability [Bus11, Rus13]. Indistinguishable [Ben22]. Individuality [dS89]. individuation [Har91]. Induction [Gol20, Gre94, Mag08, Mos94, OM94, Sho18, Ste10, Wra21, Bar94, Bis03, Coh94, Gir94, Hun94, JL94a, JL94b, Mül15, Smi94]. Inductions [Miz15, Wra15, Ruh13]. Inductive [Hun11, New94, Gir94]. Industry [Gha12]. Ineffable [Zwi12]. Inequalities [Cru24]. Inertial [Acu16]. Inference [Bri10, Isl17, Kha10, Mag08, Mar89, Pag02, Réd92]. inferences [BN12]. Inferential [Bae21, Bic22, New12, New13]. Inferentialist [CMF16]. Inferring [Lan14]. inflating [Liu15]. Influence [Kut11]. Information [Fel14, HGS24, Sar16, Søe19, Til90a, Tet89]. Informational [Ben20]. Informed [Lyo13, Wag13]. Inheritance [Bou19]. Insights [Ori24]. Inspired [Tam25]. Instituting [Mar08]. Instrumentalism [PZ20, Sea90]. instruments [Boo04, Fje91, Kut87]. Int [Flo06, Kra06]. Integral [Rad02]. Integrate [Sch09]. Integrated [YTH23]. Integrating [BR93]. Intellectual [Bro98, Mal12]. Intellectuals [PG06]. Intelligence [Kee12, Tym21, Why25, Jac93]. intelligibility [Luk93]. intensive [Leo12]. Intentional [Hoc13, RR95]. intentionality [Why90]. Intentions [Gir23]. Inter [Bae25, Mar90]. Inter-Level [Bae25]. inter-theory [Mar90]. Interaction [Kuo09, Lar06, Whi15, Mac92]. Interactions [Bok11]. Interactive [Bus19, Tso07]. Interdisciplinarity [Alv11, Pol17]. Interdisciplinary [CW13, TP23]. International [Ano15, Ano17]. Interphenomena [Bog20]. Interpretation [All17, Jan15, MA13, Per12, Thé15, Ber90, Flo06, Kra06, Pec02, Rob92, Vai98]. Interpretations [Bar20, Flo05, Kra06, Fri03]. Interpreting [RW07]. Intervention [Mak23]. Interventionism [EB14, HK14, Reu13]. Interventionist [Bau09, Par16b, Per13a]. Interventionists [Reu13]. Interventions [Was08]. Intrinsic [BG12]. Introducing [Lin95]. Introduction [And20a, BBB06, Bro99b, Bry12, Car23a, Ekw15, Eme17, Har97b, Kap10, Leg94, NSW87a, NSW87b, NSW88a, NSW88b, NSW89a, NSW89b, Pau16, Pet93b, Pio14, Rin12, TO22, Wal13, Zhu17]. Intuitions [Lor15]. Invariance [Rus14, Ear02]. invention [Bai98, Gra18]. Investigation [Auf12, Bab15, Hac95]. invisible [Bai98]. Iran [Nav07]. Iranian [PG06]. irreality [Roc98]. Irreversibility [CGL05, Jod24]. ISBN [Ala24, Ano02a, Bor13, Bus19, Dud17, Duw16, Hue25, Orr24, Pau12, Pla10, Rad02, Zha15]. Islamic [Pay15]. Islamisation [Pay15]. Isolation [GY11]. isomorphism [Sua03]. Issue [Car23a, Bro99b, Wil92]. Issues [And20b, Kou15, Mak23, Liu98]. ivory [Mam90].

J [Cla13]. J. [Bra94, Har97a]. James [Ano02a]. Jeffrey [Duw16]. Jesus [SS15]. Johan [Pla10]. John [Pla10, Gat00]. Johnson [Coh94]. Jones [Mac07, Oka00a]. Jordi [Okr98]. Journey [Boo16, Roe17]. Judgements [Tod08, Log97]. Judgment [Lan11]. July [Par24]. Jung [Kar15]. Jure} [Gol20]. Just [Lan23]. Justice [Bjö12]. Justifiable [Sha07]. Justification [Auf17, Käs15, Nee13, San96]. Justificatory [Sch09].

Kant [Hoe00, Wat92, vdB14]. Kantian [Cam05, Fje91]. Kantianism [Dim16]. Karl [PG06]. Kathleen [Ano06d]. Keynes [Bra94]. Kind [Bir12, Nee12, Tso07]. Kindness [BJS17]. Kinds [Brz16, Ful22, God14, Hul14, dS89]. Kitcher [Psi97]. Kluwer [Rad02]. Knot [Han09, Let19]. know [dS91]. Knowability [CL11]. Knowledge [Der10, Gha12, Iva14, JM16, Koc08, Kou23, Kow17, Lin22, Mam10, Mar08, O’H94, Pay15, P´ 18, Sch17, SS15, Was08, Ant97, Bol87, Kra92, Mac92, Oka00b, Pet93a, Rec17]. Kochan [How09]. Kragh [Bor13]. Krajewski [Big07, Flo06]. Kries [Hei01]. Kripke [How18]. Kronfelder [Bus19]. Kuhn [Alb06, Bab03, Ghi03, Kuu09, May22, May23, Mel23, Pol20, PS23, Ren13]. Kuhnian [Hun01, Kin17, Sim19]. Kuhnianism [Dim16].

laboratory [Mar01]. ladenness [Fje91]. Laird [Coh94]. Lakatos [Bro90, Sep25]. Lakatosian [Kuu17]. Language [Gil91, Kva16]. Last [JB15]. Latour [Bab15, Dud17]. Laudan [Feh98a, Kno02, Pay95, San96]. Lawlike [BG12]. Lawmakers [Vet11]. Laws [Ala24, Gol93, HS09, Hun90, Jae17, Kow14, Kow17, Lau15, Max09, McA97a, Mel90, Reu11, SM94, Vet11, Bro94, Car98a, Jae02, Nor00, Nug91, Pag02, Spu99, Spu01, Why90]. Layers [BK14]. leads [Ram93]. learn [Nor00]. Learning [Ste10]. least [Feh88]. Leaving [Bag17]. Lee [Tha02]. Left [Bus19]. Legacy [Buz10, LS94]. Leggett [Cru24]. Leibniz [Fut02, Gom91, Lod01, Mor11, Zha15]. Lelas [Rad02, Ano03g]. Lennox [Ano02a]. lensing [Roc98]. Lesson [Ken22, Rös89]. Letter [Ano03d, De 15a, FD10, HS15]. Level [Bae25, HK14]. Levels [Par16b]. Lexicons [Pol20]. Library [Ala24]. Life [Ano02a, Far19, Leo11, Mal12, Pau12, Sci16, Swa11, vdT14]. Lifeboat [OM16]. Light [PB14]. Like [Fas18]. Limit [All17]. Limits [CF06, Jac01, Luk93]. Lingua [Wol15]. Linguistic [Alb06, Bjö12, Pet23, Pol20]. Linguistics [Gon12]. Living [Dav99, Saa07]. Local [Ruh13]. Locke [Nar01]. Logic [Kaz14, Sha07, Kuk95]. logical [Hei01, Rei01]. Lola [Okr98]. London [Hue25]. Lonely [HS09]. Long [Ken11]. Longish [Bur19]. Look [Bru13, Vic13]. Looks [Orr24]. Looping [Tso07]. Loops [CF06]. Loose [Let19]. Lots [Kha10]. Lucas [Har97a]. Lure [Bri17].

M [Bra94, Pla10]. MA [Bus19]. Machines [Ste06]. macrocosm [Gou87]. Macroscopic [Bur19, VG15]. Made [Bus11, Tod08]. Magnetic [Col22, Fey22]. Major [Tra20]. Make [Per13b]. Make-believe [Per13b]. Makes [Koc09]. making [Bra94, Nov10, Won97]. Mallards [God14]. Mamchur [Par24]. Mandelstam [Pec02]. manifold [Hug99]. Manipulationism [Kow17]. Manipulative [Gel03]. Many [Baç10, Ken11, Les96, Vai98]. many-worlds [Les96, Vai98]. March [Gar94]. Maria [Bus19]. Marriage [Bus11]. Marxism [She07]. masculine [Til87b]. Massimi [Bic22]. Massive [Vil23]. masterpiece [Bro90]. Material [Bri10]. materialism [Jac93]. Mathematical [Ban09, Boo11, Bus11, Dan20, Duc05, Mar99, Pap88a, Van10, Wag09, Wag13, Why25, Det89, Sza03]. Mathematics [JM16, Ken13, Rus13, Tho92, Tro11, Why25]. Mathematisation [Kva16]. Mathesis [Til87b]. matter [Car87]. Matters [Yan08, Ear02, Psi00]. Maupertuis [Feh88]. Maximum [Rec17]. Maxwell [All13]. May [Tod08]. McAllister [Dav98]. Meaning [CMF14, Laz17, Mel23, Swa11, BS95, Cor92, Mil91b, Mon93b, vBG88]. Means [ES07, SC13]. measured [Cha97]. Measurement [KF22]. measurements [Bok03]. Mechanical [Käs19, DF94]. Mechanically [Per12]. Mechanicism [MHB22]. Mechanics [All13, All17, Bar20, CGL05, Del22, Duw16, Fel14, Ken22, Kut10, Lar06, Laz17, MA13, Ori24, Pan17, Sko10, Uga15, vS24, Cha97, For00, Fri03, Hod97, Hug99, Lew03, Luc95, Nor01, Pec02, Rob92, Rös89, Tem91]. Mechanism [Kuo09, Lyr08, Wei24]. Mechanisms [Mac04, Was08]. Mechanistic [Bae13, Har10, HGS24, Ill11, Web09]. Medicine [And20a, Tei13, Wil19, Qiu89]. Meets [How18]. Memoriam [Par24]. memories [Nor00]. Memory [And15]. men [Wil91]. Mental [Mis92, Kis99, New94]. Mentzou [Bro94]. Mereological [Pin06]. Messaging [Why25]. Meta [Mül15, Ruh13, Sho18]. Meta-Induction [Sho18, Mül15]. Meta-inductions [Ruh13]. Metacriterion [Fas18]. Metaphor [OM16, Feh88, vBG88]. Metaphors [Wag13]. Metaphysical [Ben17, Gra18, Mar05, Rup05, Bis03, Gro91]. Metaphysics [Ala24, Bur19, Den15, Liv21, Mac04, Max09, Mei14, Vet11, Pap88b]. Meteorology [New01]. Method [Del22, Gha17, Ruz11, Thé15, Vis14, Feh87, For93]. Methodological [JB15, Mat21, Res94, Car87, San96]. Methodology [Bol12, Bac92, BO91, For93, Lar94, Ram93]. mice [Wil91]. Michael [Pau12]. Michel [Ala24]. Microbiology [Kai15]. microcosm [Gou87]. microscope [Bai98]. Middle [Bur19]. Middle-sized [Bur19]. Mill [Nar01]. Mimesis [Ban12]. mimetic [Rue02]. Mimicking [VG15]. Mind [Fra09, HS09, Phi14, Sle09, Fla03]. Mingled [Chi10a]. minimal [Ada04]. Minkowski [Kut10]. Minkowskian [Gau05]. Miracles [Hun11, Mat98]. Mirage [Ses11]. mixed [Pap90]. Mizrahi [CT19, Isl17]. Modal [Esf09, Rea20, Rue03, Bri99]. Model [Bod06, Bor13, Che23a, Chi19, Kou15, Moo12, New12, New13, Par15, RH05, TP23, Ear08, Gir94, JL94a, JL94b, Lin95, Smi94, dCF93]. Model-based [Chi19, Kou15]. Model-Theoretic [RH05]. Modeling [Pin11a]. Modelling [Sve14, Tam25]. Models [BK14, Bul09, CL19, Cas17, GY11, HGS24, Hei06, Hoc13, JK15, Knu10, Liu97, Liu98, Liu16, Paa06, Per13b, Qiu89, Roe17, Ste06, Tug13, Was08, BJ03, Bon97, Duc05, Gar94, Har88, Kra97, Mis92, New94, Nug88, dS91]. Modern [Abr16, Jac11, Max09, Pea10, Sar16, Sch12, Sve14, Bai98, MR88, Nug88, Rad02, Rös89, Sl91]. Modernity [SS15, Rad02]. Modes [Ale09]. modest [Ci03]. Modularity [Vil23]. Modus [How09, Lyo13]. Molecular [Bae10]. Molyneux [Par20]. Monad [Mor11]. Monopoles [Col22]. monster [Hon91]. Monte [BN12]. most [Eng02]. motion [Mat87]. motivation [McF91]. move [Gro91]. Movement [De 15b]. MR0599360 [Duc05]. MR2106565 [Mam10]. MR2201925 [Kra06]. MR2253243 [Flo06]. MR2680848 [How11]. MR2845407 [Ste11]. Multidisciplinarity [Alv11].

Multiple [Bru13, Ful22, Kis99, SC13, Str17]. Multisemiosis [MS09]. Multiverse [Fri17]. music [Gou87]. Myth [Miz16].

N [Gat00]. Nagel [Che23a]. Nancy

[Okr98, Orr24]. Nath [Das12]. nationalism [Mis97]. Natural [Bai89, Bir12, BV14, Bou19, Bru13, Brz16, Des18, Far19, God14, Hul14, Nee12, Oft10, Wei24, dP15, Bon97, Kut87, Liu03, McA97a, Pag02, Pih96, RR95, dCF93]. Naturalism [Ban09, Fre06, WĆ06, Ye10, Kno02, San96, Sed96, Sla00]. naturalist [Lel89]. Naturalistic [BB21, Bor12, Kow14]. Naturalized [Koc12, Sce98]. Nature [Ala24, Bri17, Bus19, Esf09, Jae17, Jan15, Kva16, Ses11, Vet11, Whi15, Bak91, DF94, Gol93, McA97a, Rad02, SM94, Phi14]. Natures [Mel23]. Navier [PZ14]. Navigating [Bic22]. necessity [Hun01, Mac01]. need [Jan89]. Needham [Bir12, Sce99, Sce00]. Negative [BB21]. neglected [Wyl00]. Neo [Dim16, Reu13]. Neo-Kantianism [Dim16]. Neo-Russellians [Reu13]. Neoplatonism [Mar10b]. Nesić [Ben22]. net [Fre97a, Fre97b, Fre98a, Fre98b, Fre99]. Netherlands [McA97b]. network [Mis01]. Networks [Pol20, YH24]. Neural [YH24]. Neurath [Okr98, Rei01, Ueb01]. Neurobiological [Har10, Yan13]. Neurophilosophy [Sin17]. Neuroscience [Abr16, Ven21]. Neutral [Nes22, Obe90]. Neutral-Structuralist [Nes22]. neutron [Vai98]. Newcomb [Irv93, Mar96]. news [Ano97e]. Newton [Bel07, BM22, Car87, DiS92, Duc05, Feh87, Gou87, Kva16]. Newton-Smith [BM22]. Newtonian [Duc05, Die01, Duc05]. nexus [Goo90]. Niels [How18, Rio17, Bor13]. Nineteenth [Zem13, PZ14]. nineteenth-century [PZ14]. No [CE11, HS09, How09, How11, Hun11, Kas04, Flo06, Kra06, Mat98, Mil91a]. no-miracles [Mat98]. Nominal [Ray09].

Nominalism [Saa07]. Nomological [All17]. Non [BTB16, Bau09, Gur22, How16, Lyo09, Moo12, Pin10, Reu11, SS15, Sve14, YH24, Ant97, Fut02, Lar01, Réd92]. Non-Bayesian [BTB16, How16]. non-collision [Lar01]. Non-committal [Pin10]. non-commutative [Réd92]. Non-competitor [Lyo09]. Non-epistemic [Sve14]. Non-European [SS15]. Non-Factivist [Gur22]. non-observational [Ant97]. Non-reductive [Bau09, Moo12]. non-tensed [Fut02]. Non-Uniqueness [YH24]. Non-universal [Reu11]. Nonlocality [Woo07, Pea98]. Nonrelativistic [All17]. Normal [May23]. Normative [Fre06, San96, Sla00, Ste10, Jan87, Kno02, Til87a]. Normativity [BS22]. Northpoles [Fey22]. notebook [Mar01]. Notes [Ano87c, Ano87d, Ano88d, Ano88e, Ano89c, Ano89d, Ano90d, Ano90e, Ano90f, Ano91e, Ano91f, Ano91g, Ano92e, Ano92f, Ano93d, Ano93e, Ano94d, Ano94e, Ano95d, Ano97e]. Nothing [Bac11]. Notice [Ano01c, Bai98, Ano98d, Dav98]. notices [Ano97f]. Notion [Mak23, Mam16, Per13a]. Novelty [Jeb17]. Nozick [Wyl00]. Number [Sin17]. numbers [Šik96]. Nurture [Ses11, Whi15].

O [Nee02]. obeyed [Nii90]. Obituary [Ano97f, Ano98d, Ano03g, Big07, Bub03, Gat00]. objections [Myr03]. Objective [BB10, DS12]. Objectivity [Amb10, Fil16, Kuu12, Pea10, Pop16, Tul16, Wyl00]. Objects [Bur19, Haw10, Kra92, Kra97, Sza03]. Observable [She21]. Observable-Unobservable [She21]. Observation [Per21, Goo90]. observational [Ant97]. observations [Fje91]. Obsolete [Mül15]. Oct [Bus19]. Oil [VG15]. Okasha [Jon00]. Old [Fri17, Miz15, Nug96]. Oliver [Bol87].

Omniscienter [Kos02]. Once [Wil95]. One [Baç10, Ken11]. Only [FHS21, Nor00]. Ontic [Esf09]. Ontological [CACA19, Vil23, Har97c]. ontologically [Pih96]. Ontologies [Leo12, Rob95]. Ontology [Ben17, Chi10b, Laz17, Rus13, Slo05, Bel95, Har90, Har95, Pea02]. Open [Leo12, Reu13, Why25]. Opposite [Bog04]. Optics [Zem13]. Optimal [HGS24]. Optimality [Sho18]. option [Wyl00]. Order [De 15b, Jae17, Pin11b]. Organicism [Mal07, vdB14, Mar05]. oriented [CMF14, CMF16, Gao17]. Origin [Alc10, Rad02]. Origins [Ano02a, Hei01, Ort24]. Ørsted [Phi14]. Orthodoxy [dR16]. Osman [Nie08]. other [BS95, Mos94, Smi94]. Otto [Rei01, Okr98]. Outcome [BB21]. outsider [Das12]. Overall [Kra21]. Overlooked [And20b]. Overpopulation [OM16]. Overview [All10]. Oxford [Bor13, Duw16, Pla10].

P [Cla13, Pla10]. page [Ano97d]. Paleontology [Bry12]. paper [Aga93, BR93, Kli93, Lin93, Mar01, Mon93a]. paperback [Ano02a, Dud17, Orr24, Rad02, Zha15]. Paraconsistenciti [SS14]. paradigm [Bab03, Pea98]. Paradigms [Asa13, Hun01]. Paradox [CL11, Ped18, Wil23, Irv93, Mar96, Woo07]. Paradoxes [Kas04, Mas09]. Paribus} [Kow14, Kow17]. Parochialism [Wol15]. Part [SS24]. particulars [Har95]. Past [Nee13, Nor00]. path [Tha02]. Paths [Bic22]. Patrick [Tha02]. Pattern [Kou23]. Patterns [Ben17, Feh98b]. Paul [Far00, Ken22, Sce00]. peaceful [Pea98]. Peirce [PB14]. People [Sch22]. perceive [Mac00]. perception [Hac95]. perceptualist [Sed96]. Perfect [Bou19, Kli93]. Perils [Bag17]. peripheral [Das12]. Persistence [Kat13]. Perspectival [Bic22]. Perspective [Mas09, BO91, Pec02].

Perspectives [Ort24, Pea10, Rio17, Tul16, Hun01]. Perspectivist [Ruy21]. Pertinence [Mal12]. Pessimistic [Mül15, Ruh13, Wra15, Bis03]. Peter [SS14]. Phenomena [Bou24, Ste06]. Phenomenological [Bar20]. Phenomenology [The12]. Phenomenon [Wei24]. Philos [Flo06, Kra06]. Philosopher [All10, De 15b, Miz16, Orr24, PG06]. Philosophers [Pay95, Thy22, vS24, Feh90]. Philosophical [Alc10, Auf12, Bry12, Cas17, Cha11a, Ken22, Pau16, Pin11a, Sim16]. Philosophically [Wag13]. Philosophy [Alm07, And20a, Ano02a, Ano15, Ano17, Ara17, Ara20, Bab10, Bol12, Bre15, Cam05, Cha11b, Chi10a, Col20, Der10, Ekw15, Far19, Gin05, Gon15, Hol09, Irv97, Kai15, Kap10, Käs19, Ken11, Kno11, Knu10, Koc12, Lew16, Mar15, McA97b, McA08a, Nav07, Nel20, Okr98, Par20, Pio14, Pla10, Pol20, PS23, Rad15, Rin12, Rio17, Sch11, Sch09, Sei16, Sha16, SS22, Sle09, SS24, Tei13, The12, Val12, Van10, Vaz16, Wil11, Wol15, YTH23, Yan08, vdB14, Bai98, Car98b, Cut88, Far00, Kle01, Man91, Mis01, Nar01, Nie08, PS01, Sch98, vdS14]. Physical [Thé15, De 02, Gol93, Har90, Hun01, Nor00, Sza03, Why90]. Physicalism [Bau09, Pao17, Pin06, Jac93]. Physicalism* [FHS21]. physicalist [Sza03]. Physics [Gra18, Lan23, Mam16, Ort24, Pao17, Pea10, Per14, Rio17, Sar16, SS22, SS24, Vic06, Way15, Die01, Eng02, HT00, Jan87, MR88, Nug91, Nug96, Rös89, Spu99, Kno11]. picking [Miz15]. pictorial [Sed96]. Picture [Ler01]. pions [Hoe00]. Place [Das17, Mam16]. Planck [Cha97]. Planetary [Car10a, Mar92]. Planets [God14]. Platonic [Bro89]. Platonism [Bro94, Len05]. Platonist [De 15b]. please [Far00]. Pluralism [Cla97]. Pluralist [Bus19]. Pluralistic [Far00]. Poincaré [dP15]. Policy [CMF14, CMF16, Vaz16, McA01]. Policy-oriented [CMF14, CMF16]. Political [Bak91]. Politics [Gin05, Okr98]. Polity [Dud17, Zha15]. Pollock [Jac93]. pomegranites [Mar92]. Ponens} [How09, Lyo13]. Popper [Auf12, Bla95, Buz10, Del22, Ort24, PV20, Par90, PG06, Sce98, Sep25]. Popperian [Far00, Sla00]. Populations [Oft10]. Possibility [Lam12, Hoh03]. Possible [Bri17]. Post [Bus19, Pol20, Sle09]. Post-Essentialist [Bus19]. Post-‘Linguistic [Pol20]. Post-Reductionist [Sle09]. Postcolonial [SS15]. Posteriori [JM16]. Postmodern [Gin05]. Postmodernism [Bor12]. Postmodernist [Kid16]. Power [Bul09]. powerful [Har95]. Powers [Ala24]. pp [Ala24, Ano02a, Bor13, Bus19, Dud17, Duw16, Hue25, Orr24, Pau12, Pla10, Rad02, Zha15]. Practical [Ruz11]. Practice [Brz16, Cha11a, Gol20, Lew16, Mar10a, Miz20, OM17, Per21, Sch09, Van10, Wyl88]. Practices [Amb14, Boo11, Lin22, dMM23, Mar05]. Pragmatic [Ale09, BB21, vBG88]. Pragmatic-Naturalistic [BB21]. Pragmatism [Alm07, MA13, Mis90]. Praise [Far19]. Precautionary [Vaz16]. predicates [Lin95]. Prediction [Lip90, Gon95]. Preferential [RH05]. pregnancy [Won97]. prejudice [Lip90]. Premises [Tra20]. preparation [Rue03]. Present [Mam16, Cra00]. Present-day [Mam16]. Press [Ano02a, Bor13, Bus19, Duc05, Dud17, Duw16, Orr24, Pau12, Pla10, Zha15]. Presupposed [Gir23]. presuppositions [Mar05]. Prevail [Rup05]. Primacy [How18]. Primate [Alc10]. Primates [Duw16]. Primer [Cas17]. Principia [Duc05]. Principle [Dar07, PV20, Ste10, Vaz16, BS95, Feh88, Rei01, Dou00, Ear04].

Principles [Max09, San96]. priori [Art99, Fje91, Iva11, McA08c]. Privilege [Kha10]. Probabilistic [Web09, Got00]. probabilities [Pap90]. Probability [Lan14, Hei01, Log97, OM94]. probably [Hug99]. Problem [Bae25, BB21, BS22, EB14, Fas18, Fer11, Gil92, Sho18, Ste10, Ste11, TP23, YH24, Acu16, Hes92, Irv93, Jac01, Mar96, NS92, Nor01, Pag02, Rat05, Rue03, Sim92a, Sim92b, Tra92, Vah94, Vor89]. Problem-Feeding [TP23]. Problems [Tra92, Hun01]. Processes [Bur19, Kro08]. Produced [Per12]. Producing [Pay15]. Production [SS15]. Professors [Lan11]. Program [Dan20]. Programme [Koc08, Feh98a, Oka00b, Sle91, Tem91]. Programmes [Pay15, For93]. Progress [CT19, CW13, Gao17, Isl17, Mar10a, Miz17, Mül15, Row10]. Project [Aga92]. Projectability [Hun90]. Projectivism [Pop16]. Prometheanism [Bak91]. Proof [Bro90]. Proofs [Car10b, For93]. Proper [CW09, Kou23, BV24]. Properties [Lee14, Vic13, DF94, Kis99, Mel90]. property [Fri03]. Propositions [Sal93]. Prospects [Luc95, Ear08]. prothrombin [Mar01]. Provisional [EB14]. Pseudo [How16]. Pseudo-hypotheses [How16]. Pseudoscience [Bab15, BB21, Let22]. Pseudosciences [Fas18]. Psychiatric [Tso07]. Psychic [Gom91]. psycho [Why90]. psycho-physical [Why90]. psychological [Car98a]. Psychology [Käs15, Val12, Rat05, Sea90, Wil91]. Psychopaths [BJS17]. Ptolemaic [Car10a]. Ptolemy [Rec17]. Public [Ci06, Feh90]. Publishers [Rad02]. Pure [Yag17]. purpose [Pri00]. Pursuit [Sha20]. Push [VG15]. Putnam [Nor01]. Puzzle [Ban09, P´ 18]. Puzzles [Kas04].

Quanta [Bac11]. Quantities [For00]. Quantum [All10, All17, Bar20, Bog20, Bok03, Bor13, CGL05, CS24, Cru24, Del22, Duw16, EV14, Eng01, Fel14, Kas04, Ken22, KF22, Kut10, Lam13, Laz17, MA13, Ori24, Ort24, Pan17, Pea98, VG15, vS24, Acu16, DF94, For96, For00, Fri03, Hod97, Lew03, Luc95, Nor01, Pea02, Pec02, Rob92, Rös89, Vai98]. Question [Kar15, Par20]. Questioning [Nee13]. Questionnaires [CACA19]. Questions [Kra92, BV24, SH92]. Quine [Cha11b, Wei95].

R [Har97a]. Race [Kou23]. Racist [Mnc17]. Radical [Rob95, She21]. Ramsey [Cru04]. Random [Tem91]. Randomised [Kra21]. rapid [Tem96]. Rational [Kuu17, P´ 18, Til99, Fuc93]. Rationalism [Auf12]. Rationality [Aga93, Ede10, LR93, Ori24, PV20, Van93, Wor90, Wra21, vS24, BV24, Bou93b, Bro99b, Cre93, D’A90, Dav99, Fol87, Kli93, Lin93, Log97, Mar87, Mon93b, Sal93, Zhe88]. Ravens [Ped18, Wil95]. Raymond [Aga93, BR93, Fuc93, Kli93]. Re [Liu15]. Re-inflating [Liu15]. Reactivity [JB15]. Reading [Phi14]. Real [Chi10b, Kro08, BV24]. Real-World [Kro08]. Realisability [Ful22]. Realisation [Bru13]. Realism [Ala24, Asa12, Asa13, Ben17, Ben20, Bog20, Bus11, Che23b, Cru24, Eme17, Esf09, Har95, Hod97, Hun11, Iva11, Koc08, Lyo09, Lyo13, Mat21, McA08c, Miz13, Pan17, Pil23, Roh04, Saa07, Slo05, Cha04, Ci03, Cru04, Kle01, Kra92, Lel89, Luc95, McA89, Nor01, Psi00, Tuc90, Via99, Wra08, Wra21]. Realism/Anti [Wra21]. Realism/Anti-realism [Wra21]. Realist [Pin24, Fri03]. realities [Til90b]. Reality [Luc97, Roc10, Sar16, Tro11, Mam90, Sea99]. Realization [Str17, Kis99]. really [Har88, Kno02]. Reason [Gha17, Roc10, Mis96]. reasonableness [Van93]. Reasoning [Boo11, Bri10, Kou15, PB14, Sci16, Gir94, New94, dCF93]. Reasons [Bie16, Alb93, BV24, Bou93a, Cre93, Dav99, LR93, Mac00, Nad93, Piz93]. Reassessing [Ear08]. Reassessment [KF22]. rebus [SM94]. Receptive [Ven21]. Recognise [Car23b]. Reconciled [dR06, Tuc90]. Reconciliation [Gur22, PV20]. reconfiguration [Elg02]. Reconsideration [Kin17]. reconsidered [Wei16]. Reconsidering [But07, Col22]. Reconstructing [Auf17]. Reconstruction [Bor12, Kuu17, Mat87]. recourse [McA89]. Red [Chi10b]. Reduction [Che23a, ES07, KBdR05, Nee99, Nee00, RH05, Sle09, Kis99, Ram93, Sce98, Sla00]. Reductionism [CF06, Kha05, Mal07, Mar05, Rup05]. Reductionist [Sle09]. reductive [Bau09, Moo12]. Referees [Ano24]. Reference [Asa12, Bis03, Cru04, Mil91a, New01, Nor01, Psi97]. Referential [Bae10]. Referring [Bir12]. Refining [New13]. Reflections [Alc10, Har91, Ara17]. Reflective [Bac92]. Reforms [And20b]. refutations [For93]. Regularity [Bae13]. Regulation [BS22]. Rehabilitating [Bel99]. Reichenbach [Auf17, Bog20, Gon95, Kle01]. Reinflating [FL99]. Rejoinder [CM95]. Relating [Smi94]. Relation [Kin17, Pan17]. Relationalism [Rea20, Bel99]. relationism [Bri99, Die01]. Relations [CG12, CW09, Way15, Mar90]. Relative [Miz13]. Relativised [Iva11, McA08c]. Relativism [Zib94, San96, Tuc90, Fre06, Sci16]. Relativistic [But07, EV14, Die01, Jan89]. Relativity [Dor07, Fer11, Gha17, Kat13, BS95, Bro97, Moz00]. relevance [Wyl88]. Reliabilism [Koc08]. Reliability [CL19, Sol14, Yan13, Ada04]. Religion [Alc10, Run11]. religious [Mac92]. Remapping [Hue25]. Remarkable [Boo16]. remarks [Alb93, Car87]. reovirus [Tha02].

Replicability [And20b]. Reply [Ben22, Bir12, Bou93a, CT19, How09, How16, Isl17, JL94b, Koc09, Oka00a, Sim92b, Ste11, Wra21, Bro94, Flo06, Fuc93, Kra06]. Report [Ano05e, Ano06c, Kin18, Kin20, McA07, McA08b, McA09, McA10, McA11, McA12, McA13, McA14, McA15, McA16, McA17, Ano02c, Ano03f]. Represent [Mey08, BJ03]. Representation [Ban12, Fra09, Gre94, Ken13, Liu15, Mas09, Mei14, Per12, Per13b, Too12, YKMM22, Cor92, LM95, Sed96, Sua03, dS91]. representational [Hun01]. Representations [Amb14]. Representative [Amb14]. Reproducibility [Bae13]. Rereading [Kuu09]. res} [SM94]. Research [CACA19, Car23a, Car23b, CMF14, CMF16, Gin12, Jac11, Kou23, Mal07, Pet23, Ruz11, TP23, Wil23, For93, Mar05, Smi94, Tem91]. Resilience [Tho14]. Resiliency [Log97]. Resisting [Mis97]. resolved [Ueb01]. Respond [Car23b]. Response [All15, Sce99, Flo06, Jon00, Nee99, Nee00, Sce00]. Responses [Kin11]. Restoration [Lyo13]. Resultant [Moo12]. Results [Sol14]. Rethinking [Buz10, Jae17, Rad10, Wyl00]. Retroduction [PB14]. reversal [Ear02]. Review [Ala24, Ano02a, Bor13, Bus19, Ci03, Cla13, Col20, Dud17, Duw16, Hue25, Kno11, Lel88, Okr98, Orr24, Pau12, Pla10, Rad02, Zha15, vdS14, Abe91, Ili91, Mam10, Man90]. Reviewers [Ano21, Ano22, Ano23]. Reviewing [RH05]. Reviews [AKLV98, Ano00a, Ano00b, Ano00c, Ano01a, Ano01b, Ano02b, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano05b, Ano05c, Ano06a, Ano06b, BBD09, CHLL07, CBG06, CJD^{+}08, DLVW08, JSJ^{+}94, KSA07, SM91, VHO^{+}07, YTP08, Ano05a, BSU97, CIL96, HF92, HMO^{+}99, JBM^{+}95, LB98, LPR99, LLOA93, MV90, MWRB98, OKHT96, PLO94, PS92, STL^{+}96, WR93, WMB97]. revision [Jan89]. Revisited [Alb06, YKMM22, Dav99, Liu97, Wei95]. Revisiting [Kva16, Pan17]. Revolution [De 15b, Far19, NL16, And91, BV24, Bro87, Lin95, Duc05]. Revolutionary [May23, PG06]. Revolutions [Kin11]. Richard [Zha15]. Riemannian [Gau05]. Rightful [Das17]. Rigid [Ina14]. rise [Bak91, Mis97]. Rises [Auf17]. Risky [DS21]. Road [All13]. Rob [Bub03]. Robert [Mac92]. robots [McF92]. Robustness [Sol14, Log97]. Role [DL23, JK15, dP15, Bro97, Das12, Feh88, Feh90, Fje91, Kuj98, Lep88, Mam87, Til87a, Wyl88]. Roles [Ci11]. Romanticising [Hue25]. Room [Gur22]. Rorty [May23]. Routledge [Hue25]. Rule [Isl17, Koc09]. Rupik [Hue25]. Russellians [Reu13]. Russo [Ill11, Tre21].

S [Pla10]. Salmon [dR06]. Same [Miz15]. Samir [Jon00]. Satyendra [Das12]. Saul [How18]. Scale [Bae25]. Scerri [Nee99, Nee00]. Schelling [Hue25]. schemes [Bab03]. schizophrenic [Vai98]. Schlick [Ort24]. Schrödinger [Mal12]. Sci [Flo06, Kra06]. Science [Ale09, Alm07, Ano15, Ara17, Ara20, Bab10, Bab15, Ban12, Bod06, Bol12, Boo06, Bre15, CACA19, Cas17, CS24, Chi19, Ci11, Col20, Das17, Ekw15, Fey22, Gin05, Gon15, Han20, Hib16, Hol09, Hop13, Iva14, Jac11, JLL92, Kap10, Kar15, Kha05, KBdR05, Knu10, Koc12, Kou15, Kuu12, Leo12, Let22, Lew16, Liv21, Mar08, Mar12, Max09, McA08a, Mis01, Miz16, Mnc17, Mor06, Nav07, Nee13, NL16, Okr98, Pap88b, Pau12, Pay15, Pet23, Pin11a, Qin16, Rad10, Rad15, Run11, Rus13, RS10, Sch09, Sch12, Sch17, Sha16, She07, Sim19, Sle09, Sve14, The12, Tod08, Tul16, Vaz16, Vet11, Wil11, Wol15, Yag17, YTH23, YH24, dP15, vS24, Art99, Bac92, CM90, Car98b, Ci03, Cut88, D’A90, Das12, Dav96, Dav98, De 02, Elg02]. science [Feh90, Irv97, Jac01, Kra97, Law93, Len01, Lin93, Mam87, Man91, McA97b, McA02, Mil91a, Mil91b, Nar01, PS01, Rad02, Sep25, Sl91, Til87a, BBB06, Ano02a, Orr24, Rad02, Ano17, Rad02]. Science/Knowledge [Pay15]. Sciences [Alv11, CG12, CW09, Gon15, Hul14, RW07, vdT14, Bon97, Kut87, Tho92]. Scientific [Alc10, And15, And20b, Asa13, DL23, Ber96, BB21, Boo11, Bri10, Brz16, Bus11, Cha11a, Cha11b, Che23b, Ci06, CW09, Cru04, De 92, De 15b, Dim16, Dur17, Eme17, Gao17, Gij09, Gin12, God14, Gur22, Hoc13, Isl17, Iva14, Jan15, Ken13, Kid13, Kin11, Knu10, Koc08, Kut87, Lin22, Liu15, Liu16, Lyo09, Lyo13, Mäk13, Mar08, Mar10a, Mas09, Miz17, Miz20, Mül15, New12, New13, NL16, Oka00a, Pan17, Per13b, Per21, Pet23, PB14, Pil23, Roe17, Row10, RS10, Sch17, Sea87, Sim16, Sim92a, Sim92b, Sol14, Sua03, Til88, Too12, Vis14, BJ03, Boo04, Bro87, Cha04, Fje91, Fol87, For93, Hun01, Kos02, Ler01, Lin95, Los92, Mac92, Mar01, Mar05, Mar87, Pet92, Pet93a, Psi00, Ram93, SH92, Zhe88, Gil92, Ala24, Hes92, NS92, Tra92]. Scientism [dR16]. Scientometric [YTH23]. Scourge [Joy13]. Seagulls [Fas18]. Search [Iva11, Laz17, Mel23, Sha07, JLL92]. Searching [YH24]. Second [Pin11b, Sce00, Nee00]. Second-Order [Pin11b]. Selecting [Dar07]. Selection [BV14, Bou19, Bru13, Oft10, Wei24, Bai89, RR95]. Selectionist [Gon12, She21]. Selective [Ped18]. Selfhood [Nes22]. Selves [Ben22]. Semantic [Isl17, FL99, Ler01, Liu97]. semantics [Rue03]. Sennert [New01]. Sense [Bou19, Sha20]. Senses [Chi10a, Kut87]. sentences [Cru04]. Separation [Bae25]. SETIbacks [Why25]. sets [Mac00]. Sfendoni [Bro94]. Sfendoni-Mentzou [Bro94]. Shaped [PG06]. Shaping [Bou24]. Shared [Roe17]. Shech [Wra21]. shift [Pea98]. Should [Mis96, Nii90, Rup05, Vai98]. Shutters [Kas04]. Similarity [Too12, Sua03]. Simon [Aga92, Gil92, Hes92, JLL92, Los92, Mar92, NS92, Tra92]. Simple [Wyl88]. Simplicity [HGS24]. Simulation [Jeb17, Par08]. Simulations [BK14, Dur17, Kro08, BN12]. Simultaneity [Buz14]. since [Bro87]. Single [Fey22]. singularities [Lar01]. sized [Bur19]. Skeleton [Mag08]. sky [Law93]. slain [Hon91]. Smith [BM22]. so-called [Fri03]. Sober [Bog20]. Social [Bou24, Bul09, CACA19, CS24, Ci11, Gon15, Hol09, Hul14, Lew16, Lin22, Ren13, Sch17, Sha20, Tis89, Gil91, Lin93, Res94]. Society [SS15]. sociobiology [dS90]. Sociology [Koc08, Sch17, Oka00b, Pet93a, Wor90, dS90]. Sokal [Lon97]. Solipsism [Bog20]. Solution [EB14, Fas18]. Solved [Sho18]. solves [Irv93, Mar96]. solving [Jac01, Sim92a, Sim92b, Gil92, Hes92, NS92, Tra92]. Some [Alb93, Car87, Cul95, Har91, JLL92, Let19, Mak23, Myr03]. Sorites [How09]. Source [JM16]. sources [Har90]. Southpoles [Fey22]. Soviet [Bak91]. Space [Bel07, Die01, Gau05, Kut10, Ses11, Wei13, Dav99, Har97c, Hoe00, Til88]. Spaces [SS15, ZG16, Hun01]. Spaceships [Fer11]. Spacetime [Slo05, YKMM22, DiS92]. Span [Dur17]. spatial [PZ14]. speaking [Car98b]. Special [Car23a, Fer11, Bro97, Bro99b, Moz00, New01, Wil92]. Specialisation [DL23, Pol17]. Specialised [Roe17]. speciation [Vor89]. Species [Wal13, Vor89, dS89]. Sphere [Ci06]. spontaneous [Ear04, Lew03]. Springer [Ala24, Chi19]. Srdan [Rad02, Ano03g]. Stance [Pin24, Ruy21, RR95]. Standardisation [YTH23]. State [PG06, For96, Rue03]. statement [McA01]. Statements [BG12]. Statistical [All13, Rat89, Réd92]. statistics [Das12]. Status [Hun11]. Stellar [Suá13]. step [Far00]. Stewart [Nar01]. Stimulating [The12]. Stokes [PZ14]. Stone [Abr16]. Story [Boo16]. straighten [Ant97].

Strategies [Way15, Lew03, Sol14]. Strategies* [Pil23]. strategy [Gre94]. String [Vis14]. Strong [BB10, Koc08, Feh98a, Jac93, Oka00b, Sle91]. Structural [Ben17, Ben20, Dor07, Esf09, Iva11, McA08c, Slo05]. Structuralism [Cha04]. Structuralist [Nes22]. Structure [Bor13, Haw10, Kre17, Lam13, Lan14, Mei14, PS23, ZG16, Car87, Feh87, Fri03, Gin12, Nug88]. Structures [Esf09, Tug13, Bus03]. Struggle [Whi15]. Stud [Flo06, Kra06]. Studies [Ano02a, Ano15, Ano17, Bol12, Koc12, Lew16, She07, Tul16]. Study [Cru24, Miz20, Pie12, Pol20, Ruz11, Sik92]. Studying [Sch09]. Stuff [Wei13]. Style [Duc05]. Styles [Boo11, Sci16]. Sub [ES07]. Sub-types [ES07]. subjunctives [Rat89]. Substance [Mor11, Lod01]. substantival [Hoe00]. Substantivalism [But88, Goł99, Hug99]. Success [Iva14, Gel03]. supertasks [Bok03]. Supervenience [EB14, Jae02]. Support [Sol14]. Supposed [Tod08]. Surgery [Abr16]. Surrender [Tym21]. Survey [Alm07]. Sustainable [Jod24]. Sweep [She07]. Switzerland [Ala24]. Symbolic [Til88]. symbols [Tra92]. symmetric [Wei16]. Symmetries [Fil16, Ruy21, BS95, Kos00]. Symmetry [Kas04, Sko10, Che02, Ear04, Wor90]. Symposium [BBB06]. Synthese [Ala24]. Synthesis [Mam10, Feh87]. synthetic [Bou93b, Cre93, Sal93]. System [Car10a, Kuo09, Rus10, Jae02]. Systemic [Mal07]. Systems [Lam13, Pin11b, Reu13, Rou11, Mar92, Sza03].

T [Zha15]. Tainted [Sha16]. Taper

[How16]. Taxonomic [San98]. Taxonomies [Mel23, Pol20]. Taxonomy [Tra20]. Taymiyya [Ray09]. team [McA01]. Technological [Boo04, Til90b, Nii90]. Technologies [Nel20]. Technology [Bol12, Boo06, Kou15, Nel20, Obe90, Qin16, Tul16, Til90a]. teenage [Won97]. Teleology [CW13, Tet89]. tell [Nug91]. Temporal [Pal16]. tense [Moz00]. tensed [Fut02]. Terms [Max09, Cru04, Dou00]. Terra [KBdR05]. Terrestrial [Why25]. Test [Row08, Rus14]. Testing [DS21, Miz20]. Texture [Why25]. Their [CL19, Fey22, Fra09, dP15, Car87, Ueb01]. Them [CE11, dMM23]. themes [Far00]. Theologians [Kap10]. Theological [Alc10]. Theorem [Ant12]. Theoretic [RH05, Tug13, dCF93]. Theoretical [Dou00, Ina14, Kou15, Cru04, Flo99]. Theories [And15, Che23b, Dar07, Fra09, MHB22, Mar90, Max09, Miz16, The14, Eng01, Eng02, Eng05, Flo99, Fuc93, Ler01, Lew03, Liu97, Liu98, Rat89, Til88]. Theory [All10, Bol12, Buz14, Cha11b, Che23a, Dor07, EV14, Ede10, ES07, Fel14, Gha12, Goo90, How16, Hun11, Kas04, KF22, Leo12, MHB22, McA08c, Mel23, Nes22, OM17, Pin11b, Rad02, Reu11, Sha20, She21, Ste10, Svo15, Thé15, Vis14, ZG16, Acu16, Ada04, Bon97, Bou93b, BO91, Coh94, Cre93, Fje91, Flo99, Fut02, Gir94, Hac95, Hei01, Hun94, JLL92, JL94a, JL94b, Jon00, Lep88, Les96, Mar90, Mat87, Mil91a, Myr03, Nug88, Oka00b, Pag02, Sal93, Smi94, Tis89, Vai98, Zib94, Rec17]. theory-acceptance [Jon00]. theory-dependence [Ada04]. theory-ladenness [Fje91]. There [BG12, Bou19, Bre15, CE11, Che23b, Gur22, Kuk95, Mam90]. Thesis [Ill11, Kva16, Tre21, dR06, Wei95, Wor90]. Thingamajigs [Bir12, Nee12]. Think [Lan11]. Thinking [Jan15]. Third [Par15, Sia19, Rei01, Wyl00]. Thomas [Okr98, Ghi03, PS23]. Thomson [Cla13]. Thought [Bro87, Far19, Pau12, Art99, Bon97, Kuj98, Mis92, Mos94]. Threat [How09]. Three [Asa13]. threshold [dS91]. Tightrope [Ori24]. Tightrope-Walking [Ori24]. Time [Ant97, CGL05, Den15, Gau05, Kas04, Kre17, Moz00, Pau16, Rom05, Sim19, Vit16, Wei13, Zin08, Die01, Ear02, Fut02, Har97c, Loc97, Luc97, Til87b, Wei16]. Time-Asymmetric [CGL05]. time-symmetric [Wei16]. times [Kut10]. tithing [Bro98]. token [Jac93]. Tooley [Pag02]. Tools [Bac11, Hoc13, Liv21]. tower [And91, Mam90]. Tracks [Har90]. trading [Das12]. tradition [New01]. Traditions [Boo11]. Traits [CACA19]. trajectories [Acu16]. Transcendental [Pea10]. Transdisciplinarity [Alv11]. Transformation [Cam05, Sim19]. Transformed [Hop13]. transgressing [Lon97]. Translation [Kin17]. Travelling [May22]. Treatment [SS14, Tre21, Tha02]. trial [Bai89, Sea87]. Trials [Kra21]. Truth [Asa12, Baç10, BV14, CT19, Der10, Gur22, Hun11, Käs15, Miz17, Nov10, Roc10, Tod08, Bro90, Cor89, Nor01, Nug91, Pay95, Psi00, Sza03]. Truth-making [Nov10]. Truthmaking [Asa13]. Truths [CT19]. Tuning [WĆ06, Fri17]. Turn [Alb06, Lew16, Pol20]. Twentieth [Leo11]. Twentieth-Century [Leo11]. Twenty [Rio17]. Twenty-first-century [Rio17]. Two [BB10, Boo11, Gin12, Kuo09]. twofold [Mon93b]. Tychomancy [Lan14]. Tying [Let19]. types [ES07, Flo99, Kra92]. Tyranny [Mar12].

Uebel [Okr98]. unconstrained [Cla97].

Underconsideration [Kha10, Miz13, Wra08]. Underdetermination [Jon00, Pie12, Oka00b, Sar00, Zib94]. Understand [Whi15]. Understanding [Cat11, Cor89, Gao17, Gur22, Iva14, Jan15, New12, New13, Rou11, Kos02, Pen97]. unexpected [Acu16]. Unification [Ben17, CW13, Liu03]. Unificationism [dR06]. unified [Bon97]. uniformity [Hun90]. Unifying [Tho14]. Uniqueness [YH24]. unity [CM90, Dav96, Len01]. Univ [Duc05]. universal [Reu11]. universe [Gra18, Wei16, Ear08]. University [Ano02a, Bor13, Duw16, Jac11, Orr24, Pau12, Pla10]. Unmasking [Tod08]. Unmet [And20b]. Unobservable [She21]. unobservables [Cul93]. Unpacking [Tam25]. unpredictability [Rei01]. unreal [Gel03]. Updated [Kuu17]. US$129 [Rad02]. US$22.95 [Ano02a]. US$24.95 [Dud17, Zha15]. US$42 [Rad02]. US$44.95 [Duw16]. US$62.99 [Bor13]. US$64.95 [Ano02a]. US$69.95 [Dud17, Zha15]. Use [Rec17, Web09]. User [Cam14]. Uses [Gur22]. Using [CE11, SC13].

V [Ano06d]. Validity [Ale09, CACA19, Kra21, Yan13]. Value [Ara17, DL23, Ber96]. Values [DL23, Ci06, Rad10, Sve14, Mar87]. Variables [Wag09, Har97c]. variation [Kuj98]. Varieties [Wra15]. Varying [Dur17]. Venus [Rec17]. Verge [Das17]. Verification [Tra20]. verificationism [Bro88]. Versus [Whi15, Psi00, RA16, San96]. vi [Bor13]. Vickers [SS14]. View [Isl17, Liu16, Row10, Duc05, Liu97]. Viewpoints [Tre21]. Views [Cru24, PG06, Sla00]. viii [Dud17]. vindicated [Wat92]. Virtual [Bac11, Wil11]. Virtue [Sha20]. Visual [Ven21]. Vocabulary [Roe17]. void [Mac01]. Volder [Lod01]. Volume [Ano10, Ano11, Ano12b, Ano13b, Ano14a, Ano16, Qin16]. Vries [Dud17]. vs [Bai89].

W [Gat00, Hue25, Zha15]. W. [Flo06, Oka00a]. Waismann [Hei01]. Walking [Ori24]. War [vdT14, Wil92]. Warming [RS10]. Was [Kid16]. water [Nee02]. Waterbears [Bab10]. Watkins [Gat00]. Wave [All17, Laz17, Mat21]. Wave-Function [Mat21]. Wesley [dR06]. Western [Irv97]. Weyl [Gau05]. Where [Har88, HS09, Vic13]. Whigs [Bab10].

Wigner [Ban09]. Wilkes [Ano06d]. Will [Far00, Sch12]. William [BM22]. Williamson [Ill11, Tre21]. Wilson [Bai98]. Within [Zhu17]. Without [Bog20, Bou19]. Wittgenstein [Hei01, Pla10]. Władysław [Big07]. Woodward [Mak23]. words [Cor92]. work [McA02, Nee04]. workable [Kli93]. Works [Iva14, Mar16]. World [Bjö12, Kro08, Lan11, Mei14, Rou11, Bai98, Ghi03, McA97a, Par90]. Worlds [Der10, HS09, May22, Les96, Vai98]. Worldview [OM16]. Worrall [San96]. worry [Ada04, Len05]. writings [Ray09]. wrong [Kno02].

xiii [Rad02]. xiv [Duw16]. xv [Pau12]. xvi [Zha15]. xviii [Ala24]. xxii [Pla10]. xxiii [Ano02a].

yes [Mil91a]. York [Duw16].

zone [Das12]. zoo [Bis03].

References

Abela:1991:BR

[Abe91] Paul Abela. Book review. International Studies in the Philosophy of Science, 5(3):281–283, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                              //www.tandfonline.com//doi/abs/
                                              10.1080/02698599108573400.

Abraham:2016:HBS

[Abr16] Tara H. Abraham. A history of the brain: From stone age surgery to modern neuroscience. International Studies in the Philosophy of Science, 30(1): 91–92, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                              URL http://www.tandfonline.com/

doi/full/10.1080/02698595.2016.
1240477.

Acuna:2016:ITB

[Acu16] Pablo Acuña. Inertial trajectories in de Broglie–Bohm quantum theory: an unexpected problem. International Studies in the Philosophy of Science, 30(3):201–230, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1316107.

Adam:2004:WWA

[Ada04] Matthias Adam. Why worry about theory-dependence? Circularity, minimal empiricality and reliability. International Studies in the Philosophy of Science, 18(2–3):117–132, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000296486.

Agassi:1992:HCA

[Aga92] Joseph Agassi. Heuristic computerassisted, not computerized: Comments on Simon’s Project. International Studies in the Philosophy of Science, 6(1):15–18, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573404.

Agassi:1993:RCR

[Aga93] Joseph Agassi. Rationality: a comment on Raymond Boudon’s paper. International Studies in the Philosophy of Science, 7(1):21–23, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/

                                        10.1080/02698599308573440. See

[Bou93b] and reply [Bou93a].

Andreou:1998:BR

[AKLV98] Chrisoula Andreou, Andris Krumins, Brendan Larvor, and Andre Vellino. Book reviews. International Studies in the Philosophy of Science, 12(3):281–292, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573602.

Alai:2024:BRS

[Ala24] Mario Alai. Book review: Scientific Realism and Laws of Nature. A Metaphysics of Causal Powers, by Michel Ghins, Cham, Switzerland, Springer, Synthese Library, 2024, pp. xviii + 211, EUR 124,79 (hard cover), ISBN 978-3-031-54226-8. International Studies in the Philosophy of Science, 37(3):150–154, 2024. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2024.2402886.

Albert:1993:SRR

[Alb93] Hans Albert. Some remarks on reasons in explaining human action. International Studies in the Philosophy of Science, 7(1):25–27, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573441.

Albedah:2006:GCK

[Alb06] Amani Albedah. A Gadamerian critique of Kuhn’s linguistic turn: Incommensurability revisited. International Studies in the Philosophy of Science, 20(3):323–345, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600961000.

Alcorta:2010:BPS

[Alc10] Candace S. Alcorta. The believing primate: Scientific, philosophical, and theological reflections on the origin of religion. International Studies in the Philosophy of Science, 24(2):233–236, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
484562.

Alegre:2009:EVP

[Ale09] Maria Caamaño Alegre. Experimental validity and pragmatic modes in empirical science. International Studies in the Philosophy of Science, 23(1):19–45, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590902843377.

Allori:2010:QTP

[All10] Valia Allori. Quantum theory: a philosopher’s overview. International Studies in the Philosophy of Science, 24(3):330–333, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522416.

Allori:2013:RMD

[All13] Valia Allori. The road to Maxwell’s Demon: Conceptual foundations of statistical mechanics. International Studies in the Philosophy of Science, 27(4):453–456, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868179.

Allori:2015:R

[All15] Valia Allori. Response. International Studies in the Philosophy of Science, 29(1):94–98, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1079077.

Allori:2017:NAN

[All17] Valia Allori. A new argument for the nomological interpretation of the wave function: The Galilean group and the classical limit of nonrelativistic quantum mechanics. International Studies in the Philosophy of Science, 31(2):177–188, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2018.1424759.

Almeder:2007:PPS

[Alm07] Robert Almeder. Pragmatism and philosophy of science: a critical survey. International Studies in the Philosophy of Science, 21(2):171–195, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701498100.

Alvargonzalez:2011:MIT

[Alv11] David Alvargonzález. Multidisciplinarity, interdisciplinarity, transdisciplinarity, and the sciences. International Studies in the Philosophy of Science, 25(4):387–403, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.623366.

Ambrosio:2010:O

[Amb10] Chiara Ambrosio. Objectivity. International Studies in the Philosophy of Science, 24(1):125–128, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903467184.

Ambrosio:2014:IRR

[Amb14] Chiara Ambrosio. Iconic representations and representative practices. International Studies in the Philosophy of Science, 28(3):255–275, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.959831.

Andersson:1991:TEC

[And91] Dr Gunnar Andersson. The tower experiment and the Copernican revolution. International Studies in the Philosophy of Science, 5(2):143–152, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573386.

Andreas:2015:FMA

[And15] Holger Andreas. A finite memory argument for an axiomatic conception of scientific theories. International Studies in the Philosophy of Science, 29(2):113–127, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119410.

Andreoletti:2020:CCI

[And20a] Mattia Andreoletti. Care & cure. an introduction to philosophy of medicine. International Studies in the Philosophy of Science, 33(1):59–62, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2020.1796038.

Andreoletti:2020:RCS

[And20b] Mattia Andreoletti. Replicability crisis and scientific reforms: Overlooked issues and unmet challenges. International Studies in the Philosophy of Science, 33(3):135–151, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2021.1943292.

Anonymous:1986:EBa

[Ano87a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 1(1):??, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598608573277.

Anonymous:1986:EBb

[Ano87b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 1(2):??, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573287.

Anonymous:1986:NCa

[Ano87c] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 1(1):142, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573286.

Anonymous:1986:NCb

[Ano87d] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 1(2):262–263, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573296.

Anonymous:1987:A

[Ano88a] Anonymous. Addendum. International Studies in the Philosophy of Science, 2(2):2, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573308.

Anonymous:1987:EBa

[Ano88b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 2(1):??, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573297.

Anonymous:1987:EBb

[Ano88c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 2(2):??, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573307.

Anonymous:1987:NCa

[Ano88d] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 2(1):115, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573306.

Anonymous:1987:NCb

[Ano88e] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 2(2):243–244, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573318.

Anonymous:1988:EBa

[Ano89a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 3(1):??, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573319.

Anonymous:1988:EBb

[Ano89b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 3(2):??, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598908573329.

Anonymous:1988:NCa

[Ano89c] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 3(1):117–118, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573328.

Anonymous:1988:NCb

[Ano89d] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 3(2):248–249, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573340.

Anonymous:1990:EBa

[Ano90a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 4(1):??, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573341.

Anonymous:1990:EBb

[Ano90b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 4(2):??, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573351.

Anonymous:1990:EBc

[Ano90c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 4(3):??, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573362.

Anonymous:1990:NCa

[Ano90d] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 4(1):113–114, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573350.

Anonymous:1990:NCb

[Ano90e] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 4(2):226, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573361.

Anonymous:1990:NCc

[Ano90f] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 4(3):325, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573372.

Anonymous:1991:EBa

[Ano91a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 5(1):??, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573373.

Anonymous:1991:EBb

[Ano91b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 5(2):??, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573382.

Anonymous:1991:EBc

[Ano91c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 5(3):??, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573392.

Anonymous:1991:E

[Ano91d] Anonymous. Errata. International Studies in the Philosophy of Science, 5(2):192, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573391.

Anonymous:1991:NCa

[Ano91e] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 5(1):93, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573381.

Anonymous:1991:NCb

[Ano91f] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 5(2):191, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573390.

Anonymous:1991:NCc

[Ano91g] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 5(3):285, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573401.

Anonymous:1992:A

[Ano92a] Anonymous. Apology. International Studies in the Philosophy of Science, 6(3):173–174, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573427.

Anonymous:1992:EBa

[Ano92b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 6(1):??, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573402.

Anonymous:1992:EBb

[Ano92c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 6(2):??, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573418.

Anonymous:1992:EBc

[Ano92d] Anonymous. Editorial board. International Studies in the Philosophy of Science, 6(3):??, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573425.

Anonymous:1992:NCa

[Ano92e] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 6(2):163, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573424.

Anonymous:1992:NCb

[Ano92f] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 6(3):255, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573436.

Anonymous:1993:EBa

[Ano93a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 7(1):??, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573437.

Anonymous:1993:EBb

[Ano93b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 7(2):??, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573456.

Anonymous:1993:EBc

[Ano93c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 7(3):??, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573465.

Anonymous:1993:NCa

[Ano93d] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 7(2):208, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573464.

Anonymous:1993:NCb

[Ano93e] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 7(3):273, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573471.

Anonymous:1994:EBa

[Ano94a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 8(1):??, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573472.

Anonymous:1994:EBb

[Ano94b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 8(2):??, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573486.

Anonymous:1994:EBc

[Ano94c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 8(3):??, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573493.

Anonymous:1994:NCa

[Ano94d] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 8(2):169, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573492.

Anonymous:1994:NCb

[Ano94e] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 8(3):277, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573501.

Anonymous:1995:EBa

[Ano95a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 9(1):??, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573502.

Anonymous:1995:EBb

[Ano95b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 9(2):??, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573510.

Anonymous:1995:EBc

[Ano95c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 9(3):??, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573516.

Anonymous:1995:NC

[Ano95d] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 9(3):198, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573518.

Anonymous:1996:E

[Ano96a] Anonymous. Editorial. International Studies in the Philosophy of Science, 10(2):101, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573532.

Anonymous:1996:EBa

[Ano96b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 10(1):??, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573525.

Anonymous:1996:EBb

[Ano96c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 10(2):??, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573531.

Anonymous:1996:EBc

[Ano96d] Anonymous. Editorial board. International Studies in the Philosophy of Science, 10(3):??, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573538.

Anonymous:1997:EBa

[Ano97a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 11(1):??, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573545.

Anonymous:1997:EBb

[Ano97b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 11(2):??, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573555.

Anonymous:1997:EBc

[Ano97c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 11(3):??, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573566.

Anonymous:1997:EP

[Ano97d] Anonymous. Editor’s page. International Studies in the Philosophy of Science, 11(2):117, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573556.

Anonymous:1997:NN

[Ano97e] Anonymous. Notes and news. International Studies in the Philosophy of Science, 11(1):109, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573554.

Anonymous:1997:ON

[Ano97f] Anonymous. Obituary notices. International Studies in the Philosophy of Science, 11(3):315, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573575.

Anonymous:1998:EBa

[Ano98a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 12(1):??, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573576.

Anonymous:1998:EBb

[Ano98b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 12(2):??, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573585.

Anonymous:1998:EBc

[Ano98c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 12(3):??, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573595.

Anonymous:1998:ONC

[Ano98d] Anonymous. Obituary notice: [Carl Hempel (1905–1997)]. International Studies in the Philosophy of Science, 12(2):205, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573594.

Anonymous:1999:EBa

[Ano99a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 13(1):??, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573603.

Anonymous:1999:EBb

[Ano99b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 13(2):??, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573610.

Anonymous:1999:EBc

[Ano99c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 13(3):??, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573620.

Anonymous:2000:BRa

[Ano00a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 14(1):83–99, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900111927.

Anonymous:2000:BRb

[Ano00b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 14(2):199–210, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590050045764.

Anonymous:2000:BRc

[Ano00c] Anonymous. Book reviews. International Studies in the Philosophy of Science, 14(3):327–336, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/026985900437845.

Anonymous:2001:BRa

[Ano01a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 15(1):99–120, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590020029332.

Anonymous:2001:BRb

[Ano01b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 15(3):311–318, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120080235.

Anonymous:2001:CN

[Ano01c] Anonymous. Critical notice. International Studies in the Philosophy of Science, 15(1):93–98, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590020029323.

Anonymous:2002:BRAa

[Ano02a] Anonymous. Book review: Aristotle’s Philosophy of Biology: Studies in the Origins of Life Science, James G. Lennox, Cambridge, Cambridge University Press, 2001. xxiii + 321 pp, £40.00, US$64.95 (hardback), ISBN 0-521-65027-5; £14.95, US$22.95 (paperback), ISBN 0-521-65976-0. International Studies in the Philosophy of Science, 16(3):291–295, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013364.

Anonymous:2002:BRa

[Ano02b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 16(2):187–198, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145115.

Anonymous:2002:ER

[Ano02c] Anonymous. Editor’s report, 2001.

International Studies in the Philosophy of Science, 16(2):109–110, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145052.

Anonymous:2003:BRa

[Ano03a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 17(1):95–106, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590305231.

Anonymous:2003:BRb

[Ano03b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 17(2):195–203, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859031000160658.

Anonymous:2003:BRc

[Ano03c] Anonymous. Book reviews. International Studies in the Philosophy of Science, 17(3):313–319, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859032000181583.

Anonymous:2003:CLE

[Ano03d] Anonymous. Correspondence: Letter to the editor. International Studies in the Philosophy of Science, 17(2):191–194, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859031000160649.

Anonymous:2003:E

[Ano03e] Anonymous. Editorial. International Studies in the Philosophy of Science, 17(1):3–5, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590305234.

Anonymous:2003:EER

[Ano03f] Anonymous. Editorial: Editor’s report, 2002. International Studies in the Philosophy of Science, 17(2):115–116, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859031000160577.

Anonymous:2003:OSL

[Ano03g] Anonymous. Obituary: Srdan Lelas (1939–2003). International Studies in the Philosophy of Science, 17(3):309–311, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859032000169497.

Anonymous:2004:BRa

[Ano04a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 18(1):95–108, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590412331289288.

Anonymous:2004:BRb

[Ano04b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 18(2–3):231–239, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000296549.

Anonymous:2004:E

[Ano04c] Anonymous. Editorial. International Studies in the Philosophy of Science, 18(2–3):115–116, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000296477.

Anonymous:2005:BRa

[Ano05a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 19(1):85–100, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500051183.

Anonymous:2005:BRb

[Ano05b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 19(2):213–222, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500249555.

Anonymous:2005:BRc

[Ano05c] Anonymous. Book reviews. International Studies in the Philosophy of Science, 19(3):335–344, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500462505.

Anonymous:2005:C

[Ano05d] Anonymous. Correspondence. International Studies in the Philosophy of Science, 19(2):209–212, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500249548.

Anonymous:2005:ER

[Ano05e] Anonymous. Editor’s report, 2004. International Studies in the Philosophy of Science, 19(2):101–103, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500249381.

Anonymous:2006:BRa

[Ano06a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 20(1):113–124, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600641107.

Anonymous:2006:BRb

[Ano06b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 20(2):233–242, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814514.

Anonymous:2006:ER

[Ano06c] Anonymous. Editor’s report, 2005.

International Studies in the Philosophy of Science, 20(2):125–127, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814258.

Anonymous:2006:KVW

[Ano06d] Anonymous. Kathleen V. Wilkes, 1946–2003. International Studies in the Philosophy of Science, 20(1):109–111, 2006. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600641081.

Anonymous:2010:CAI

[Ano10] Anonymous. Contents and author index, volume 24, 2010. International Studies in the Philosophy of Science, 24(4):453–458, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.543361.

Anonymous:2011:VCA

[Ano11] Anonymous. Volume contents & author index. International Studies in the Philosophy of Science, 25(4):427– 431, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
624715.

Anonymous:2012:EB

[Ano12a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 26(4):??, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.757085.

Anonymous:2012:VCA

[Ano12b] Anonymous. Volume contents and author index. International Studies in the Philosophy of Science, 26(4):471– 475, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
749616.

Anonymous:2013:EB

[Ano13a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 27(4):??, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.868628.

Anonymous:2013:VCA

[Ano13b] Anonymous. Volume contents and author index, 2013. International Studies in the Philosophy of Science, 27(4):463–467, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.868187.

Anonymous:2014:CV

[Ano14a] Anonymous. Contents of volume 28, 2014. International Studies in the Philosophy of Science, 28(4):449–453, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
979666.

Anonymous:2014:EB

[Ano14b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 28(4):??, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.1007619.

Anonymous:2015:ISP

[Ano15] Anonymous. International studies in the philosophy of science. International Studies in the Philosophy of Science, 29(4):448–452, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1195154.

Anonymous:2016:VCA

[Ano16] Anonymous. Volume contents and author index, 2016. International Studies in the Philosophy of Science, 30(4):419–423, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1335936.

Anonymous:2017:ISP

[Ano17] Anonymous. International Studies in the Philosophy of Science. International Studies in the Philosophy of Science, 31(4):439–443, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565202.

Anonymous:2021:RA

[Ano21] Anonymous. Reviewers acknowledgement. International Studies in the Philosophy of Science, 34(4):??, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2123491.

Anonymous:2022:RA

[Ano22] Anonymous. Reviewers acknowledgement. International Studies in the Philosophy of Science, 35(3–4):??, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2235854.

Anonymous:2023:RA

[Ano23] Anonymous. Reviewers acknowledgement. International Studies in the Philosophy of Science, 36(4):335–336, 2023. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2024.
2303181.

Anonymous:2024:ER

[Ano24] Anonymous. External referees, 2024.

International Studies in the Philosophy of Science, 37(4):291–292, 2024. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2024.2436801.

Antonopoulos:1997:TNO

[Ant97] Constantin Antonopoulos. Time as non-observational knowledge: How to straighten out ΔEΔt ≥ h. International Studies in the Philosophy of Science, 11(2):165–183, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573561.

Antonelli:2012:FT

[Ant12] G. Aldo Antonelli. Frege’s Theorem.

International Studies in the Philosophy of Science, 26(2):219–222, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.703481.

Arabatzis:2017:WIH

[Ara17] Theodore Arabatzis. What’s in it for the historian of science? Reflections on the value of philosophy of science for history of science. International Studies in the Philosophy of Science, 31(1): 69–82, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.
1370924.

Arabatzis:2020:CPS

[Ara20] Theodore Arabatzis. 20th Century philosophy of science in focus. International Studies in the Philosophy of Science, 33(1):53–57, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2020.1784585.

Arthur:1999:TEP

[Art99] Richard Arthur. On thought experiments as a priori science. International Studies in the Philosophy of Science, 13(3):215–229, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573622.

Asay:2012:TRR

[Asa12] Jamin Asay. Truth, reference and realism. International Studies in the Philosophy of Science, 26(3):345–348, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2012.
                                        731735.

Asay:2013:TPS

[Asa13] Jamin Asay. Three paradigms of scientific realism: a truthmaking account. International Studies in the Philosophy of Science, 27(1):1–21, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.783971.

Aufrecht:2012:PCR

[Auf12] Monica Aufrecht. Popper’s critical rationalism: a philosophical investigation. International Studies in the Philosophy of Science, 26(2):223–225, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
703482.

Aufrecht:2017:RFR

[Auf17] Monica Aufrecht. Reichenbach Falls — and rises? Reconstructing the discovery/justification distinction. International Studies in the Philosophy of Science, 31(2):151–176, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1424760.

Babich:2003:FDK

[Bab03] Babette E. Babich. From Fleck’s Denkstil to Kuhn’s paradigm: Conceptual schemes and incommensurability. International Studies in the Philosophy of Science, 17(1):75–92, 2003. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698590305236.

Babich:2010:TCP

[Bab10] Babette Babich. Towards a critical philosophy of science: Continental beginnings and bugbears, whigs, and waterbears. International Studies in the Philosophy of Science, 24(4):343–391, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
543349.

Babich:2015:CSP

[Bab15] Babette Babich. Calling science pseudoscience: Fleck’s archaeologies of fact and Latour’s ‘biography of an investigation’ in AIDS denialism and homeopathy. International Studies in the Philosophy of Science, 29(1):1–39, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1071550.

Baccarini:1992:REM

[Bac92] Elvio Baccarini. Reflective equilibrium and methodology of science. International Studies in the Philosophy of Science, 6(3):175–180, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573428.

Bac:2010:TOM

[Baç10] Murat Baç. Truth as one and many. International Studies in the Philosophy of Science, 24(1):122–125, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903467176.

BacelarValente:2011:VQN

[Bac11] Mario Bacelar Valente. Are virtual Quanta nothing but formal tools? International Studies in the Philosophy of Science, 25(1):39–53, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.552417.

Baetu:2010:RCG

[Bae10] Tudor M. Baetu. The referential convergence of gene concepts based on classical and molecular analyses. International Studies in the Philosophy of Science, 24(4):411–427, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.543351.

Baetu:2013:CER

[Bae13] Tudor M. Baetu. Chance, experimental reproducibility, and mechanistic regularity. International Studies in the Philosophy of Science, 27(3):253– 271, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
825492.

Baetu:2021:DIA

[Bae21] Tudor M. Baetu. In defence of an inferential account of extrapolation. International Studies in the Philosophy of Science, 34(2):81–100, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.2015678.

Baetu:2025:FSS

[Bae25] Tudor M. Baetu. Failures of scale separation in biology and the problem of inter-level causation. International Studies in the Philosophy of Science, 38(1):1–19, 2025. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2025.2455681.

Bagatur:2017:EPL

[Bag17] Sine Bagatur. The econocracy: The perils of leaving economics to the experts. International Studies in the Philosophy of Science, 31(3):324–327, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2018.
                                        1463690.

Baigrie:1988:NSV

[Bai89] Brian S. Baigrie. Natural selection vs trial and error elimination. International Studies in the Philosophy of Science, 3(2):157–172, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598908573333.

Baigrie:1998:CNC

[Bai98] Brian S. Baigrie. Critical notice: Catherine Wilson’s The invisible world: Early modern philosophy and the invention of the microscope. International Studies in the Philosophy of Science, 12(2):165–174, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573590.

Bakhurst:1991:PED

[Bak91] David Bakhurst. Political emancipation and the domination of nature: The rise and fall of Soviet Prometheanism. International Studies in the Philosophy of Science, 5(3):215–226, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573395.

Bangu:2009:WPM

[Ban09] Sorin Bangu. Wigner’s puzzle for mathematical naturalism. International Studies in the Philosophy of Science, 23(3):245–263, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903196015.

Bantinaki:2012:BMC

[Ban12] Katerina Bantinaki. Beyond mimesis and convention: Representation in art and science. International Studies in the Philosophy of Science, 26(1):114– 118, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
653120.

Bara:1994:DI

[Bar94] Bruno G. Bara. Developing induction. International Studies in the Philosophy of Science, 8(1):31–34, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573475.

Barzegar:2020:PAE

[Bar20] Ali Barzegar. A phenomenological approach to epistemic interpretations of quantum mechanics. International Studies in the Philosophy of Science, 33(3):175–187, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1950483.

Baumgartner:2009:ICE

[Bau09] Michael Baumgartner. Interventionist causal exclusion and nonreductive physicalism. International Studies in the Philosophy of Science, 23(2):161–178, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903006909.

Bandyopadhyay:2010:TDS

[BB10] Prasanta S. Bandyopadhyay and Gordon Brittan, Jr. Two dogmas of strong objective Bayesianism. International Studies in the Philosophy of Science, 24(1):45–65, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903467119.

Blancke:2021:PNO

[BB21] Stefaan Blancke and Maarten Boudry. Pseudoscience as a negative outcome of scientific dialogue: a pragmaticnaturalistic approach to the demarcation problem. International Studies in the Philosophy of Science, 34(3):183– 198, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2022.
                                        2057777.

Bod:2006:ISA

[BBB06] Rens Bod, Mieke Boon, and Marcel Boumans. Introduction to the symposium ‘Applying Science’. International Studies in the Philosophy of Science, 20(1):1–3, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600640943.

Bergstrom:2009:BR

[BBD09] Lars Bergström, Lisa Bortolotti, and Willem B. Drees. Book reviews. International Studies in the Philosophy of Science, 23(1):109–118, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590902843419.

Belot:1995:DO

[Bel95] Gordon Belot. Determinism and ontology. International Studies in the Philosophy of Science, 9(1):85–101, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698599508573508.

Belot:1999:RR

[Bel99] Gordon Belot. Rehabilitating relationalism. International Studies in the Philosophy of Science, 13(1):35–52, 1999. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573606.

Belkind:2007:NCA

[Bel07] Ori Belkind. Newton’s conceptual argument for absolute space. International Studies in the Philosophy of Science, 21(3):271–293, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701589551.

Beni:2017:SRM

[Ben17] Majid Davoody Beni. Structural realism, metaphysical unification, and the ontology and epistemology of patterns. International Studies in the Philosophy of Science, 31(3):285–300, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463691.

Beni:2020:CIS

[Ben20] Majid D. Beni. Causal informational structural realism. International Studies in the Philosophy of Science, 33(2):117–134, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1848346.

Beni:2022:CII

[Ben22] Majid D. Beni. The curious incident of indistinguishable selves: a reply to Nesić. International Studies in the Philosophy of Science, 35(3–4):261–268, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2153215.

Bergstrom:1990:EIA

[Ber90] Lars Bergström. Explanation and interpretation of action. International Studies in the Philosophy of Science, 4(1):3–15, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573342.

Bergstrom:1996:SV

[Ber96] Lars Bergström. Scientific value. International Studies in the Philosophy of Science, 10(3):189–202, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573539.

Blondeau:2012:TIC

[BG12] Julien Blondeau and Michel Ghins. Is there an intrinsic criterion for causal lawlike statements? International Studies in the Philosophy of Science, 26(4):381–401, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.748496.

Bica:2022:NMP

[Bic22] Daian Bica. Navigating Massimi’s perspectival garden with inferential forking paths. International Studies in the Philosophy of Science, 35(3–4):291– 303, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2023.
2191168.

Bielik:2016:RW

[Bie16] Lukáš Bielik. Reasons why. International Studies in the Philosophy of Science, 30(4):407–409, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331974.

Bigaj:2007:OWK

[Big07] Tomasz Bigaj. Obituary: Władysław Krajewski, 1919–2006. International Studies in the Philosophy of Science, 21(1):91–93, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701305818.

Bird:2012:RNK

[Bir12] Alexander Bird. Referring to natural kind thingamajigs, and what they are: a reply to Needham. International Studies in the Philosophy of Science, 26(1):103–109, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.653111. See

[Nee12].

Bishop:2003:PIF

[Bis03] Michael A. Bishop. The pessimistic induction, the flight to reference and the metaphysical zoo. International Studies in the Philosophy of Science, 17(2):161–178, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859031000160621.

Bailer-Jones:2003:WSM

[BJ03] Daniela M. Bailer-Jones. When scientific models represent. International Studies in the Philosophy of Science, 17(1):59–74, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590305238.

Bjorkman:2012:LJE

[Bjö12] Beyza Björkman. Linguistic justice for Europe and for the world. International Studies in the Philosophy of Science, 26(3):354–357, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.731738.

Brzovic:2017:KP

[BJS17] Zdenka Brzović, Marko Jurjako, and Predrag Sustar. The kindness of psychopaths. International Studies in the Philosophy of Science, 31(2):189–211, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2018.
1424761.

Boyer-Kassem:2014:LMC

[BK14] Thomas Boyer-Kassem. Layers of models in computer simulations. International Studies in the Philosophy of Science, 28(4):417–436, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979670.

Blachowicz:1995:ECP

[Bla95] James Blachowicz. Elimination, correction and Popper’s evolutionary epistemology. International Studies in the Philosophy of Science, 9(1):5–17, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573503.

Blachowicz:1996:AA

[Bla96] James Blachowicz. Ampliative abduction. International Studies in the Philosophy of Science, 10(2):141–157, 1996. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573535.

Brown:2022:WHN

[BM22] James Robert Brown and Cheryl Misak. William H. Newton-Smith (1943–2023). International Studies in the Philosophy of Science, 35(2):205– 208, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2023.
2208416.

Beisbart:2012:WMC

[BN12] Claus Beisbart and John D. Norton. Why Monte Carlo simulations are inferences and not experiments. International Studies in the Philosophy of Science, 26(4):403–422, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.748497.

Boylan:1991:CET

[BO91] Thomas A. Boylan and Pascal F. O’Gorman. The critique of equilibrium theory in economic methodology: a constructive empiricist perspective. International Studies in the Philosophy of Science, 5(2):131–142, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573385.

Bod:2006:TGM

[Bod06] Rens Bod. Towards a general model of applying science. International Studies in the Philosophy of Science, 20(1):5–25, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600640950.

Bogen:2004:ACO

[Bog04] Jim Bogen. Analysing causality: The opposite of counterfactual is factual. International Studies in the Philosophy of Science, 18(1):3–26, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590412331289233.

Boge:2020:RIR

[Bog20] Florian J. Boge. Realism without interphenomena: Reichenbach’s cube, Sober’s evidential realism, and quantum solipsism. International Studies in the Philosophy of Science, 33(4):231– 246, 2020. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2021.
1964335.

Bokulich:2003:QMS

[Bok03] Alisa Bokulich. Quantum measurements and supertasks. International Studies in the Philosophy of Science, 17(2):127–136, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859031000160595.

Bokulich:2011:ICB

[Bok11] Peter Bokulich. Interactions and the consistency of black hole complementarity. International Studies in the Philosophy of Science, 25(4):371–386, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
623365.

Bolotovsky:1986:BDN

[Bol87] Boris M. Bolotovsky. The birth and development of new knowledge: Oliver Heaviside. International Studies in the Philosophy of Science, 1 (2):143–175, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573289.

Boldyrev:2012:PSS

[Bol12] Ivan A. Boldyrev. Philosophy of science or science and technology studies? Economic methodology and auction theory. International Studies in the Philosophy of Science, 26(3):289– 307, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
731732.

Boniolo:1997:UTM

[Bon97] Giovanni Boniolo. On a unified theory of models and thought experiments in natural sciences. International Studies in the Philosophy of Science, 11(2):121–142, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573558.

Boon:2004:TIS

[Boo04] Mieke Boon. Technological instruments in scientific experimentation. International Studies in the Philosophy of Science, 18(2–3):221–230, 2004. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000296530.

Boon:2006:HSA

[Boo06] Mieke Boon. How science is applied in technology. International Studies in the Philosophy of Science, 20(1):27–47, 2006. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590600640992.

Boon:2011:TSR

[Boo11] Mieke Boon. Two styles of reasoning in scientific practices: Experimental and mathematical traditions. International Studies in the Philosophy of Science, 25(3):255–278, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.605248.

Boon:2016:RJS

[Boo16] Eva Boon. A remarkable journey: The story of evolution. International Studies in the Philosophy of Science, 30(1):88–90, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1240475.

Borutti:2012:APN

[Bor12] Silvana Borutti. After postmodernism: a naturalistic reconstruction of the humanities. International Studies in the Philosophy of Science, 26(4):467–470, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.748503.

Borrelli:2013:BRN

[Bor13] Arianna Borrelli. Book review: Niels Bohr and the Quantum Atom: The Bohr Model of Atomic Structure 1913–1925 [Helge Kragh, Oxford, Oxford University Press, 2012. vi + 410 pp. ISBN 978-0-19-965498-7, £35.50, US$62.99 (hardback)]. International Studies in the Philosophy of Science, 27(2):222–224, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.813250.

Boudon:1993:MGR

[Bou93a] Raymond Boudon. More on “good reasons”: Reply to critics. International Studies in the Philosophy of Science, 7(1):87–102, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573455. See

[Bou93b, Aga93, BR93, Fuc93, Kli93, Lin93, Mon93a].

Boudon:1993:TST

[Bou93b] Raymond Boudon. Toward a synthetic theory of rationality. International Studies in the Philosophy of Science, 7(1):5–19, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573439.

Bourrat:2019:WSC

[Bou19] Pierrick Bourrat. In what sense can there be evolution by natural selection without perfect inheritance? International Studies in the Philosophy of Science, 32(1):13–31, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1623485.

Boumans:2024:SSP

[Bou24] Marcel Boumans. Shaping social phenomena. International Studies in the Philosophy of Science, 37(3):95–108, 2024. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2024.
2399420.

Ben-Rafael:1993:ICC

[BR93] Eliezer Ben-Rafael. Integrating cooperation and conflict: Comments on Raymond Boudon’s paper. International Studies in the Philosophy of Science, 7(1):29–31, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573442. See

[Bou93b] and reply [Bou93a].

Brady:1994:AJM

[Bra94] Michael E. Brady. On the application of J. M. Keynes’s approach to decision making. International Studies in the Philosophy of Science, 8(2):99–112, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573487.

Bradley:2017:WBC

[Bra17] Darren Bradley. What is Bayesian confirmation for? International Studies in the Philosophy of Science, 31(3):229–241, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2018.1463692.

Brenner:2015:TCB

[Bre15] Anastasios Brenner. Is there a cultural barrier between historical epistemology and analytic philosophy of science? International Studies in the Philosophy of Science, 29(2):201–214, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1119421.

Brighouse:1999:ICM

[Bri99] Carolyn Brighouse. Incongruent counterparts and modal relationism. International Studies in the Philosophy of Science, 13(1):53–68, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573607.

Brigandt:2010:SRM

[Bri10] Ingo Brigandt. Scientific reasoning is material inference: Combining confirmation, discovery, and explanation. International Studies in the Philosophy of Science, 24(1):31–43, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903467101.

Brioschi:2017:NEL

[Bri17] Maria Regina Brioschi. Nature as event: The lure of the possible. International Studies in the Philosophy of Science, 31(4):427–429, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565208.

Brown:1986:TES

[Bro87] James Robert Brown. Thought experiments since the scientific revolution. International Studies in the Philosophy of Science, 1(1):1–15, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573279.

Brown:1987:EBV

[Bro88] James Robert Brown. Einstein’s brand of verificationism. International Studies in the Philosophy of Science, 2(1):33–54, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573301.

Brown:1988:PEW

[Bro89] James Robert Brown. Platonic explanation: Or, what abstract entities can do for you. International Studies in the Philosophy of Science, 3(1):51–67, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573324.

Brown:1990:PTL

[Bro90] James Robert Brown. Proof and truth in Lakatos’s masterpiece. International Studies in the Philosophy of Science, 4(2):117–130, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573352.

Brown:1994:PLR

[Bro94] James Robert Brown. Platonism and laws: a reply to Demetra Sfendoni-Mentzou. International Studies in the Philosophy of Science, 8(3):243–246, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573498. See

[SM94].

Brown:1997:RSR

[Bro97] Harvey R. Brown. On the role of special relativity in general relativity. International Studies in the Philosophy of Science, 11(1):67–81, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573551. Festschrift

for J. R. Lucas.

Brown:1998:IT

[Bro98] James Robert Brown. Intellectual tithing. International Studies in the Philosophy of Science, 12(1):5–6, 1998. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573577.

Brown:1999:D

[Bro99a] James Robert Brown. “Dubrovnik”. International Studies in the Philosophy of Science, 13(2):101, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573611.

Brown:1999:ISI

[Bro99b] James Robert Brown. Introduction to the special issue on rationality. International Studies in the Philosophy of Science, 13(3):213, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573621.

Brown:2001:E

[Bro01] James Robert Brown. Editorial. International Studies in the Philosophy of Science, 15(2):125–126, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120058988.

Brunnander:2013:NSM

[Bru13] Björn Brunnander. Natural selection and multiple realisation: a closer look. International Studies in the Philosophy of Science, 27(1):73–83, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.783972.

Brysse:2012:PPI

[Bry12] Keynyn Brysse. Paleontology: a philosophical introduction. International Studies in the Philosophy of Science, 26(1):111–114, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.653112.

Brzovi:2016:NKC

[Brz16] Zdenka Brzovi. Natural kinds and classification in scientific practice. International Studies in the Philosophy of Science, 30(4):412–415, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331975.

Brown:1995:MRP

[BS95] Harvey R. Brown and Roland Sypel. On the meaning of the relativity principle and other symmetries. International Studies in the Philosophy of Science, 9(3):235–253, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573522.

Bolton:2022:RNP

[BS22] Derek Bolton and Predrag Sustar. Regulation and the normativity problem. International Studies in the Philosophy of Science, 35(2):135–151, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2149050.

Belot:1997:BR

[BSU97] Gordon Belot, Spas Spassov, and Alasdair Urquhart. Book reviews. International Studies in the Philosophy of Science, 11(3):305–313, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573574.

Bandyopadhyay:2016:NBA

[BTB16] Prasanta S. Bandyopadhyay, Mark L. Taper, and Gordon Brittan, Jr. Non-Bayesian accounts of evidence: Howson’s counterexample countered.

International Studies in the Philosophy of Science, 30(3):291–298, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.

1316109. See reply [How16].

Bub:2003:ORC

[Bub03] Jeffrey Bub. Obituary: Rob Clifton (1964–2002). International Studies in the Philosophy of Science, 17(1):93– 94, 2003. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698590305237.

Bulle:2009:UWC

[Bul09] Nathalie Bulle. Under what conditions can formal models of social action claim explanatory power? International Studies in the Philosophy of Science, 23(1):47–64, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590902843385.

Bursten:2019:MMM

[Bur19] Julia R. S. Bursten. Macroscopic metaphysics: Middle-sized objects and longish processes. International Studies in the Philosophy of Science, 32(1): 63–64, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.
1615659.

Busch:2003:WSC

[Bus03] Jacob Busch. What structures could not be. International Studies in the Philosophy of Science, 17(3):211–225, 2003. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859032000169433.

Busch:2011:SRI

[Bus11] Jacob Busch. Scientific realism and the indispensability argument for mathematical realism: a marriage made in Hell. International Studies in the Philosophy of Science, 25(4):307– 325, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
623361.

Buskell:2019:BRW

[Bus19] Andrew Buskell. Book review: What’s Left of Human Nature? A Post-Essentialist, Pluralist, and Interactive Account of a Contested Concept, by Maria Kronfelder, Cambridge, MA, The MIT Press, Oct. 2018, 336 pp, $45.00, £38.00 (hardback), ISBN 978-0-262-03841-6. International Studies in the Philosophy of Science, 32(2):137– 140, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.
1704616.

Butterfield:1987:SD

[But88] Jeremy Butterfield. Substantivalism and determinism. International Studies in the Philosophy of Science, 2(1):10–32, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573300.

Butterfield:2007:RRC

[But07] Jeremy Butterfield. Reconsidering relativistic causality. International Studies in the Philosophy of Science, 21(3):295–328, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701589585.

Buzzoni:2010:RPH

[Buz10] Marco Buzzoni. Rethinking Popper and his legacy. International Studies in the Philosophy of Science, 24(3):309– 321, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
522413.

Buzzoni:2014:ATC

[Buz14] Marco Buzzoni. The agency theory of causality, anthropomorphism, and simultaneity. International Studies in the Philosophy of Science, 28(4):375– 395, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
979668.

Boudry:2014:NSD

[BV14] Maarten Boudry and Michael Vlerick. Natural selection does care about truth. International Studies in the Philosophy of Science, 28(1):65–77, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915651.

Beck:2024:GRR

[BV24] Pieter T. L. Beck and Maarten Van Dyck. Good reasons, real questions and proper aims: Hasok Chang on the rationality of the chemical revolution. International Studies in the Philosophy of Science, 37(3):127–145, 2024. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2024.2418278.

Caamano-Alegre:2019:OTV

[CACA19] María Caamaño-Alegre and José Caamaño-Alegre. From ontological traits to validity challenges in social science: The cases of economic experiments and research questionnaires. International Studies in the Philosophy of Science, 32(2):101–127, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1682773.

Camilleri:2005:HTK

[Cam05] Kristian Camilleri. Heisenberg and the transformation of Kantian philosophy. International Studies in the Philosophy of Science, 19(3):271–287, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590500462273.

Campaner:2014:CUG

[Cam14] Raffaella Campaner. Causation: a user’s guide. International Studies in the Philosophy of Science, 28(2):221– 223, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/

doi/full/10.1080/02698595.2014.
932532.

Carrier:1986:NIS

[Car87] Martin Carrier. Newton’s ideas on the structure of matter and their impact on eighteenth-century chemistry: Some historical and methodological remarks. International Studies in the Philosophy of Science, 1(1):85–105, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573283.

Carrier:1998:DPL

[Car98a] Martin Carrier. In defense of psychological laws. International Studies in the Philosophy of Science, 12(3):217–232, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573597.

Carrier:1998:PSG

[Car98b] Martin Carrier. The philosophy of science in German-speaking countries. International Studies in the Philosophy of Science, 12(1):45–86, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573582.

Carman:2010:DPD

[Car10a] Christián C. Carman. On the determination of planetary distances in the Ptolemaic system. International Studies in the Philosophy of Science, 24(3):257–265, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.522408.

Carter:2010:DPA

[Car10b] Jessica Carter. Diagrams and proofs in analysis. International Studies in the Philosophy of Science, 24(1):1–14, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590903467085.

Carrier:2023:FRHa

[Car23a] Martin Carrier. Fake research and harmful findings: Introduction to the special issue. International Studies in the Philosophy of Science, 36(3):167– 171, 2023. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2023.
                                        2284624.

Carrier:2023:FRHb

[Car23b] Martin Carrier. Fake research: How can we recognise it and respond to it? International Studies in the Philosophy of Science, 36(3):247–264, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2023.2187206.

Cassini:2017:HDS

[Cas17] Alejandro Cassini. How to do science with models: a philosophical primer. International Studies in the Philosophy of Science, 31(1):105–107, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1370926.

Cat:2011:UU

[Cat11] Jordi Cat. On understanding understanding. International Studies in the Philosophy of Science, 25(4):405–411, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
623368.

Cat:2006:BR

[CBG06] Jordi Cat, Jonathan Bain, and John Gascoigne. Book reviews. International Studies in the Philosophy of Science, 20(3):347–357, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600961018.

Cartwright:2011:HCU

[CE11] Nancy Cartwright and Sophia Efstathiou. Hunting causes and using them: Is there no bridge from here to there? International Studies in the Philosophy of Science, 25(3):223– 241, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
605245.

Carrier:2006:ELL

[CF06] Martin Carrier and Patrick Finzer.

Explanatory loops and the limits of genetic reductionism. International Studies in the Philosophy of Science, 20(3):267–283, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600960937.

Campaner:2012:EAC

[CG12] Raffaella Campaner and Maria Carla Galavotti. Evidence and the assessment of causal relations in the health sciences. International Studies in the Philosophy of Science, 26(1):27–45, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
653113.

Castagnino:2005:TAI

[CGL05] Mario Castagnino, Manuel Gadella, and Olimpia Lombardi. Time’s arrow and irreversibility in time-asymmetric quantum mechanics. International Studies in the Philosophy of Science, 19(3):223–243, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500460806.

Chang:1997:CPC

[Cha97] Hasok Chang. Can Planck’s constant be measured with classical mechanics? International Studies in the Philosophy of Science, 11(3):223–243, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573568.

Chakravartty:2004:SFS

[Cha04] Anjan Chakravartty. Structuralism as a form of scientific realism. International Studies in the Philosophy of Science, 18(2–3):151–171, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000296503.

Chang:2011:PGS

[Cha11a] Hasok Chang. The philosophical grammar of scientific practice. International Studies in the Philosophy of Science, 25(3):205–221, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.605244.

Chatti:2011:EST

[Cha11b] Saloua Chatti. Extensionalism and scientific theory in Quine’s philosophy. International Studies in the Philosophy of Science, 25(1):1–21, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.552415.

Chester:2002:SI

[Che02] Marvin Chester. Is symmetry identity? International Studies in the Philosophy of Science, 16(2):111–124, 2002. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145061.

Chen:2023:ENM

[Che23a] Bohang Chen. Ernest Nagel’s model of reduction and theory change. International Studies in the Philosophy of Science, 36(1):19–37, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2204792.

Chen:2023:TEA

[Che23b] Ruey-Lin Chen. Are there experimental arguments independent of theories? In defense of a Hackingian approach to the scientific realism debate. International Studies in the Philosophy of Science, 36(4):279–297, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2294465.

Chimisso:2010:FSP

[Chi10a] Cristina Chimisso. The five senses: a philosophy of mingled bodies. International Studies in the Philosophy of Science, 24(2):226–228, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.484559.

Chirimuuta:2010:RRE

[Chi10b] Mazviita Chirimuuta. The red and the real: An essay on color ontology. International Studies in the Philosophy of Science, 24(3):339–342, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522419.

Chiffi:2019:SHM

[Chi19] Daniele Chiffi. Springer handbook of model-based science. International Studies in the Philosophy of Science, 32(1):65–67, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1615662.

Carter:2007:BR

[CHLL07] Jessica Carter, Jussi Haukioja, Mariska E. M. P. J. Leunissen, and Brendan Larvor. Book reviews. International Studies in the Philosophy of Science, 21(2):213–225, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701498175.

Ci:2003:RAM

[Ci03] Jiwei Ci. Review article: From modest realism to a democratic conception of science. International Studies in the Philosophy of Science, 17(3):301–307, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859032000169488.

Ci:2006:CSV

[Ci06] Jiwei Ci. Can scientific values be extended to the public sphere? International Studies in the Philosophy of Science, 20(2):219–231, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814506.

Ci:2011:SSD

[Ci11] Jiwei Ci. Social science and the diversity of its roles for democracy. International Studies in the Philosophy of Science, 25(2):179–190, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574860.

Caruana:1996:BR

[CIL96] Louis Caruana, Sam Ingus, and Brendan Larvor. Book reviews. International Studies in the Philosophy of Science, 10(2):173–179, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573537.

Code:2008:BR

[CJD^{+}08] Lorraine Code, Struan Jacobs, Deepanwita Dasgupta, Charles R. Twardy, and Rafaela Hillerbrand. Book reviews. International Studies in the Philosophy of Science, 22(1):97– 114, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698590802280969.

Costa-Leite:2011:NEK

[CL11] Alexandre Costa-Leite. New essays on the knowability paradox. International Studies in the Philosophy of Science, 25(2):194–196, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.574865.

Carrier:2019:CMH

[CL19] Martin Carrier and Johannes Lenhard. Climate models: How to assess their reliability. International Studies in the Philosophy of Science, 32(2):81–100, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2019.
                                        1644722.

Clarke:1997:PU

[Cla97] Steve Clarke. Pluralism unconstrained. International Studies in the Philosophy of Science, 11(2):143–146, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573559.

Clarke:2013:BRH

[Cla13] Imogen Clarke. Book review: A History of the Electron: J. J. and G. P. Thomson. International Studies in the Philosophy of Science, 27(1):104–107, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
783973.

Carrier:1990:US

[CM90] Martin Carrier and Jürgen Mittelstrass. The unity of science. International Studies in the Philosophy of Science, 4(1):17–31, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573343.

Carrier:1995:R

[CM95] Martin Carrier and Jürgen Mittelstrass. Rejoinder. International Studies in the Philosophy of Science, 9(1):103–104, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573509.

Claveau:2014:MCG

[CMF14] François Claveau and Luis Mireles-Flores. On the meaning of causal generalisations in policy-oriented economic research. International Studies in the Philosophy of Science, 28(4):397–416, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.979669.

Claveau:2016:CGP

[CMF16] François Claveau and Luis Mireles-Flores. Causal generalisations in policy-oriented economic research: An inferentialist analysis. International Studies in the Philosophy of Science, 30(4):383–398, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1331976.

Cohen:1994:JLT

[Coh94] L. Jonathan Cohen. Johnson– Laird’s theory of induction. International Studies in the Philosophy of Science, 8(1):35–36, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573476.

Colaco:2020:RAE

[Col20] David Colaço. Review of advances in experimental philosophy of science. International Studies in the Philosophy of Science, 33(1):57–59, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2020.1795781.

Collodel:2022:EEM

[Col22] Matteo Collodel. Ehrenhaft’s experiments on magnetic monopoles: Reconsidering the Feyerabend–Ehrenhaft connection. International Studies in the Philosophy of Science, 35(1):69– 94, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2123099.

Corbi:1988:UTE

[Cor89] Josep Corbí. Understanding, truth, and explanation. International Studies in the Philosophy of Science, 3(1):19–34, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573322.

Cordeschi:1992:FWR

[Cor92] Roberto Cordeschi. A few words on representation and meaning. International Studies in the Philosophy of Science, 6(1):19–21, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573405.

Craig:2000:EP

[Cra00] William Craig. The extent of the present. International Studies in the Philosophy of Science, 14(2):165–185, 2000. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590050045746.

Crespi:1993:GRS

[Cre93] Franco Crespi. The good reasons of the synthetic theory of rationality. International Studies in the Philosophy of Science, 7(1):33–35, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573443.

Cruse:2004:SRR

[Cru04] Pierre Cruse. Scientific realism, Ramsey sentences and the reference of theoretical terms. International Studies in the Philosophy of Science, 18(2–3):133–149, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000296495.

Crull:2024:RQF

[Cru24] Elise Crull. Realism with quantum faces: The Leggett–Garg inequalities as a case study for Feyerabend’s views. International Studies in the Philosophy of Science, 37(4):195–217, 2024. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2024.2371633.

Cat:2024:QHD

[CS24] Jordi Cat and Jamie Shaw. From quantum holism to the disunity of science and social activism: The Cat– Feyerabend correspondence. International Studies in the Philosophy of Science, 37(4):243–290, 2024. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2024.2424154.

Cevolani:2019:WAT

[CT19] Gustavo Cevolani and Luca Tambolo. Why adding truths is not enough: A reply to Mizrahi on progress as approximation to the truth. International Studies in the Philosophy of Science, 32(2):129–135, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.

1623502. See [Miz17].

Culjak:1993:BU

[Cul93] Zvonimir Culjak. Bosković’s unobservables. International Studies in the Philosophy of Science, 7(3):211–224, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698599308573466.

Culjak:1995:SAE

[Cul95] Zvonimir Culjak. Some aspects of explanation in boskovic. International Studies in the Philosophy of Science, 9(1):73–84, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573507.

Cutting:1987:GBP

[Cut88] Gary Cutting. Gaston Bachelard’s philosophy of science. International Studies in the Philosophy of Science, 2(1):55–71, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573302.

Clarke:2009:SIP

[CW09] Steve Clarke and Adrian Walsh. Scientific imperialism and the proper relations between the sciences. International Studies in the Philosophy of Science, 23(2):195–207, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903007170.

Clarke:2013:IPD

[CW13] Steve Clarke and Adrian Walsh. Imperialism, progress, developmental teleology, and interdisciplinary unification. International Studies in the Philosophy of Science, 27(3):341–351, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2013.
                                        825493.

DAgostino:1990:ARS

[D’A90] Fred D’Agostino. The aimless rationality of science. International Studies in the Philosophy of Science, 4(1):33–50, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573344.

Danne:2020:EMP

[Dan20] Nicholas Danne. An extramathematical program explanation of color experience. International Studies in the Philosophy of Science, 33(3):153–173, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2021.1951637.

Darrigol:2007:FPS

[Dar07] Olivier Darrigol. A Faradayan principle for selecting classical field theories. International Studies in the Philosophy of Science, 21(1):35–55, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701305768.

Dasgupta:2012:CPT

[Das12] Deepanwita Dasgupta. Creating a peripheral trading zone: Satyendra Nath Bose and Bose–Einstein statistics, doing science in the role of an outsider. International Studies in the Philosophy of Science, 26(3):259–287, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.731731.

Dasgupta:2017:RPS

[Das17] Deepanwita Dasgupta. The rightful place of science: Science on the verge. International Studies in the Philosophy of Science, 31(1):108–110, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1370927.

Davies:1996:EDU

[Dav96] David Davies. Explanatory disunities and the unity of science. International Studies in the Philosophy of Science, 10(1):5–21, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573526.

Davies:1998:MAS

[Dav98] David Davies. McAllister’s aesthetics in science: a critical notice. International Studies in the Philosophy of Science, 12(1):25–32, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573580.

Davies:1999:LSR

[Dav99] David Davies. Living in the “space of reasons”: The “rationality debate” revisited. International Studies in the Philosophy of Science, 13(3):231–244, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573623.

daCosta:1993:MTA

[dCF93] Newton C. A. da Costa and Steven French. A model theoretic approach to ‘natural’ reasoning. International Studies in the Philosophy of Science, 7(2):177–190, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573462.

Douven:2002:DAA

[DD02] Igor Douven and Henk W. De Regt. A Davidsonian argument against incommensurability. International Studies in the Philosophy of Science, 16(2):157–169, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145098.

DeMey:1992:SDC

[De 92] Marc De Mey. Scientific discovery: Cold fusion of ideas? International Studies in the Philosophy of Science, 6(1):23–27, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573406.

DeRegt:2002:BPS

[De 02] Henk W. De Regt. Beauty in physical science circa 2000. International Studies in the Philosophy of Science, 16(1):95–103, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120118855.

DeGroot:2015:LE

[De 15a] Jean De Groot. Letter to the Editor.

International Studies in the Philosophy of Science, 29(4):431–433, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1195146.

DePace:2015:CPA

[De 15b] Anna De Pace. Copernicus: Platonist astronomer–philosopher. Cosmic order, the movement of the Earth, and the scientific revolution. International Studies in the Philosophy of Science, 29(3):337–340, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1179044.

DelSanto:2022:BMD

[Del22] Flavio Del Santo. Beyond method:

The diatribe between Feyerabend and Popper over the foundations of quantum mechanics. International Studies in the Philosophy of Science, 35(1):5– 22, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2031430.

Deng:2015:DMT

[Den15] Natalja Deng. Debates in the metaphysics of time. International Studies in the Philosophy of Science, 29(3):340–344, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1179045.

Deroy:2010:WTP

[Der10] Ophelia Deroy. Worlds of truth: a philosophy of knowledge. International Studies in the Philosophy of Science, 24(4):446–448, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.543357.

DesRoches:2018:CCC

[Des18] C. Tyler DesRoches. On the concept and conservation of critical natural capital. International Studies in the Philosophy of Science, 32(3–4):207– 228, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2020.
                                        1788347.

Detlefsen:1988:FHM

[Det89] Michael Detlefsen. Fregean hierarchies and mathematical explanation. International Studies in the Philosophy of Science, 3(1):97–116, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573327.

Dotson:1994:BPQ

[DF94] Allen Dotson and Henry Folse. Bearers of properties in the quantum mechanical description of nature. International Studies in the Philosophy of Science, 8(3):179–194, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573494.

Dieks:2001:STR

[Die01] Dennis Dieks. Space–time relationism in Newtonian and relativistic physics. International Studies in the Philosophy of Science, 15(1):5–17, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590020029279.

Dimitrakos:2016:KNK

[Dim16] Thodoris Dimitrakos. Kuhnianism and neo-Kantianism: On Friedman’s account of scientific change. International Studies in the Philosophy of Science, 30(4):361–382, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331977.

DiSalle:1992:ENE

[DiS92] Robert DiSalle. Einstein, Newton and the empirical foundations of spacetime geometry. International Studies in the Philosophy of Science, 6(3):181–189, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573429.

Benedetto:2023:SVD

[DL23] Matteo De Benedetto and Michele Luchetti. Specialisation by value divergence: The role of epistemic values in the branching of scientific disciplines. International Studies in the Philosophy of Science, 36(2):121–141, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2023.2233822.

Dowell:2008:BR

[DLVW08] Janice Dowell, Soazig Le Bihan, Janet Vertesi, and Sally Wyatt. Book reviews. International Studies in the Philosophy of Science, 22(3):327–339, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590802567431.

deMelo-Martin:2023:HAP

[dMM23] Inmaculada de Melo-Martín. On the harms of agnotological practices and how to address them. International Studies in the Philosophy of Science, 36(3):211–228, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2023.2178838.

Dorato:2007:RTB

[Dor07] Mauro Dorato. Relativity theory between structural and dynamical explanations. International Studies in the Philosophy of Science, 21(1):95–102, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590701305891.

Douven:2000:TTP

[Dou00] Igor Douven. Theoretical terms and the Principle of the Benefit of Doubt. International Studies in the Philosophy of Science, 14(2):135–146, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590050045728.

Dowe:1992:EDC

[Dow92] Phil Dowe. An empiricist defence of the causal account of explanation. International Studies in the Philosophy of Science, 6(2):123–128, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573420.

dePaz:2015:PCH

[dP15] María de Paz. Poincaré’s classification of hypotheses and their role in natural science. International Studies in the Philosophy of Science, 29(4):369– 382, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1195142.

deRegt:2006:WSC

[dR06] Henk W. de Regt. Wesley Salmon’s complementarity thesis: Causalism and unificationism reconciled? International Studies in the Philosophy of Science, 20(2):129–147, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814308.

deRidder:2016:SNO

[dR16] Jeroen de Ridder. Scientism: The new orthodoxy. International Studies in the Philosophy of Science, 30(1):93– 95, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2016.
                                        1240479.

deSousa:1988:KKI

[dS89] Ronald de Sousa. Kinds of kinds: Individuality and biological species. International Studies in the Philosophy of Science, 3(2):119–135, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598908573331.

deSousa:1990:SS

[dS90] Ronald de Sousa. The sociology of sociobiology. International Studies in the Philosophy of Science, 4(3):271–283, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573367.

deSousa:1991:DEK

[dS91] Ronald de Sousa. Does the eye know calculus? The threshold of representation in classical and connectionist models. International Studies in the Philosophy of Science, 5(2):171–185, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573388.

Dziurosz-Serafinowicz:2012:DOB

[DS12] Patryk Dziurosz-Serafinowicz. In defence of objective Bayesianism. International Studies in the Philosophy of Science, 26(3):348–351, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.731736.

Diez:2021:EFR

[DS21] José Díez and Albert Solé. On the elusive formalisation of the risky condition for hypothesis testing. International Studies in the Philosophy of Science, 34(4):199–219, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2084705.

Ducheyne:2005:MMN

[Duc05] Steffen Ducheyne. Mathematical models in Newton’s Principia: a new view of the ‘Newtonian style’. Comment on: The Newtonian revolution [Cambridge Univ. Press, Cambridge, 1980; MR0599360] by I. B. Cohen. International Studies in the Philosophy of Science, 19(1):1–19, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500051035.

Dudhwala:2017:BRB

[Dud17] Farzana Dudhwala. Book review: Bruno Latour [Gerard de Vries, Cambridge, Polity Press, 2016, viii + 221 pp., ISBN 978-0-7456-5062-3, £55.00, US$69.95, EUR 76.28 (hardback); ISBN 978-0-7456-5063-0, £16.99, US$24.95, EUR 23.56 (paperback)]. International Studies in the Philosophy of Science, 31(1):103–105, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.
1370928.

Duran:2017:VES

[Dur17] Juan M. Durán. Varying the explanatory span: Scientific explanation for computer simulations. International Studies in the Philosophy of Science, 31(1):27–45, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1370929.

Duwell:2016:BRB

[Duw16] Armond Duwell. Book review: Bananaworld: Quantum Mechanics for Primates [Jeffrey Bub, New York, Oxford University Press, 2016, xiv + 273 pp. ISBN 978-0-19-871853-6, US$44.95, £25.00 (hardback)]. International Studies in the Philosophy of Science, 30(3):309–311, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1316110.

Earman:2002:WTR

[Ear02] John Earman. What time reversal invariance is and why it matters. International Studies in the Philosophy of Science, 16(3):245–264, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013328.

Earman:2004:CPS

[Ear04] John Earman. Curie’s Principle and spontaneous symmetry breaking. International Studies in the Philosophy of Science, 18(2–3):173–198, 2004. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000311299.

Earman:2008:RPG

[Ear08] John Earman. Reassessing the prospects for a growing block model of the Universe. International Studies in the Philosophy of Science, 22(2):135–164, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802496680.

Eronen:2014:ISN

[EB14] Markus I. Eronen and Daniel S.

Brooks. Interventionism and supervenience: a new problem and provisional solution. International Studies in the Philosophy of Science, 28(2):185–202, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
932529.

Eder:2010:DTR

[Ede10] Anna-Maria A. Eder. Decision theory and rationality. International Studies in the Philosophy of Science, 24(3):326–329, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522415.

Ekweke:2015:IPS

[Ekw15] Evaristus Egwuatu Ekweke. An introduction to the philosophy of science. International Studies in the Philosophy of Science, 29(4):435–438, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1195147.

Elgin:2002:CRA

[Elg02] Catherine Z. Elgin. Creation as reconfiguration: Art in the advancement of science. International Studies in the Philosophy of Science, 16(1):13–25, 2002. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120118792.

Emedolu:2017:CIS

[Eme17] Christian C. Emedolu. A critical introduction to scientific realism. International Studies in the Philosophy of Science, 31(2):213–215, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2018.1424764.

Engler:2001:QFT

[Eng01] Gideon Engler. Quantum field theories and aesthetic disparity. International Studies in the Philosophy of Science, 15(1):51–63, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590020029305.

Engler:2002:EMB

[Eng02] Gideon Engler. Einstein and the most beautiful theories in physics. International Studies in the Philosophy of Science, 16(1):27–37, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120118800.

Engler:2005:EHT

[Eng05] Gideon Engler. Einstein, his theories, and his aesthetic considerations. International Studies in the Philosophy of Science, 19(1):21–30, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500051068.

Esfeld:2007:TRM

[ES07] Michael Esfeld and Christian Sachse.

Theory reduction by means of functional sub-types. International Studies in the Philosophy of Science, 21(1):1–17, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701305735.

Eser:2009:WB

[Ese09] Uta Eser. What is biodiversity? International Studies in the Philosophy of Science, 23(3):330–334, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903197823.

Esfeld:2009:MNS

[Esf09] Michael Esfeld. The modal nature of structures in ontic structural realism.

International Studies in the Philosophy of Science, 23(2):179–194, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903006917.

Earman:2014:RCA

[EV14] John Earman and Giovanni Valente. Relativistic causality in algebraic quantum field theory. International Studies in the Philosophy of Science, 28(1):1–48, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915652.

Farrell:2000:WPF

[Far00] Robert P. Farrell. Will the Popperian Feyerabend please step forward: Pluralistic, Popperian themes in the philosophy of Paul Feyerabend. International Studies in the Philosophy of Science, 14(3):257–266, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437764.

Fard:2019:PNP

[Far19] Mohammad Reza Haghighi Fard. In praise of natural philosophy: A revolution for thought and life. International Studies in the Philosophy of Science, 32(1):67–69, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1615663.

Fasce:2018:PLS

[Fas18] Angelo Fasce. Are pseudosciences like seagulls? A discriminant metacriterion facilitates the solution of the demarcation problem. International Studies in the Philosophy of Science, 32(3–4):155–175, 2018. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1767891.

Fuller:2010:LE

[FD10] Steve Fuller and Willem B. Drees.

Letter to the Editor. International Studies in the Philosophy of Science, 24(2):217–221, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.484546.

Feher:1986:MAS

[Feh87] Marta Fehér. The method of analysissynthesis and the structure of causal explanation in Newton. International Studies in the Philosophy of Science, 1(1):60–84, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573282.

Feher:1987:RMA

[Feh88] Maria Feher. The role of metaphor and analogy in the birth of the principle of least action of Maupertuis (1698–1759). International Studies in the Philosophy of Science, 2(2):175–188, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573313.

Feher:1990:RAP

[Feh90] Mart Fehér. The role accorded to the public by philosophers of science. International Studies in the Philosophy of Science, 4(3):229–240, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573363.

Feher:1998:BAA

[Feh98a] Márta Fehér. Bad arguments against a good case (Laudan’s attack on the strong programme). International Studies in the Philosophy of Science, 12(3):233–238, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573598.

Feher:1998:PAG

[Feh98b] Marta Fehér. Patterns of argumentation in Galileo’s Discorsi. International Studies in the Philosophy of Science, 12(1):17–24, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573579.

Felline:2014:QIT

[Fel14] Laura Felline. Quantum information theory and the foundations of quantum mechanics. International Studies in the Philosophy of Science, 28(3):349– 352, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2014.
                                        953348.

Fernflores:2011:BSP

[Fer11] Francisco Fernflores. Bell’s spaceships problem and the foundations of special relativity. International Studies in the Philosophy of Science, 25(4):351– 370, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
623364.

Feyerabend:2022:SMN

[Fey22] Paul K. Feyerabend. Single magnetic northpoles and southpoles and their importance for science. International Studies in the Philosophy of Science, 35(1):95–117, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.

2158430. Ten lectures delivered at the University of Vienna during the summer semester of 1947.

Firt:2021:HDO

[FHS21] Erez Firt, Meir Hemmo, and Orly Shenker. Hempel’s Dilemma: Not only for physicalism*. International Studies in the Philosophy of Science, 34(2):101–129, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2041969.

Filomeno:2016:FEO

[Fil16] Aldo Filomeno. Fundamentality, effectiveness, and objectivity of gauge symmetries. International Studies in the Philosophy of Science, 30(1):19– 37, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2016.
                                        1240459.

Fjelland:1991:TLO

[Fje91] Ragnar Fjelland. The theoryladenness of observations, the role of scientific instruments, and the Kantian a priori. International Studies in the Philosophy of Science, 5(3):269–280, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573399.

French:1999:RSA

[FL99] Steven French and James Ladyman. Reinflating the semantic approach. International Studies in the Philosophy of Science, 13(2):103–121, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573612.

Flannery:2003:AAF

[Fla03] Maura C. Flannery. Agnes Arber: form in the mind and the eye. International Studies in the Philosophy of Science, 17(3):281–300, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859032000169479.

Flores:1999:ETT

[Flo99] Francisco Flores. Einstein’s theory of theories and types of theoretical explanation. International Studies in the Philosophy of Science, 13(2):123–134, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573613.

Flores:2005:IEE

[Flo05] Francisco Flores. Interpretations of Einstein’s equation E = mc^{2}. International Studies in the Philosophy of Science, 19(3):245–260, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500462257. See re-

ply [Kra06] and response [Flo06].

Flores:2006:IER

[Flo06] Francisco Flores. On the interpretation of the equation E = mc^{2}: response to W. Krajewski: “On the interpretation of the equation E = mc^{2}: reply to F. Flores” [Int. Stud. Philos. Sci. 20 (2006), no. 2, 215–216; MR2253243]. International Studies in the Philosophy of Science, 20(2):217–218, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814480. See

[Kra06].

Foley:1986:ERS

[Fol87] Richard Foley. Epistemic rationality and scientific rationality. International Studies in the Philosophy of Science, 1(2):233–250, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573294.

Forrai:1993:MPR

[For93] Gábor Forrai. From the method of proofs and refutations to the methodology of scientific research programmes. International Studies in the Philosophy of Science, 7(2):161–175, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        /doi/abs/10.1080/02698599308573461.

Forge:1996:EQS

[For96] John Forge. Explanation and the quantum state. International Studies in the Philosophy of Science, 10(3):203–215, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573540.

Forge:2000:QQM

[For00] John Forge. Quantities in quantum mechanics. International Studies in the Philosophy of Science, 14(1):43– 56, 2000. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        /doi/abs/10.1080/026985900111891.

Franssen:2009:CMT

[Fra09] Maarten Franssen. Creations of the mind: Theories of artifacts and their representation. International Studies in the Philosophy of Science, 23(3):334–337, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903197856.

Freedman:1997:WNNa

[Fre97a] Karyn Freedman. What’s new on the net. International Studies in the Philosophy of Science, 11(2):205–206, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573564.

Freedman:1997:WNNb

[Fre97b] Karyn Freedman. What’s new on the net. International Studies in the Philosophy of Science, 11(3):303–304, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573573.

Freedman:1998:WNNa

[Fre98a] Karyn Freedman. What’s new on the net. International Studies in the Philosophy of Science, 12(1):87–89, 1998. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573583.

Freedman:1998:WNNb

[Fre98b] Karyn Freedman. What’s new on the net. International Studies in the Philosophy of Science, 12(2):193–195, 1998. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573592.

Freedman:1999:WNN

[Fre99] Karyn Freedman. What’s new on the net. International Studies in the Philosophy of Science, 13(2):193–194, 1999. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573618.

Freedman:2006:NNE

[Fre06] Karyn L. Freedman. Normative naturalism and epistemic Relativism. International Studies in the Philosophy of Science, 20(3):309–322, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600960952.

Frigg:2003:PSR

[Fri03] Roman Frigg. On the property structure of realist collapse interpretations of quantum mechanics and the socalled “counting anomaly”. International Studies in the Philosophy of Science, 17(1):43–57, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590305232.

Friederich:2017:FTO

[Fri17] Simon Friederich. Fine-tuning as old evidence, double counting, and the multiverse. International Studies in the Philosophy of Science, 31(4):363– 377, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2019.
                                        1565214.

Fuchs:1993:AET

[Fuc93] Stephan Fuchs. Against essentialism in theories of rational action: a reply to Raymond Boudon. International Studies in the Philosophy of Science, 7(1):37–39, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/

10.1080/02698599308573444. See

[Bou93b] and reply [Bou93a].

Fuller:2022:DFK

[Ful22] Gareth Fuller. A defence of functional kinds: Multiple realisability and explanatory counterfactuals. International Studies in the Philosophy of Science, 35(2):119–133, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2144436.

Futch:2002:LNT

[Fut02] Michael J. Futch. Leibniz’s nontensed theory of time. International Studies in the Philosophy of Science, 16(2):125–139, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145070.

Gao:2015:AEE

[Gao15] Jie Gao. Advances in experimental epistemology. International Studies in the Philosophy of Science, 29(1):101– 105, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1071554.

Gao:2017:USP

[Gao17] Shan Gao. Understanding scientific progress: Aim-oriented empiricism. International Studies in the Philosophy of Science, 31(4):435–438, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565211.

Garnham:1994:MM

[Gar94] Alan Garnham. March of the models. International Studies in the Philosophy of Science, 8(1):37–39, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573477.

Gattei:2000:OJW

[Gat00] Stefano Gattei. Obituary: John W. N. Watkins, 1924–1999. International Studies in the Philosophy of Science, 14(1):77–81, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/026985900111918.

Gauthier:2005:HWM

[Gau05] Yvon Gauthier. Hermann Weyl on Minkowskian space–time and Riemannian geometry. International Studies in the Philosophy of Science, 19(3):261–269, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500462265.

Gelfert:2003:MSU

[Gel03] Axel Gelfert. Manipulative success and the unreal. International Studies in the Philosophy of Science, 17(3):245–263, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859032000169451.

Ghassib:2012:TKI

[Gha12] Hisham Ghassib. A theory of the knowledge industry. International Studies in the Philosophy of Science, 26(4):447–456, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.748499.

Ghassib:2017:RME

[Gha17] Hisham Ghassib. Reason and method in Einstein’s relativity. International Studies in the Philosophy of Science, 31(4):331–342, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565204.

Ghins:2003:TKE

[Ghi03] Michel Ghins. Thomas Kuhn on the existence of the world. International Studies in the Philosophy of Science, 17(3):265–279, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859032000169460.

Gijsbers:2009:DAS

[Gij09] Victor Gijsbers. Depth: An account of scientific explanation. International Studies in the Philosophy of Science, 23(2):225–228, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903007212.

Gillett:1991:LSE

[Gil91] Grant Gillett. Language, social ecology and experience. International Studies in the Philosophy of Science, 5(3):195–203, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573393.

Gillies:1992:CSD

[Gil92] Donald A. Gillies. Comments on ‘Scientific Discovery as Problem Solving’ by Herbert A. Simon. International Studies in the Philosophy of Science, 6(1):29–31, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573407. See

[Sim92a, Sim92b].

Ginev:2005:APP

[Gin05] Dimitri Ginev. Against the politics of postmodern philosophy of science. International Studies in the Philosophy of Science, 19(2):191–208, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500249522.

Ginev:2012:TAH

[Gin12] Dimitri Ginev. Two accounts of the hermeneutic fore-structure of scientific research. International Studies in the Philosophy of Science, 26(4):423–445, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
748498.

Girotto:1994:MTI

[Gir94] Vittorio Girotto. Is the model theory of induction also a theory of inductive reasoning? International Studies in the Philosophy of Science, 8(1):41–43, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573478.

Girel:2023:WEP

[Gir23] Mathias Girel. What exactly is presupposed by agnotology? The challenge of intentions. International Studies in the Philosophy of Science, 36(3):229–246, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2257111.

Godman:2014:SEN

[God14] Marion Godman. Scientific enquiry and natural kinds: From planets to mallards. International Studies in the Philosophy of Science, 28(3):343–346, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
953346.

Goldstick:1993:LNP

[Gol93] D. Goldstick. Laws of nature and physical existents. International Studies in the Philosophy of Science, 7(3):255–265, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573469.

Golosz:1999:FAS

[Goł99] Jerzy Gołosz. On Field’s argument for substantivalism. International Studies in the Philosophy of Science, 13(1):5–16, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573604.

Goldstick:2020:FJP

[Gol20] Daniel Goldstick. De Facto and De Jure in the practice of induction. International Studies in the Philosophy of Science, 33(2):105–116, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2020.1866478.

Gombay:1991:PDL

[Gom91] A. Gombay. Psychic determinism: Leibniz and Freud. International Studies in the Philosophy of Science, 5(3):205–213, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573394.

Gonzalez:1995:RCP

[Gon95] Wenceslao J. González. Reichenbach’s concept of prediction. International Studies in the Philosophy of Science, 9(1):37–58, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573505.

Gontier:2012:SAE

[Gon12] Nathalie Gontier. Selectionist approaches in evolutionary linguistics: An epistemological analysis. International Studies in the Philosophy of Science, 26(1):67–95, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.653114.

Gonzalez:2015:CEP

[Gon15] Wenceslao J. Gonzalez. From the characterization of ‘European philosophy of science’ to the case of philosophy of the social sciences. International Studies in the Philosophy of Science, 29(2):167–188, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119418.

Gooding:1990:TOE

[Goo90] David Gooding. Theory and observation: The experimental nexus. International Studies in the Philosophy of Science, 4(2):131–148, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573353.

Gottlob:2000:NAP

[Got00] Rainer Gottlob. New aspects of the probabilistic evaluation of hypotheses and experience. International Studies in the Philosophy of Science, 14(2):147–163, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590050045737.

Gouk:1986:NMM

[Gou87] Penelope Gouk. Newton and music: From the microcosm to the macrocosm. International Studies in the Philosophy of Science, 1(1):36–59, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573281.

Grammenos:2018:MEP

[Gra18] Theophanes Grammenos. Metaphysical experiments — physics and the invention of the universe. International Studies in the Philosophy of Science, 32(3–4):232–234, 2018. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2020.1767893.

Green:1994:IRS

[Gre94] David W. Green. Induction: Representation, strategy and argument. International Studies in the Philosophy of Science, 8(1):45–50, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573479.

Grobler:1991:VFM

[Gro91] Adam Grobler. Van Fraassen’s metaphysical move. International Studies in the Philosophy of Science, 5(1):21–34, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573375.

Gurova:2022:UTT

[Gur22] Lilia Gurova. The uses of truth: Is there room for reconciliation of factivist and non-factivist accounts of scientific understanding? International Studies in the Philosophy of Science, 35(3–4):211–221, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2022.2084329.

Grune-Yanoff:2011:ICM

[GY11] Till Grüne-Yanoff. Isolation is not characteristic of models. International Studies in the Philosophy of Science, 25(2):119–137, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574854.

Hacker:1995:HTP

[Hac95] P. M. S. Hacker. Helmholtz’s theory of perception: An investigation into its conceptual framework. International Studies in the Philosophy of Science, 9(3):199–214, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573519.

Hansson:2009:CGK

[Han09] Sven Ove Hansson. Cutting the Gordian knot of demarcation. International Studies in the Philosophy of Science, 23(3):237–243, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903196007.

Hansson:2020:DDD

[Han20] Sven Ove Hansson. Disciplines, doctrines, and deviant science. International Studies in the Philosophy of Science, 33(1):43–52, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1831258.

Harre:1987:WMA

[Har88] R. Harre. Where models and analogies really count. International Studies in the Philosophy of Science, 2 (2):118–133, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573310.

Harre:1990:TAS

[Har90] R. Harré. Tracks and affordances: The sources of a physical ontology. International Studies in the Philosophy of Science, 4(2):149–158, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573354.

Harre:1991:SRI

[Har91] Rom Harré. Some reflections on the individuation of events. International Studies in the Philosophy of Science, 5(1):49–63, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573377.

Harre:1995:ROP

[Har95] Rom Harré. Realism and an ontology of powerful particulars. International Studies in the Philosophy of Science, 9(3):285–300, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573524.

Harre:1997:FJR

[Har97a] Rom Harré, editor. Festschrift for J. R. Lucas. Carfax Publishing, Taylor & Francis Ltd., Basingstoke, UK, 1997. ISSN 0269-8595 (print), 1469-9281 (electronic). 1–108 pp. Internat. Stud. Philos. Sci. 11 (1997), no. 1.

Harre:1997:I

[Har97b] Rom Harré. Introduction. International Studies in the Philosophy of Science, 11(1):5, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573546.

Harre:1997:ODS

[Har97c] Rom Harré. The ontological duality of space–time variables. International Studies in the Philosophy of Science, 11(1):83–96, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573552. Festschrift

for J. R. Lucas.

Harbecke:2010:MCN

[Har10] Jens Harbecke. Mechanistic constitution in neurobiological explanations. International Studies in the Philosophy of Science, 24(3):267–285, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522409.

Hawley:2010:SO

[Haw10] Katherine Hawley. The structure of objects. International Studies in the Philosophy of Science, 24(3):336–339, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
522418.

Heidelberger:2001:OLT

[Hei01] Michael Heidelberger. Origins of the logical theory of probability: Von Kries, Wittgenstein, Waismann. International Studies in the Philosophy of Science, 15(2):177–188, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120059040.

Heidelberger:2006:AMF

[Hei06] Michael Heidelberger. Applying models in fluid dynamics. International Studies in the Philosophy of Science, 20(1):49–67, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600641016.

Hesse:1992:CHS

[Hes92] Mary Hesse. Comment on Herbert Simon, “Scientific discovery as problem solving”. International Studies in the Philosophy of Science, 6(1):33–34, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573408. See

[Sim92a].

Harre:1992:BR

[HF92] Rom Harré and Marta Fehér. Book reviews. International Studies in the Philosophy of Science, 6(3):249–253, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        /doi/abs/10.1080/02698599208573435.

Harbecke:2024:BIC

[HGS24] Jens Harbecke, Jonas Grunau, and Philip Samanek. Are the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) applicable in determining the optimal fit and simplicity of mechanistic models? International Studies in the Philosophy of Science, 37(1–2):17–36, 2024. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2024.
2304487.

Hibbert:2016:WIS

[Hib16] Ruth Hibbert. What is an immature science? International Studies in the Philosophy of Science, 30(1):1– 17, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2016.
1240433.

Hoffmann-Kolss:2014:IHL

[HK14] Vera Hoffmann-Kolss. Interventionism and higher-level causation. International Studies in the Philosophy of Science, 28(1):49–64, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915653.

Henderson:1999:BR

[HMO^{+}99] P. Henderson, Matthew MacKenzie, Dawn Ogden, Jutta Rockmann, Spas Spassov, and John Taylor. Book reviews. International Studies in the Philosophy of Science, 13(1):83–96, 1999. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573609.

Hochstein:2013:IME

[Hoc13] Eric Hochstein. Intentional models as essential scientific tools. International Studies in the Philosophy of Science, 27(2):199–217, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.813251.

Hodgson:1997:RQM

[Hod97] Peter Hodgson. Realism and quantum mechanics. International Studies in the Philosophy of Science, 11(1):53–65, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573550. Festschrift

for J. R. Lucas.

Hoefer:2000:KHE

[Hoe00] Carl Hoefer. Kant’s hands and Earman’s pions: Chirality arguments for substantival space. International Studies in the Philosophy of Science, 14(3):237–256, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/026985900437755.

Hohwy:2003:CEP

[Hoh03] Jakob Hohwy. Capacities, explanation and the possibility of disunity. International Studies in the Philosophy of Science, 17(2):179–190, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859031000160630.

Holst:2009:WPS

[Hol09] Cathrine Holst. What is philosophy of social science? International Studies in the Philosophy of Science, 23(3):313–321, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903197781.

Hon:1991:CME

[Hon91] Giora Hon. Can the monster errour be slain? International Studies in the Philosophy of Science, 5(3):257–268, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573398.

Hope:2013:STD

[Hop13] Kristin Lofthus Hope. Science transformed? Debating claims of an epochal break. International Studies in the Philosophy of Science, 27(2):228–231, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
813252.

Howson:2009:SNT

[How09] Colin Howson. Sorites is no threat to Modus Ponens: a reply to Kochan. International Studies in the Philosophy of Science, 23(2):209–212, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903007188. See

[Koc08, Koc09].

Howson:2011:NAH

[How11] Colin Howson. No answer to Hume [discussion of MR2680848]. International Studies in the Philosophy of Science, 25(3):279–284, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.605249. See

replies [Ste11, BTB16].

Howson:2016:HPH

[How16] Colin Howson. How pseudohypotheses defeat a non-Bayesian theory of evidence: Reply to Bandyopadhyay, Taper, and Brittan. International Studies in the Philosophy of Science, 30(3):299–306, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1316111.

Howson:2018:PCS

[How18] Colin Howson. The primacy of the classical? Saul Kripke meets Niels Bohr. International Studies in the Philosophy of Science, 32(3–4):141–153, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2020.
                                        1767890.

Haufe:2009:WNM

[HS09] Chris Haufe and Matthew H. Slater. Where no mind has gone before: Exploring laws in distant and lonely worlds. International Studies in the Philosophy of Science, 23(3):265–276, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590903196049.

Hemmo:2015:LE

[HS15] Meir Hemmo and Orly Shenker. Letter to the Editor. International Studies in the Philosophy of Science, 29(1): 91–93, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1079075.

Huttemanna:2000:EP

[HT00] Andreas Hüttemanna and Orestis Terzidis. Emergence in physics. International Studies in the Philosophy of Science, 14(3):267–281, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437773.

Hueck:2025:BRR

[Hue25] Christoph J. Hueck. Book review: Remapping Biology with Goethe, Schelling, and Herder. Romanticising Evolution, by Gregory Rupik, London, Routledge, 2024, 184 pp., 4 B/W Illustrations, £135.00, ISBN 978-1-032-58279-5, https://doi.org/10.4324/9781003441809. International Studies in the Philosophy of Science, 38(1):79–81, 2025. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2024.2426193.

Huggett:1999:WMS

[Hug99] Nick Huggett. Why manifold substantivalism is probably not a consequence of classical mechanics. International Studies in the Philosophy of Science, 13(1):17–34, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573605.

Hulton:2014:NCH

[Hul14] Ana Hulton. Natural categories and human kinds: Classification in the natural and social sciences. International Studies in the Philosophy of Science, 28(1):102–105, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915654.

Hunt:1990:LPU

[Hun90] G. M. K. Hunt. Laws: Projectability and uniformity. International Studies in the Philosophy of Science, 4(3):241–246, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573364.

Hunt:1994:CTI

[Hun94] Earl Hunt. Comments on a theory of induction. International Studies in the Philosophy of Science, 8(1):51–55, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573480.

Hung:2001:KPR

[Hun01] Edwin H.-C. Hung. Kuhnian paradigms as representational spaces: New perspectives on the problems of incommensurability, scientific explanation, and physical necessity. International Studies in the Philosophy of Science, 15(3):275–292, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080217.

Hunt:2011:TSI

[Hun11] Shelby D. Hunt. Theory status, inductive realism, and approximate truth: No miracles, no charades. International Studies in the Philosophy of Science, 25(2):159–178, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574858.

Iliescu:1991:BR

[Ili91] Adrian-Paul Iliescu. Book review. International Studies in the Philosophy of Science, 5(1):89–92, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573380.

Illari:2011:MED

[Ill11] Phyllis McKay Illari. Mechanistic evidence: Disambiguating the Russo–Williamson thesis. International Studies in the Philosophy of Science, 25(2):139–157, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574856.

Inan:2014:RDT

[Ina14] Ilhan Inan. Rigid designation and theoretical identities. International Studies in the Philosophy of Science, 28(2):217–220, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.932531.

Irvine:1993:HBP

[Irv93] A. D. Irvine. How Braess’ paradox solves Newcomb’s problem. International Studies in the Philosophy of Science, 7(2):141–160, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573460. See re-

ply [Mar96].

Irvine:1997:PSW

[Irv97] A. D. Irvine. Philosophy of science in Western Canada. International Studies in the Philosophy of Science, 11(3):289–301, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573572.

IslasMondragon:2017:IRA

[Isl17] Damián Islas Mondragón. The inference rule of addition and the semantic view of scientific progress: Reply to Mizrahi. International Studies in the Philosophy of Science, 31(4):421– 425, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.

1565212. See [Miz17].

Ivanova:2011:FRP

[Iva11] Milena Ivanova. Friedman’s relativised A Priori and structural realism: In search of compatibility. International Studies in the Philosophy of Science, 25(1):23–37, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.552416.

Ivanova:2014:ESS

[Iva14] Milena Ivanova. Explaining science’s success: Understanding how scientific knowledge works. International Studies in the Philosophy of Science, 28(1):105–108, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915655.

Jacquette:1993:PTP

[Jac93] Dale Jacquette. Pollock on token physicalism, agent materialism and strong artificial intelligence. International Studies in the Philosophy of Science, 7(2):127–140, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573459.

Jacobs:2001:LPS

[Jac01] Struan Jacobs. Limits to problem solving in science. International Studies in the Philosophy of Science, 15(3):231–242, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080172.

Jacob:2011:CAR

[Jac11] Merle Jacob. The commodification of academic research: Science and the modern university. International Studies in the Philosophy of Science, 25(4):423–426, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.623373.

Jaeger:2002:HSB

[Jae02] Lydia Jaeger. Humean supervenience and best-system laws. International Studies in the Philosophy of Science, 16(2):141–155, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590220145089.

Jaeger:2017:ROA

[Jae17] Lydia Jaeger. Rethinking order: After the laws of nature. International Studies in the Philosophy of Science, 31(2):215–218, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2018.1424765.

Janich:1986:NFP

[Jan87] Peter Janich. The normative foundation of physics. International Studies in the Philosophy of Science, 1 (2):251–261, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573295.

Janich:1988:DBN

[Jan89] Peter Janich. Does biology need a relativistic revision? International Studies in the Philosophy of Science, 3(2):190–198, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598908573335.

Jansson:2015:NST

[Jan15] Lina Jansson. The nature of scientific thinking: On interpretation, explanation, and understanding. International Studies in the Philosophy of Science, 29(2):218–221, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119425.

Jimenez-Buedo:2015:LDG

[JB15] María Jiménez-Buedo. The last dictator game? Dominance, reactivity, and the methodological artefact in experimental economics. International Studies in the Philosophy of Science, 29(3):295–310, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1179042.

Jones:1995:BR

[JBM^{+}95] W. Jones, James Robert Brown, W. J. Mander, Władysław Krajewski, John M. Preston, Stathis Psillos, Katherine Hawley, and John Taylor. Book reviews. International Studies in the Philosophy of Science, 9(2):157–188, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573515.

Jebeile:2017:CSE

[Jeb17] Julie Jebeile. Computer simulation, experiment, and novelty. International Studies in the Philosophy of Science, 31(4):379–395, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2019.1565205.

Jebeile:2015:EMR

[JK15] Julie Jebeile and Ashley Graham Kennedy. Explaining with models: The role of idealizations. International Studies in the Philosophy of Science, 29(4):383–392, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1195143.

Johnson-Laird:1994:MTI

[JL94a] Philip N. Johnson-Laird. A model theory of induction. International Studies in the Philosophy of Science, 8(1):5–29, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573474.

Johnson-Laird:1994:RCM

[JL94b] Philip N. Johnson-Laird. Reply to the commentators on a model theory of induction. International Studies in the Philosophy of Science, 8(1):73–96, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573485.

Johnson-Laird:1992:SHS

[JLL92] Philip Johnson-Laird and Paolo Legrenzi. Science as heuristic search: Some comments on Simon’s theory. International Studies in the Philosophy of Science, 6(1):35–39, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573409.

Johansen:2016:CSP

[JM16] Mikkel Willum Johansen and Morten Misfeldt. Computers as a source of a posteriori knowledge in mathematics. International Studies in the Philosophy of Science, 30(2):111–127, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265862.

Jodoin:2024:CIS

[Jod24] Laurent Jodoin. A conceptualisation of irreversibility for sustainable development. International Studies in the Philosophy of Science, 37(3):109–126, 2024. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2024.
2409057.

Jones:2000:UET

[Jon00] Ward E. Jones. Underdetermination and the explanation of theory-acceptance: a response to Samir Okasha. International Studies in the Philosophy of Science, 14(3):299–304, 2000. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437791.  See

[Oka00b].

Joyce:2013:CDH

[Joy13] Kelly Joyce. Chagas disease: History of a continent’s scourge. International Studies in the Philosophy of Science, 27(4):459–461, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868180.

Jones:1994:BR

[JSJ^{+}94] W. Jones, Helen Steward, Fiona Jenkins, Lee Mcintyre, Mary Hesse, C. W. Kilmister, Richard J. Hall, and Timothy Williamson. Book reviews. International Studies in the Philosophy of Science, 8(2):147–168, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573491.

Jraissati:2012:CCC

[JWDD12] Yasmina Jraissati, Elley Wakui, Lieven Decock, and Igor Douven. Constraints on colour category formation. International Studies in the Philosophy of Science, 26(2):171–196, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.703479.

Kaiser:2015:PM

[Kai15] Marie I. Kaiser. Philosophy of microbiology. International Studies in the Philosophy of Science, 29(2):224– 228, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1119427.

Kapusta:2010:PST

[Kap10] Paweł Kapusta. Philosophy of science for theologians: An introduction. International Studies in the Philosophy of Science, 24(4):443–446, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.543356.

Kardaun:2015:JQS

[Kar15] Maria Kardaun. Jung and the question of science. International Studies in the Philosophy of Science, 29(1):108– 111, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1071556.

Kastner:2004:SBN

[Kas04] Ruth E. Kastner. Shutters, boxes, but no paradoxes: Time symmetry puzzles in quantum theory. International Studies in the Philosophy of Science, 18(1):89–94, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590412331289279.

Kastner:2015:ECD

[Käs15] Lena Kästner. Epistemic cognition and development: The psychology of justification and truth. International Studies in the Philosophy of Science, 29(4):444–447, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1195150.

Kastner:2019:NMP

[Käs19] Lena Kästner. The new mechanical philosophy. International Studies in the Philosophy of Science, 32(1):69– 72, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2019.
                                        1615664.

Katzav:2013:DCP

[Kat13] Joel Katzav. Dispositions, causes, persistence as is, and general relativity. International Studies in the Philosophy of Science, 27(1):41–57, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.783974.

Kazumi:2014:DCC

[Kaz14] Inoue Kazumi. Dialectical contradictions and classical formal logic. International Studies in the Philosophy of Science, 28(2):113–132, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.932526.

Kleinhans:2005:TIE

[KBdR05] Maarten G. Kleinhans, Chris J. J. Buskes, and Henk W. de Regt. Terra Incognita: explanation and reduction in Earth science. International Studies in the Philosophy of Science, 19(3):289–317, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500462356.

Keet:2012:AIB

[Kee12] C. Maria Keet. Artificial intelligence: The basics. International Studies in the Philosophy of Science, 26(3):351– 354, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/

doi/full/10.1080/02698595.2012.
731737.

Kendig:2011:DPB

[Ken11] Catherine Kendig. Debates in philosophy of biology: One long argument, or many? International Studies in the Philosophy of Science, 25(1): 73–81, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
552419.

Kennedy:2013:MSR

[Ken13] Ashley Graham Kennedy. Mathematics and scientific representation. International Studies in the Philosophy of Science, 27(1):95–98, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.783975.

Kent:2022:PFD

[Ken22] Rory Kent. Paul Feyerabend and the dialectical character of quantum mechanics: a lesson in philosophical dadaism. International Studies in the Philosophy of Science, 35(1):51– 67, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2075687.

Kuby:2022:FQT

[KF22] Daniel Kuby and Patrick Fraser. Feyerabend on the quantum theory of measurement: a reassessment. International Studies in the Philosophy of Science, 35(1):23–49, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2022.2067811.

Khalidi:2005:AFR

[Kha05] Muhammad Ali Khalidi. Against functional reductionism in cognitive science. International Studies in the Philosophy of Science, 19(3):319–333, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590500462448.

Khalifa:2010:DPB

[Kha10] Kareem Khalifa. Default privilege and bad lots: Underconsideration and explanatory inference. International Studies in the Philosophy of Science, 24(1):91–105, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903467135.

Kidd:2013:HCI

[Kid13] Ian James Kidd. Historical contingency and the impact of scientific imperialism. International Studies in the Philosophy of Science, 27(3):315–324, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2013.
                                        825494.

Kidd:2016:WFP

[Kid16] Ian James Kidd. Was Feyerabend a postmodernist? International Studies in the Philosophy of Science, 30(1): 55–68, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/

doi/full/10.1080/02698595.2016.
1240463.

Kindi:2011:CSR

[Kin11] Vasso Kindi. The challenge of scientific revolutions: Van Fraassen’s and Friedman’s responses. International Studies in the Philosophy of Science, 25(4):327–349, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.623363.

Kindi:2017:RRB

[Kin17] Vasso Kindi. A reconsideration of the relation between Kuhnian incommensurability and translation. International Studies in the Philosophy of Science, 31(4):397–414, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565206.

Kindi:2018:ER

[Kin18] Vasso Kindi. Editorial report 2019.

International Studies in the Philosophy of Science, 32(3–4):235–236, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1814040.

Kindi:2019:E

[Kin19] Vasso Kindi. Editorial. International Studies in the Philosophy of Science, 32(1):1, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1650560.

Kindi:2020:ER

[Kin20] Vasso Kindi. Editorial report 2020. International Studies in the Philosophy of Science, 33(4):259–260, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2020.1986773.

Kistler:1999:MRR

[Kis99] Max Kistler. Multiple realization, reduction and mental properties. International Studies in the Philosophy of Science, 13(2):135–149, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573614.

Klein:2001:CRH

[Kle01] Carsten Klein. Conventionalism and realism in Hans Reichenbach’s philosophy of geometry. International Studies in the Philosophy of Science, 15(3):243–251, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120080181.

Kliemt:1993:PWR

[Kli93] Hartmut Kliemt. Perfect and workable rationality: a comment on Raymond Boudon’s paper. International Studies in the Philosophy of Science, 7(1):41–43, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573445. See

[Bou93b] and reply [Bou93a].

Knowles:2002:WRW

[Kno02] Jonathan Knowles. What’s really wrong with Laudan’s normative naturalism. International Studies in the Philosophy of Science, 16(2):171–186, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145106.

Knox:2011:BRA

[Kno11] Eleanor Knox. Book review: The Ashgate Companion to Contemporary Philosophy of Physics. International Studies in the Philosophy of Science, 25(2):199–202, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574868.

Knuuttila:2010:SMP

[Knu10] Tarja Knuuttila. Scientific models in philosophy of science. International Studies in the Philosophy of Science, 24(4):437–440, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.543353.

Kochan:2008:RRS

[Koc08] Jeff Kochan. Realism, reliabilism, and the ‘strong programme’ in the sociology of scientific knowledge. International Studies in the Philosophy of Science, 22(1):21–38, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802280878. See re-

ply [How09] and response [Koc09].

Kochan:2009:EMR

[Koc09] Jeff Kochan. The exception makes the rule: Reply to Howson. International Studies in the Philosophy of Science, 23(2):213–216, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903007196. See

[Koc08, How09].

Kochan:2012:SSN

[Koc12] Jeff Kochan. Science studies as naturalized philosophy. International Studies in the Philosophy of Science, 26(1):121–124, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.653115.

Kosso:2000:FAS

[Kos00] Peter Kosso. Fundamental and accidental symmetries. International Studies in the Philosophy of Science, 14(2):109–121, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590050045700.

Kosso:2002:OBS

[Kos02] Peter Kosso. The omniscienter: Beauty and scientific understanding. International Studies in the Philosophy of Science, 16(1):39–48, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120118819.

Kouw:2015:MBR

[Kou15] Matthijs Kouw. Model-based reasoning in science and technology: Theoretical and cognitive issues. International Studies in the Philosophy of Science, 29(1):105–108, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1071555.

Kourany:2023:RGT

[Kou23] Janet A. Kourany. Race and gender: Toward a proper pattern of knowledge and ignorance in research. International Studies in the Philosophy of Science, 36(3):173–192, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2280809.

Kowalenko:2014:CPL

[Kow14] Robert Kowalenko. Ceteris Paribus laws: a naturalistic account. International Studies in the Philosophy of Science, 28(2):133–155, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.932527.

Kowalenko:2017:MCP

[Kow17] Robert Kowalenko. Manipulationism, Ceteris Paribus laws, and the bugbear of background knowledge. International Studies in the Philosophy of Science, 31(3):261–283, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463693.

Krajewski:1992:QOK

[Kra92] Władysław Krajewski. Questions of the objects of knowledge and types of realism. International Studies in the Philosophy of Science, 6(3):205–213, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573431.

Krajewski:1997:IOM

[Kra97] Władysław Krajewski. Ideal objects as models in science. International Studies in the Philosophy of Science, 11(2):185–190, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573562.

Krajewski:2006:IER

[Kra06] Władysław Krajewski. On the interpretation of the equation E = mc^{2}: reply to F. Flores: “Interpretations of Einstein’s equation E = mc^{2}” [Int. Stud. Philos. Sci. 19 (2005), no. 3, 245–260; MR2201925]. International Studies in the Philosophy of Science, 20(2):215–216, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814449. See

[Flo05, Flo06].

Krauss:2021:AOV

[Kra21] Alexander Krauss. Assessing the overall validity of randomised controlled trials. International Studies in the Philosophy of Science, 34(3):159– 182, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2021.
2002676.

Kremakova:2017:AAC

[Kre17] Milena Kremakova. Accelerating academia: The changing structure of academic time. International Studies in the Philosophy of Science, 31(2):221–224, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1424766.

Krohs:2008:HDC

[Kro08] Ulrich Krohs. How digital computer simulations explain real-world processes. International Studies in the Philosophy of Science, 22(3):277–292, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590802567324.

Krause:2007:BR

[KSA07] Décio Krause, Eric Schliesser, and Hanne Andersen. Book reviews. International Studies in the Philosophy of Science, 21(3):345–357, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701589619.

Kujundzic:1998:RVT

[Kuj98] Nebojsa Kujundzic. The role of variation in thought experiments. International Studies in the Philosophy of Science, 12(3):239–243, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573599.

Kukla:1993:EB

[Kuk93] Andre Kukla. Epistemic boundedness. International Studies in the Philosophy of Science, 7(2):121–126, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573458.

Kukla:1995:TLI

[Kuk95] André Kukla. Is there a logic of incoherence? International Studies in the Philosophy of Science, 9(1):59–71, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573506.

Kuorikoski:2009:TCM

[Kuo09] Jaakko Kuorikoski. Two concepts of mechanism: Componential causal system and abstract form of interaction. International Studies in the Philosophy of Science, 23(2):143–160, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903006875.

Kutschmann:1986:SIS

[Kut87] Werner Kutschmann. Scientific instruments and the senses: Towards an anthropological historiography of the natural sciences. International Studies in the Philosophy of Science, 1(1):106–123, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598608573284.

Kutach:2010:CBM

[Kut10] Douglas Kutach. A connection between Minkowski and Galilean space– times in quantum mechanics. International Studies in the Philosophy of Science, 24(1):15–29, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903467093.

Kutach:2011:BI

[Kut11] Douglas Kutach. Backtracking influence. International Studies in the Philosophy of Science, 25(1):55–71, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.552418.

Kuukkanen:2009:RK

[Kuu09] Jouni-Matti Kuukkanen. Rereading Kuhn. International Studies in the Philosophy of Science, 23(2):217–224, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590903007204.

Kuukkanen:2012:AOS

[Kuu12] Jouni-Matti Kuukkanen. Autonomy and objectivity of science. International Studies in the Philosophy of Science, 26(3):309–334, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.731733.

Kuukkanen:2017:LRR

[Kuu17] Jouni-Matti Kuukkanen. Lakatosian rational reconstruction updated. International Studies in the Philosophy of Science, 31(1):83–102, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1370930.

Kvasz:2016:RMT

[Kva16] Ladislav Kvasz. Revisiting the mathematisation thesis: Galileo, Descartes, Newton, and the language of nature. International Studies in the Philosophy of Science, 30(4):399–406, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331978.

Lam:2012:GP

[Lam12] Vincent Lam. Geometric possibility. International Studies in the Philosophy of Science, 26(2):226–229, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.703483.

Lam:2013:ESQ

[Lam13] Vincent Lam. The entanglement structure of quantum field systems. International Studies in the Philosophy of Science, 27(1):59–72, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.783976.

Langfeldt:2011:HPT

[Lan11] Liv Langfeldt. How professors think: Inside the curious world of academic judgment. International Studies in the Philosophy of Science, 25(1):92– 95, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
552423.

Landes:2014:TIP

[Lan14] Jürgen Landes. Tychomancy: Inferring probability from causal structure. International Studies in the Philosophy of Science, 28(4):446–448, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979675.

Lange:2023:ECJ

[Lan23] Marc Lange. Explanations by constraint: Not just in physics. International Studies in the Philosophy of Science, 36(4):265–277, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2298085.

Larvor:1994:HME

[Lar94] Brendan Larvor. History, methodology and early algebra. International Studies in the Philosophy of Science, 8(2):113–124, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573488.

Laraudogoitia:2001:ICG

[Lar01] Jon Pérez Laraudogoitia. Indeterminism, classical gravitation and non-collision singularities. International Studies in the Philosophy of Science, 15(3):269–274, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080208.

Laraudogoitia:2006:GIC

[Lar06] Jon Pérez Laraudogoitia. Global interaction in classical mechanics. International Studies in the Philosophy of Science, 20(2):173–183, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600814357.

Laudisa:2015:LD

[Lau15] Federico Laudisa. Laws are not descriptions. International Studies in the Philosophy of Science, 29(3):251– 270, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1179038.

Lawton:1993:ASC

[Law93] Philip Lawton. Art, science, and the clear blue sky. International Studies in the Philosophy of Science, 7(2):107–119, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573457.

Lazarovici:2017:MWF

[Laz17] Dustin Lazarovici. The meaning of the wave function: In search of the ontology of quantum mechanics. International Studies in the Philosophy of Science, 31(3):321–324, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2018.1463694.

Leng:1998:BR

[LB98] Mary Leng and James Robert Brown. Book reviews. International Studies in the Philosophy of Science, 12(2):197–203, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573593.

Leech:2014:P

[Lee14] Jessica Leech. Properties. International Studies in the Philosophy of Science, 28(4):439–442, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979673.

Legrenzi:1994:I

[Leg94] Paolo Legrenzi. Introduction. International Studies in the Philosophy of Science, 8(1):3, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573473.

Lelas:1987:RA

[Lel88] Srdjan Lelas. Review article. International Studies in the Philosophy of Science, 2(2):237–242, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573317.

Lelas:1988:ENR

[Lel89] Srdan Lelas. Evolutionary naturalist realism: Can this blend be coherent? International Studies in the Philosophy of Science, 3(2):136–156, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573332.

Lennox:2001:AUD

[Len01] James G. Lennox. Aristotle on the unity and disunity of science. International Studies in the Philosophy of Science, 15(2):133–144, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120059004.

Leng:2005:PAP

[Len05] Mary Leng. Platonism and anti-Platonism: Why worry? International Studies in the Philosophy of Science, 19(1):65–84, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500051159.

Leonelli:2011:ECT

[Leo11] Sabina Leonelli. An epistemology of the concrete: Twentieth-century histories of life. International Studies in the Philosophy of Science, 25(4):420– 422, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
623372.

Leonelli:2012:CTD

[Leo12] Sabina Leonelli. Classificatory theory in data-intensive science: The case of open biomedical ontologies. International Studies in the Philosophy of Science, 26(1):47–65, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.653119.

Leplin:1987:RET

[Lep88] Jarrett Leplin. The role of experiment in theory construction. International Studies in the Philosophy of Science, 2(1):72–83, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573303.

Leroux:2001:PTF

[Ler01] Jean Leroux. “Picture theories”

as forerunners of the semantic approach to scientific theories. International Studies in the Philosophy of Science, 15(2):189–197, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120059059.

Leslie:1996:DEM

[Les96] John Leslie. A difficulty for Everett’s many-worlds theory. International Studies in the Philosophy of Science, 10(3):239–246, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573542.

Letrud:2019:GKD

[Let19] Kåre Letrud. The Gordian knot of demarcation: Tying up some loose ends. International Studies in the Philosophy of Science, 32(1):3–11, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1618031.

Letrud:2022:ISD

[Let22] Kåre Letrud. Incorrigible science and doctrinal pseudoscience. International Studies in the Philosophy of Science, 35(3–4):269–278, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2023.2218264.

Lewis:2003:FSD

[Lew03] Peter J. Lewis. Four strategies for dealing with the counting anomaly in spontaneous collapse theories of quantum mechanics. International Studies in the Philosophy of Science, 17(2):137–142, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859031000160603.

Lewowicz:2016:SAP

[Lew16] Lucía Lewowicz. Science after the practice turn in the philosophy, history, and social studies of science. International Studies in the Philosophy of Science, 30(2):190–193, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2016.1265863.

Lindenberg:1993:RBF

[Lin93] Siegwart Lindenberg. The rationality of beliefs at the frontier of social science: Comments on Boudon’s paper. International Studies in the Philosophy of Science, 7(1):51–53, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/

10.1080/02698599308573447. See

[Bou93b] and reply [Bou93a].

Ling:1995:INP

[Lin95] Charles X. Ling. Introducing new predicates to model scientific revolution. International Studies in the Philosophy of Science, 9(1):19–36, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573504.

Lindholm:2022:SPS

[Lin22] Juho Lindholm. Scientific practices as social knowledge. International Studies in the Philosophy of Science, 35(3–4):223–242, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2196930.

Lipton:1990:PP

[Lip90] Peter Lipton. Prediction and prejudice. International Studies in the Philosophy of Science, 4(1):51–65, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573345.

Liu:1997:MTS

[Liu97] Chuang Liu. Models and theories.

I. The semantic view revisited. International Studies in the Philosophy of Science, 11(2):147–164, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573560.

Liu:1998:MTI

[Liu98] Chuang Liu. Models and theories. II. Issues and applications. International Studies in the Philosophy of Science, 12(2):111–128, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573587.

Liu:2003:GGU

[Liu03] Chuang Liu. Gauge gravity and the unification of natural forces. International Studies in the Philosophy of Science, 17(2):143–159, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859031000160612.

Liu:2015:RIC

[Liu15] Chuang Liu. Re-inflating the conception of scientific representation. International Studies in the Philosophy of Science, 29(1):41–59, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.979671.

Liu:2016:ANF

[Liu16] Chuang Liu. Against the new fictionalism: a hybrid view of scientific models. International Studies in the Philosophy of Science, 30(1):39–54, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2016.1240460.

Livanios:2021:TMM

[Liv21] Vassilis Livanios. The tools of metaphysics and the metaphysics of science.

International Studies in the Philosophy of Science, 34(4):249–252, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2098463.

Lipton:1993:BR

[LLOA93] Peter Lipton, Brendan Larvor, Hans Oberdiek, and Paul Abela. Book reviews. International Studies in the Philosophy of Science, 7(2):191–207, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573463.

Lanz:1995:RGC

[LM95] Peter Lanz and David Mcfarland. On representation, goals and cognition. International Studies in the Philosophy of Science, 9(2):121–133, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573512.

Lockwood:1997:TG

[Loc97] Michael Lockwood. As time goes by.

International Studies in the Philosophy of Science, 11(1):35–51, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573549. Festschrift

for J. R. Lucas.

Lodge:2001:DES

[Lod01] Paul Lodge. The debate over extended substance in Leibniz’s correspondence with De Volder. International Studies in the Philosophy of Science, 15(2):155–165, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120059022.

Logue:1997:RRR

[Log97] James Logue. Resiliency, robustness and rationality of probability judgements. International Studies in the Philosophy of Science, 11(1):21–34, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573548. Festschrift

for J. R. Lucas.

Longino:1997:AST

[Lon97] Helen E. Longino. Alan Sokal’s “transgressing boundaries. International Studies in the Philosophy of Science, 11(2):119–120, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573557.

Lorenzatti:2015:I

[Lor15] Joel J. Lorenzatti. Intuitions. International Studies in the Philosophy of Science, 29(2):221–224, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1119426.

Losee:1992:HSS

[Los92] John Losee. Herbert Simon on scientific discovery. International Studies in the Philosophy of Science, 6(1):41–43, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573410.

Leng:1999:BR

[LPR99] Mary Leng, Kent A. Peacock, and Andrew Reynolds. Book reviews. International Studies in the Philosophy of Science, 13(2):195–204, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573619.

Legrenzi:1993:RGR

[LR93] Paolo Legrenzi and Rino Rumiati. Rationality as “good reasons for”. International Studies in the Philosophy of Science, 7(1):45–49, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573446.

LaFollette:1994:AEL

[LS94] Hugh LaFollette and Niall Shanks.

Animal experimentation: The legacy of Claude Bernard. International Studies in the Philosophy of Science, 8(3):195–210, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573495.

Lucas:1995:PRQ

[Luc95] J. R. Lucas. Prospects for realism in quantum mechanics. International Studies in the Philosophy of Science, 9(3):225–234, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573521.

Lucas:1997:CRT

[Luc97] J. R. Lucas. Comments: Reality and time. International Studies in the Philosophy of Science, 11(1):97–108, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573553. Festschrift

for J. R. Lucas.

Lukes:1993:LI

[Luk93] Steven Lukes. The limits of intelligibility. International Studies in the Philosophy of Science, 7(1):55–59, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573448.

Lyons:2009:NCC

[Lyo09] Timothy D. Lyons. Non-competitor conditions in the scientific realism debate. International Studies in the Philosophy of Science, 23(1):65–84, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590902843393.

Lyons:2013:HIM

[Lyo13] Timothy D. Lyons. A historically informed Modus Ponens against scientific realism: Articulation, critique, and restoration. International Studies in the Philosophy of Science, 27(4):369–392, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868181.

Lyre:2008:DHM

[Lyr08] Holger Lyre. Does the Higgs mechanism exist? International Studies in the Philosophy of Science, 22(2):119–133, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802496664.

Maleeh:2013:PBC

[MA13] Reza Maleeh and Parisa Amani.

Pragmatism, Bohr, and the Copenhagen interpretation of quantum mechanics. International Studies in the Philosophy of Science, 27(4):353–367, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
868182.

MacIntosh:1992:RBE

[Mac92] J. J. MacIntosh. Robert Boyle’s epistemology: The interaction between scientific and religious knowledge. International Studies in the Philosophy of Science, 6(2):91–121, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573419.

MacCallum:2000:CRW

[Mac00] David MacCallum. Conclusive reasons that we perceive sets. International Studies in the Philosophy of Science, 14(1):25–42, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900111882.

MacIntosh:2001:BBC

[Mac01] J. J. MacIntosh. Boyle, Bentley and Clarke on God, necessity, frigorifick atoms and the void. International Studies in the Philosophy of Science, 15(1):33–50, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590020029297.

Machamer:2004:ACM

[Mac04] Peter Machamer. Activities and causation: The metaphysics and epistemology of mechanisms. International Studies in the Philosophy of Science, 18(1):27–39, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590412331289242.

Machamer:2007:DBJ

[Mac07] Peter Machamer. Daniela Bailer-Jones, 1969–2006. International Studies in the Philosophy of Science, 21(2):211–212, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701498159.

Magnus:2008:DIS

[Mag08] P. D. Magnus. Demonstrative induction and the skeleton of inference. International Studies in the Philosophy of Science, 22(3):303–315, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590802567373.

Maki:2013:SID

[Mäk13] Uskali Mäki. Scientific imperialism: Difficulties in definition, identification, and assessment. International Studies in the Philosophy of Science, 27(3):325–339, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.825496.

Makovnik:2023:SEI

[Mak23] Dalibor Makovník. Some explanatory issues with Woodward’s notion of intervention. International Studies in the Philosophy of Science, 36(4):299– 315, 2023. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2023.
2294455.

Malaterre:2007:ORC

[Mal07] Christophe Malaterre. Organicism and reductionism in cancer research: Towards a systemic approach. International Studies in the Philosophy of Science, 21(1):57–73, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701305792.

Malaterre:2012:WLI

[Mal12] Christophe Malaterre. What is life?

The intellectual pertinence of Erwin Schrödinger. International Studies in the Philosophy of Science, 26(2):229– 231, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
703484.

Mamchur:1986:HRA

[Mam87] Elena Mamchur. The heuristic role of aesthetics in science. International Studies in the Philosophy of Science, 1(2):209–222, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573292.

Mamchur:1990:TIT

[Mam90] E. Mamchur. Is there an ivory tower in reality? International Studies in the Philosophy of Science, 4(1):101–111, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573349.

Mamchur:2010:CSG

[Mam10] Elena Mamchur. Contradictions, synthesis, and the growth of knowledge [book review of MR2106565]. International Studies in the Philosophy of Science, 24(4):429–435, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.543352.

Mamchur:2016:ANP

[Mam16] Elena Mamchur. Aristotle’s notion of ‘place’ in the context of presentday physics. International Studies in the Philosophy of Science, 30(4):319– 326, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2017.
                                        1331979.

Mander:1990:BR

[Man90] W. J. Mander. Book review. International Studies in the Philosophy of Science, 4(3):321–323, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573371.

Mander:1991:FHB

[Man91] W. J. Mander. F. H. Bradley and the philosophy of science. International Studies in the Philosophy of Science, 5(1):65–78, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573378.

Marinov:1986:CVS

[Mar87] Marin Marinov. Cognitive values and scientific rationality. International Studies in the Philosophy of Science, 1(2):223–232, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573293.

Marinov:1988:IBE

[Mar89] Marin Marinov. Inference to the best explanation: Van Fraassen and the case of the ‘fifth force’. International Studies in the Philosophy of Science, 3(1):35–50, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573323.

Martinovic:1990:TIT

[Mar90] Ivica Martinović. Theories and intertheory relations in Bosković. International Studies in the Philosophy of Science, 4(3):247–262, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573365.

Marconi:1992:PSP

[Mar92] Diego Marconi. Planetary systems and pomegranites: Doubts about Simon’s essay. International Studies in the Philosophy of Science, 6(1):45–47, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573411.

Marinoff:1996:HBP

[Mar96] Louis Marinoff. How Braess’ paradox solves Newcomb’s problem: Not! International Studies in the Philosophy of Science, 10(3):217–237, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573541. See

[Irv93].

Marquis:1999:MEM

[Mar99] Jean-Pierre Marquis. Mathematical engineering and mathematical change. International Studies in the Philosophy of Science, 13(3):245–259, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573624.

Marcum:2001:CSP

[Mar01] James A. Marcum. Constructing a scientific paper: Howell’s prothrombin laboratory notebook and paper. International Studies in the Philosophy of Science, 15(3):293–310, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080226.

Marcum:2005:MPS

[Mar05] James A. Marcum. Metaphysical presuppositions and scientific practices: Reductionism and organicism in cancer research. International Studies in the Philosophy of Science, 19(1):31–45, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590500051076.

Marcum:2008:ISD

[Mar08] James A. Marcum. Instituting science: Discovery or construction of scientific knowledge? International Studies in the Philosophy of Science, 22(2):185–210, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802496755.

Marcum:2010:HSP

[Mar10a] James A. Marcum. Horizon for scientific practice: Scientific discovery and progress. International Studies in the Philosophy of Science, 24(2):187– 215, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
484545.

Martijn:2010:N

[Mar10b] Marije Martijn. Neoplatonism. International Studies in the Philosophy of Science, 24(1):115–118, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903467150.

Martin:2012:TS

[Mar12] Brian Martin. The tyranny of science. International Studies in the Philosophy of Science, 26(1):118–121, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.653116.

Martin:2015:PB

[Mar15] August W. M. Martin. Philosophy of biology. International Studies in the Philosophy of Science, 29(4):441– 444, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1195149.

Martinez:2016:FIE

[Mar16] María Laura Martínez. Foucauldian imprints in the early works of Ian Hacking. International Studies in the Philosophy of Science, 30(1):69–84, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2016.
                                        1240470.

Massimi:2009:SRP

[Mas09] Michela Massimi. Scientific representation: Paradoxes of perspective. International Studies in the Philosophy of Science, 23(3):323–327, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903197799.

Matuszewski:1986:GTM

[Mat87] Roman Matuszewski. On Galileo’s theory of motion: an attempt at a coherent reconstruction. International Studies in the Philosophy of Science, 1(1):124–141, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573285.

Matheson:1998:WNM

[Mat98] Carl Matheson. Why the nomiracles argument fails. International Studies in the Philosophy of Science, 12(3):263–279, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573601.

Matarese:2021:MAW

[Mat21] Vera Matarese. On the methodological arguments for wave-function realism. International Studies in the Philosophy of Science, 34(2):63–80, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.2024786.

Maxwell:2009:MSA

[Max09] Nicholas Maxwell. The metaphysics of science: An account of modern science in terms of principles, laws and theories. International Studies in the Philosophy of Science, 23(2):228–232, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590903007220.

Mayoral:2022:TAK

[May22] Juan V. Mayoral. Travelling around Kuhn’s worlds. International Studies in the Philosophy of Science, 35(3–4):279–286, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2022.2163134.

Mayoral:2023:NRK

[May23] Juan V. Mayoral. The normal and the revolutionary: Kuhn’s conversations with Rorty. International Studies in the Philosophy of Science, 36(2):105–120, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2023.2248546.

McAllister:1988:ERR

[McA89] James W. McAllister. The explanative recourse to realism. International Studies in the Philosophy of Science, 3(1):2–18, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573321.

McAllister:1997:LNN

[McA97a] James W. McAllister. Laws of nature, natural history, and the description of the world. International Studies in the Philosophy of Science, 11(3):245–258, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573569.

McAllister:1997:PSN

[McA97b] James W. McAllister. Philosophy of science in The Netherlands. International Studies in the Philosophy of Science, 11(2):191–204, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573563.

McAllister:2001:NET

[McA01] James W. McAllister. New editorial team and policy statement. International Studies in the Philosophy of Science, 15(3):229–230, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080163.

McAllister:2002:RWA

[McA02] James W. McAllister. Recent work on aesthetics of science. International Studies in the Philosophy of Science, 16(1):7–11, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120118783.

McAllister:2007:ER

[McA07] James W. McAllister. Editor’s report, 2006. International Studies in the Philosophy of Science, 21(2):119–122, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590701497995.

McAllister:2008:CEP

[McA08a] James W. McAllister. Contours of a European philosophy of science.

International Studies in the Philosophy of Science, 22(1):1–3, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802280852.

McAllister:2008:ER

[McA08b] James W. McAllister. Editor’s report, 2007. International Studies in the Philosophy of Science, 22(2):115–117, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590802496649.

McArthur:2008:TCS

[McA08c] Dan McArthur. Theory change, structural realism, and the relativised A Priori. International Studies in the Philosophy of Science, 22(1):5–20, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590802280860.

McAllister:2009:ER

[McA09] James W. McAllister. Editor’s report, 2008. International Studies in the Philosophy of Science, 23(2):119–121, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590903006818.

McAllister:2010:ER

[McA10] James W. McAllister. Editor’s report, 2009. International Studies in the Philosophy of Science, 24(3):237– 239, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
522406.

McAllister:2011:ER

[McA11] James W. McAllister. Editor’s report, 2010. International Studies in the Philosophy of Science, 25(3):203– 204, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
605243.

McAllister:2012:ER

[McA12] James W. McAllister. Editor’s report, 2011. International Studies in the Philosophy of Science, 26(3):237– 239, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
731729.

McAllister:2013:ER

[McA13] James W. McAllister. Editor’s report, 2012. International Studies in the Philosophy of Science, 27(3):233– 234, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
825495.

McAllister:2014:ER

[McA14] James W. McAllister. Editor’s report, 2013. International Studies in the Philosophy of Science, 28(3):231– 233, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
953340.

McAllister:2015:ER

[McA15] James W. McAllister. Editor’s report, 2014. International Studies in the Philosophy of Science, 29(3):229– 231, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1179033.

McAllister:2016:ER

[McA16] James W. McAllister. Editor’s report, 2015. International Studies in the Philosophy of Science, 30(3):197– 199, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.
1316112.

McAllister:2017:ER

[McA17] James W. McAllister. Editor’s report, 2016. International Studies in the Philosophy of Science, 31(3):225– 227, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2018.
1464099.

McFarland:1991:DMC

[McF91] David McFarland. Defining motivation and cognition in animals. International Studies in the Philosophy of Science, 5(2):153–170, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573387.

McFarland:1992:ACB

[McF92] David McFarland. Animals as cost-based robots. International Studies in the Philosophy of Science, 6(2):133–153, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573422.

Meier:2014:SWM

[Mei14] Thomas Meier. The structure of the world: Metaphysics and representation. International Studies in the Philosophy of Science, 28(4):443–445, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
979674.

Mellor:1990:LCP

[Mel90] D. H. Mellor. Laws, chances and properties. International Studies in the Philosophy of Science, 4(2):159–170, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573355.

Mellor:2009:DC

[Mel09] D. H. Mellor. Dispositions and causes.

International Studies in the Philosophy of Science, 23(3):327–330, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903197807.

Melogno:2023:KNC

[Mel23] Pablo Melogno. Kuhn’s ‘The Natures of Conceptual Change’: the search for a theory of meaning and the birth of taxonomies (1980–1994). International Studies in the Philosophy of Science, 36(2):87–103, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2198861.

Meynell:2008:WFD

[Mey08] Letitia Meynell. Why Feynman diagrams represent. International Studies in the Philosophy of Science, 22(1):39–59, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590802280902.

Maltrana:2022:ETT

[MHB22] Diego Maltrana, Manuel Herrera, and Federico Benitez. Einstein’s theory of theories and mechanicism. International Studies in the Philosophy of Science, 35(2):153–170, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2022.2130661.

Miller:1991:ICT

[Mil91a] Arthur I. Miller. Have incommensurability and causal theory of reference anything to do with actual science? — Incommensurability, no; causal theory, yes. International Studies in the Philosophy of Science, 5(2):97–108, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573383.

Miller:1991:IMC

[Mil91b] Arthur I. Miller. Imagery and meaning, the cognitive science connection. International Studies in the Philosophy of Science, 5(1):35–48, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573376.

Misak:1990:PB

[Mis90] Cheryl Misak. Pragmatism and bivalence. International Studies in the Philosophy of Science, 4(2):171–179, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573356.

Miscevic:1992:MMT

[Mis92] Nenad Miscević. Mental models and thought experiments. International Studies in the Philosophy of Science, 6(3):215–226, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573432.

Miscevic:1996:SRF

[Mis96] Nenad Miscević. Should reason be fragmented? International Studies in the Philosophy of Science, 10(1):23–36, 1996. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573527.

Miscevic:1997:RRN

[Mis97] Nenad Miscević. Resisting the rise of nationalism. International Studies in the Philosophy of Science, 11(3):221–222, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573567.

Miscevic:2001:SCP

[Mis01] Nenad Miscevic. Science, commonsense and philosophy: a defense of continuity (a critique of “network apriorism”). International Studies in the Philosophy of Science, 15(1):19–31, 2001. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590020029288.

Mitchell:1988:CBF

[Mit89] Sandra D. Mitchell. The causal background of functional explanation. International Studies in the Philosophy of Science, 3(2):213–229, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573337.

Mizrahi:2013:AUR

[Miz13] Moti Mizrahi. The argument from underconsideration and relative realism. International Studies in the Philosophy of Science, 27(4):393–407, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.868183.

Mizrahi:2015:HIN

[Miz15] Moti Mizrahi. Historical inductions: New cherries, same old cherry-picking. International Studies in the Philosophy of Science, 29(2):129–148, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119413.

Mizrahi:2016:HSG

[Miz16] Moti Mizrahi. The history of science as a graveyard of theories: a philosophers’ myth? International Studies in the Philosophy of Science, 30(3):263– 278, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.
1316113.

Mizrahi:2017:SPW

[Miz17] Moti Mizrahi. Scientific progress:

Why getting closer to truth is not enough. International Studies in the Philosophy of Science, 31(4):415–419, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.

1565213. See replies [Isl17, CT19].

Mizrahi:2020:HTS

[Miz20] Moti Mizrahi. Hypothesis testing in scientific practice: An empirical study. International Studies in the Philosophy of Science, 33(1):1–21, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1788348.

Mncube:2017:SR

[Mnc17] Zinhle Mncube. Is science racist? International Studies in the Philosophy of Science, 31(3):319–321, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463695.

Mondadori:1993:CBP

[Mon93a] Marco Mondadori. Comments on Boudon’s paper. International Studies in the Philosophy of Science, 7(1):61–63, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573449. See

[Bou93b] and reply [Bou93a].

Montesano:1993:TMR

[Mon93b] Aldo Montesano. On the twofold meaning of rationality in economics. International Studies in the Philosophy of Science, 7(1):65–67, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573450.

Moore:2012:NRM

[Moo12] Dwayne Moore. A non-reductive model of component forces and resultant force. International Studies in the Philosophy of Science, 26(4):359– 380, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
748495.

Morrison:2006:ASA

[Mor06] Margaret Morrison. Applying science and applied science: What’s the difference? International Studies in the Philosophy of Science, 20(1):81–91, 2006. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590600641057.

Mormino:2011:LBS

[Mor11] Gianfranco Mormino. Leibniz: Body, substance, monad. International Studies in the Philosophy of Science, 25(2):191–194, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574862.

Mosconi:1994:IDO

[Mos94] Giuseppe Mosconi. Induction, deduction and other forms of thought. International Studies in the Philosophy of Science, 8(1):61–64, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573482.

Mozersky:2000:TTS

[Moz00] Joshua M. Mozersky. Time, tense and special relativity. International Studies in the Philosophy of Science, 14(3):221–236, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437746.

Martinas:1987:AAM

[MR88] Katalin Martinás and László Ropolyi.

Analogies: Aristotelian and modern physics. International Studies in the Philosophy of Science, 2(1):1–9, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573299.

Murthi:2009:MI

[MS09] S. K. Arun Murthi and Sundar Sarukkai. Multisemiosis and incommensurability. International Studies in the Philosophy of Science, 23(3):297–311, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903197773.

Muller:2015:PMI

[Mül15] Florian Müller. The pessimistic metainduction: Obsolete through scientific progress? International Studies in the Philosophy of Science, 29(4):393– 412, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1195144.

McAllister:1990:BR

[MV90] James W. McAllister and Boris M. Velichkovsky. Book reviews. International Studies in the Philosophy of Science, 4(2):221–225, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573360.

Mozersky:1998:BR

[MWRB98] Joshua Mozersky, Rodney Watkins, Andrew Reynolds, and James Robert Brown. Book reviews. International Studies in the Philosophy of Science, 12(1):91–102, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573584.

Myrvold:2003:SEO

[Myr03] Wayne C. Myrvold. On some early objections to Bohm’s theory. International Studies in the Philosophy of Science, 17(1):7–24, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590305233.

Nadeau:1993:BAG

[Nad93] Robert Nadeau. A bad argument for good reasons. International Studies in the Philosophy of Science, 7(1):69–73, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573451.

Nartonis:2001:LSM

[Nar01] David K. Nartonis. Locke–Stewart– Mill: philosophy of science at Dartmouth College, 1771–1854. International Studies in the Philosophy of Science, 15(2):167–175, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120059031.

Navabi:2007:PSI

[Nav07] Ali Akbar Navabi. Philosophy of science in Iran. International Studies in the Philosophy of Science, 21(1):75–89, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590701305800.

Needham:1999:RAC

[Nee99] Paul Needham. Reduction and abduction in chemistry — a response to Scerri. International Studies in the Philosophy of Science, 13(2):169–184, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573616. See

[Sce98, Sce99, Sce00].

Needham:2000:RCS

[Nee00] Paul Needham. Reduction in chemistry — a second response to Scerri. International Studies in the Philosophy of Science, 14(3):317–323, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/026985900437827.

Needham:2002:DWH

[Nee02] Paul Needham. The discovery that water is H_{2}O. International Studies in the Philosophy of Science, 16(3):205–226, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859022000013300.

Needham:2004:WDA

[Nee04] Paul Needham. When did atoms begin to do any explanatory work in chemistry? International Studies in the Philosophy of Science, 18(2–3):199–219, 2004. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/0269859042000296521.

Needham:2012:NKT

[Nee12] Paul Needham. Natural kind thingamajigs. International Studies in the Philosophy of Science, 26(1):97–101, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2012.

653117. See reply [Bir12].

Needham:2013:QJP

[Nee13] Paul Needham. Questioning the justification of past science. International Studies in the Philosophy of Science, 27(1):85–93, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.783977.

Nell:2020:ETC

[Nel20] Miranda Nell. Exceptional technologies: A continental philosophy of technology. International Studies in the Philosophy of Science, 33(1):62–65, 2020. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2020.
1829833.

Nesic:2022:TNS

[Nes22] Janko Nesić. Towards a neutralstructuralist theory of consciousness and selfhood. International Studies in the Philosophy of Science, 35(3–4):243–259, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2092825.

Newstead:1994:IRD

[New94] Stephen E. Newstead. Inductive reasoning, deductive reasoning and mental models. International Studies in the Philosophy of Science, 8(1):65–67, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573483.

Newman:2001:CAT

[New01] William R. Newman. Corpuscular alchemy and the tradition of Aristotle’s Meteorology, with special reference to Daniel Sennert. International Studies in the Philosophy of Science, 15(2):145–153, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120059013.

Newman:2012:IMS

[New12] Mark Newman. An inferential model of scientific understanding. International Studies in the Philosophy of Science, 26(1):1–26, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.653118.

Newman:2013:RIM

[New13] Mark Newman. Refining the inferential model of scientific understanding. International Studies in the Philosophy of Science, 27(2):173–197, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.813253.

Nielsen:2008:POB

[Nie08] Katherine Nielsen. The philosophy of Osman bin Bakar. International Studies in the Philosophy of Science, 22(1):81–95, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590802280936.

Niiniluoto:1990:STI

[Nii90] Ilkka Niiniluoto. Should technological imperatives be obeyed? International Studies in the Philosophy of Science, 4(2):181–189, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573357.

Nnaji:2016:CSD

[NL16] John Nnaji and José Luis Luján. The content of science debate in the historiography of the scientific revolution. International Studies in the Philosophy of Science, 30(2):99–109, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265864.

Norton:2000:WCW

[Nor00] John D. Norton. What can we learn about physical laws from the fact that we have memories only of the past? International Studies in the Philosophy of Science, 14(1):11–23, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900111873.

Norris:2001:PRR

[Nor01] Christopher Norris. Putnam on realism, reference and truth: The problem with quantum mechanics. International Studies in the Philosophy of Science, 15(1):65–91, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590020029314.

Norton:2007:DDB

[Nor07] John D. Norton. Disbelief as the dual of belief. International Studies in the Philosophy of Science, 21(3):231–252, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590701589536.

Novak:2010:TTM

[Nov10] Zsolt Novák. Truth and truthmaking. International Studies in the Philosophy of Science, 24(3):323–326, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2010.
                                        522414.

Newton-Smith:1992:SDP

[NS92] W. H. Newton-Smith. “Scientific discovery as problem solving” by H. A. Simon. International Studies in the Philosophy of Science, 6(1):49–52, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573412. See

[Sim92a].

Newton-Smith:1995:E

[NS95] W. H. Newton-Smith. Editorial. International Studies in the Philosophy of Science, 9(3):197, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573517.

Newton-Smith:1986:Ia

[NSW87a] W. H. Newton-Smith and K. V. Wilkes. Introduction. International Studies in the Philosophy of Science, 1(1):2, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573278.

Newton-Smith:1986:Ib

[NSW87b] W. H. Newton-Smith and K. V.

Wilkes. Introduction. International Studies in the Philosophy of Science, 1(2):141–142, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573288.

Newton-Smith:1987:Ia

[NSW88a] W. H. Newton-Smith and K. V.

Wilkes. Introduction. International Studies in the Philosophy of Science, 2(1):5, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573298.

Newton-Smith:1987:Ib

[NSW88b] W. H. Newton-Smith and K. V.

Wilkes. Introduction. International Studies in the Philosophy of Science, 2(2):117, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573309.

Newton-Smith:1988:Ia

[NSW89a] W. H. Newton-Smith and K. V.

Wilkes. Introduction. International Studies in the Philosophy of Science, 3(1):1, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573320.

Newton-Smith:1988:Ib

[NSW89b] W. H. Newton-Smith and K. V. Wilkes. Introduction. International Studies in the Philosophy of Science, 3(2):1, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598908573330.

Nugayev:1987:GSM

[Nug88] R. M. Nugayev. The genesis and structure of models in the modern theory of gravity. International Studies in the Philosophy of Science, 2(1):84–104, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573304.

Nugayev:1991:FLP

[Nug91] Renat Nugayev. The fundamental laws of physics can tell the truth. International Studies in the Philosophy of Science, 5(1):79–87, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573379.

Nugayev:1996:WDN

[Nug96] Rinat M. Nugayev. Why did the new physics force out the old? International Studies in the Philosophy of Science, 10(2):127–140, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573534.

Oberdiek:1990:TAN

[Obe90] Hans Oberdiek. Technology: Autonomous or neutral. International Studies in the Philosophy of Science, 4(1):67–77, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573346.

Oftedal:2010:DPN

[Oft10] Gry Oftedal. Darwinian populations and natural selection. International Studies in the Philosophy of Science, 24(3):333–336, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522417.

OHear:1994:KEC

[O’H94] Anthony O’Hear. Knowledge in an evolutionary context. International Studies in the Philosophy of Science, 8(2):125–138, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573489.

Okasha:2000:ESB

[Oka00a] Samir Okasha. The explanation of scientific belief: Reply to W. E. Jones. International Studies in the Philosophy of Science, 14(3):305–306, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437809.

Okasha:2000:UTD

[Oka00b] Samir Okasha. The underdetermination of theory by data and the “strong programme” in the sociology of knowledge. International Studies in the Philosophy of Science, 14(3):283–297, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437782. See re-

sponse [Jon00].

Okasha:1996:BR

[OKHT96] Samir Okasha, Joel Katzav, Matthias Hild, and John Taylor. Book reviews. International Studies in the Philosophy of Science, 10(3):257–270, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573544.

Okruhlik:1998:BRO

[Okr98] Kathleen Okruhlik. Book review: Otto Neurath: Philosophy Between Science and Politics by Nancy Cartwright, Jordi Cat, Lola Fleck and Thomas E. Uebel. International Studies in the Philosophy of Science, 12(2):175–191, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573591.

Over:1994:IP

[OM94] D. E. Over and K. I. Manktelow. Induction and probability. International Studies in the Philosophy of Science, 8(1):57–60, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573481.

Okyere-Manu:2016:OLM

[OM16] Beatrice Okyere-Manu. Overpopulation and the lifeboat metaphor: a critique from an African worldview. International Studies in the Philosophy of Science, 30(3):279–289, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1316114.

Okyere-Manu:2017:EET

[OM17] Beatrice Okyere-Manu. Environmental ethics: From theory to practice. International Studies in the Philosophy of Science, 31(1):110–112, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1370931.

Oriti:2024:TWR

[Ori24] Daniele Oriti. Tightrope-walking rationality in action: Feyerabendian insights for the foundations of quantum mechanics. International Studies in the Philosophy of Science, 37(4):161– 193, 2024. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2024.
2351917.

Orr:2024:BRP

[Orr24] Avery Orr. Book review: A Philosopher Looks at Science, by Nancy Cartwright, Cambridge, Cambridge University Press, 2022, 222 pp., $12.99, £9.99 (paperback), ISBN 978-1-009-20188-9. International Studies in the Philosophy of Science, 37(3):147–150, 2024. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2024.
                                        2379743.

OrtizdeLandazuri:2024:SPC

[Ort24] Manuel C. Ortiz de Landázuri. Schlick and Popper on causality and quantum physics: Origins and perspectives of the debate. International Studies in the Philosophy of Science, 37(3):75– 94, 2024. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2024.
                                        2379742.

PerezOtero:2018:EPA

[P´ 18] Manuel Pérez Otero. An epistemic puzzle about knowledge and rational credence. International Studies in the Philosophy of Science, 32(3–4):195– 206, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2019.
                                        1704615.

Paavola:2006:HHA

[Paa06] Sami Paavola. Hansonian and Harmanian abduction as models of discovery. International Studies in the Philosophy of Science, 20(1):93–108, 2006. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590600641065.

Pages:2002:DTA

[Pag02] Joan Pagès. The Dretske–Tooley– Armstrong theory of natural laws and the inference problem. International Studies in the Philosophy of Science, 16(3):227–243, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013319.

Palacios:2016:CTA

[Pal16] Patricia Palacios. Chance and temporal asymmetry. International Studies in the Philosophy of Science, 30(2):185–187, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265865.

Panagiotatou:2017:SRQ

[Pan17] Maria Panagiotatou. Scientific realism and quantum mechanics: Revisiting a controversial relation. International Studies in the Philosophy of Science, 31(3):243–259, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463696.

Paoletti:2017:EPB

[Pao17] Michele Paolini Paoletti. Empirical physicalism and the boundaries of physics. International Studies in the Philosophy of Science, 31(4):343–362, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.
1565207.

Papineau:1987:MF

[Pap88a] David Papineau. Mathematical fictionalism. International Studies in the Philosophy of Science, 2 (2):151–174, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573312.

Pappas:1987:SMB

[Pap88b] George S. Pappas. Science and metaphysics in Berkeley. International Studies in the Philosophy of Science, 2(1):105–114, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573305.

Papineau:1990:CMP

[Pap90] David Papineau. Causes and mixed probabilities. International Studies in the Philosophy of Science, 4(1):79–88, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573347.

Parusnikova:1990:PWH

[Par90] Dr Zuzana Parusnikova. Popper’s world 3 & human creativity. International Studies in the Philosophy of Science, 4(3):263–269, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573366.

Parker:2008:FHE

[Par08] Wendy S. Parker. Franklin, Holmes, and the epistemology of computer simulation. International Studies in the Philosophy of Science, 22(2):165–183, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590802496722.

Parigi:2015:EAD

[Par15] Silvia Parigi. Effluvia, action at a distance, and the challenge of the third causal model. International Studies in the Philosophy of Science, 29(4):351– 368, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1195141.

Parageau:2016:C

[Par16a] Sandrine Parageau. Cavendish. International Studies in the Philosophy of Science, 30(3):307–309, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1316115.

Parkkinen:2016:EDC

[Par16b] Veli-Pekka Parkkinen. Experiment, downward causation, and interventionist levels of explanation. International Studies in the Philosophy of Science, 30(3):245–261, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1316116.

Parigi:2020:MQH

[Par20] Silvia Parigi. Molyneux’s question and the history of philosophy. International Studies in the Philosophy of Science, 33(3):189–193, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1940464.

Paramonov:2024:MEM

[Par24] Andrei Paramonov. In memoriam Elena Mamchur 8 July, 1935–14 December, 2023. International Studies in the Philosophy of Science, 37(1–2): 69–73, 2024. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2024.
2341367.

Paul:2012:BRL

[Pau12] Herman Paul. Book review: The Life and Thought of Herbert Butterfield: History, Science and God [by Michael Bentley, Cambridge, Cambridge University Press, 2011. xv + 381 pp. ISBN 978-1-107-00397-2]. International Studies in the Philosophy of Science, 26(2):232–235, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.703485.

Paul:2016:TPI

[Pau16] Emily Paul. Time: a philosophical introduction. International Studies in the Philosophy of Science, 30(1):85– 88, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2016.
1240472.

Paya:1995:PAT

[Pay95] A. Paya. Philosophers against “truth”: The cases of Harré and Laudan. International Studies in the Philosophy of Science, 9(3):255–284, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573523.

Paya:2015:CAP

[Pay15] Ali Paya. A critical assessment of the programmes of producing ‘Islamic science’ and ‘Islamisation of science/ knowledge’. International Studies in the Philosophy of Science, 29(3):311– 335, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1179043.

Pietarinen:2014:NLP

[PB14] Ahti-Veikko Pietarinen and Francesco Bellucci. New light on Peirce’s conceptions of retroduction, deduction, and scientific reasoning. International Studies in the Philosophy of Science, 28(4):353–373, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979667.

Peacock:1998:EPS

[Pea98] Kent A. Peacock. On the edge of a paradigm shift: Quantum nonlocality and the breakdown of peaceful coexistence. International Studies in the Philosophy of Science, 12(2):129–150, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573588.

Peacock:2002:BBQ

[Pea02] Kent A. Peacock. Bub and the barriers to quantum ontology. International Studies in the Philosophy of Science, 16(3):285–289, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013346.

Pearce:2010:COT

[Pea10] Jacob V. Pearce. Constituting objectivity: Transcendental perspectives on modern physics. International Studies in the Philosophy of Science, 24(4):440–443, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.543354.

Pechenkin:2002:MIQ

[Pec02] A. A. Pechenkin. Mandelstam’s interpretation of quantum mechanics in comparative perspective. International Studies in the Philosophy of Science, 16(3):265–284, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013337.

Peden:2018:SCA

[Ped18] William Peden. The selective confirmation answer to the paradox of the ravens. International Studies in the Philosophy of Science, 32(3–4):177– 193, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2020.
1768014.

Penrose:1997:UU

[Pen97] Roger Penrose. On understanding understanding. International Studies in the Philosophy of Science, 11(1):7–20, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573547. Festschrift

for J. R. Lucas.

Perini:2012:IIB

[Per12] Laura Perini. Image interpretation:

Bridging the gap from mechanically produced image to representation. International Studies in the Philosophy of Science, 26(2):153–170, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.703478.

Pernu:2013:DIN

[Per13a] Tuomas K. Pernu. Does the interventionist notion of causation deliver us from the fear of epiphenomenalism? International Studies in the Philosophy of Science, 27(2):157–172, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.813254.

Pero:2013:MMB

[Per13b] Francesca Pero. Models as makebelieve: Imagination, fiction and scientific representation. International Studies in the Philosophy of Science, 27(4):447–450, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.868184.

Perovic:2014:CBF

[Per14] Slobodan Perović. Causation and its basis in fundamental physics. International Studies in the Philosophy of Science, 28(3):347–349, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.953347.

Perovic:2021:OES

[Per21] Slobodan Perović. Observation, experiment, and scientific practice. International Studies in the Philosophy of Science, 34(1):1–20, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2021.1978038.

Petroni:1992:WHS

[Pet92] Angelo M. Petroni. Why have a heuristic of scientific discovery? International Studies in the Philosophy of Science, 6(1):53–55, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573413.

Petroni:1993:CSD

[Pet93a] Angelo M. Petroni. Conventionalism, scientific discovery and the sociology of knowledge. International Studies in the Philosophy of Science, 7(3):225–240, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573467.

Petroni:1993:I

[Pet93b] Angelo M. Petroni. Introduction. International Studies in the Philosophy of Science, 7(1):3, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573438.

Peters:2023:LDS

[Pet23] Uwe Peters. Linguistic discrimination in science: Can English disfluency help debias scientific research? International Studies in the Philosophy of Science, 36(1):61–79, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2251676.

Paya:2006:PRS

[PG06] Ali Paya and Mohammad Amin Ghaneirad. The philosopher and the revolutionary state: How Karl Popper’s ideas shaped the views of Iranian intellectuals. International Studies in the Philosophy of Science, 20(2):185–213, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814381.

Phillips:2014:HCO

[Phi14] Denise Phillips. Hans Christian Ørsted: Reading Nature’s mind. International Studies in the Philosophy of Science, 28(3):341–343, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.953345.

Pietsch:2012:HUC

[Pie12] Wolfgang Pietsch. Hidden underdetermination: a case study in classical electrodynamics. International Studies in the Philosophy of Science, 26(2):125– 151, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2012.
                                        703477.

Pihlstrom:1996:GON

[Pih96] Sami Pihlström. Getting ontologically natural. International Studies in the Philosophy of Science, 10(3):247–256, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573543.

Pils:2023:SRB

[Pil23] Raimund Pils. Scientific realism and blocking strategies*. International Studies in the Philosophy of Science, 36(1):1–17, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2022.2133418.

Pineda:2006:MCP

[Pin06] David Pineda. A mereological characterization of physicalism. International Studies in the Philosophy of Science, 20(3):243–266, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600960929.

Pineda:2010:NCC

[Pin10] David Pineda. Non-committal causal explanations. International Studies in the Philosophy of Science, 24(2):147– 170, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2010.
                                        484540.

Pincock:2011:FSP

[Pin11a] Christopher Pincock. Fictions in science: Philosophical essays on modeling and idealization. International Studies in the Philosophy of Science, 25(2):196–199, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574867.

Pinto:2011:EEN

[Pin11b] Ana Teixeira Pinto. Emergence and embodiment: New essays on second-order systems theory. International Studies in the Philosophy of Science, 25(1):86–89, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.552421.

Pincock:2024:DRS

[Pin24] Christopher Pincock. Defending a realist stance. International Studies in the Philosophy of Science, 37(1–2):1– 15, 2024. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2024.
2303576.

Piotrowska:2014:GPI

[Pio14] Monika Piotrowska. Genetics and philosophy: An introduction. International Studies in the Philosophy of Science, 28(2):223–226, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.932533.

Pizzorno:1993:AYC

[Piz93] Alessandro Pizzorno. All you can do with reasons. International Studies in the Philosophy of Science, 7(1):75–80, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573452.

Plaud:2010:BRW

[Pla10] Sabine Plaud. Book review: Wittgenstein and Analytic Philosophy: Essays for P. M. S. Hacker, [by Hans-Johan Glock and John Hyman (Eds), Oxford, Oxford University Press, 2009, xxii + 315 pp., ISBN 978-0-19-921323-8, £40.00 (hardback)]. International Studies in the Philosophy of Science, 24(2):223–226, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.484558.

Percival:1994:BR

[PLO94] Philip Percival, J. R. Lucas, and Anthony O’Hear. Book reviews. International Studies in the Philosophy of Science, 8(3):265–276, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573500.

Politi:2017:SII

[Pol17] Vincenzo Politi. Specialisation, interdisciplinarity, and incommensurability. International Studies in the Philosophy of Science, 31(3):301–317, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463697.

Politi:2020:TNL

[Pol20] Vincenzo Politi. Taxonomies, networks, and lexicons: A study of Kuhn’s post-‘linguistic turn’ philosophy. International Studies in the Philosophy of Science, 33(2):87–103, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1865784.

Popa:2016:CPA

[Pop16] Elena Popa. Causal projectivism, agency, and objectivity. International Studies in the Philosophy of Science, 30(2):147–163, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265866.

Price:2000:GPC

[Pri00] Carolyn Price. General-purpose content. International Studies in the Philosophy of Science, 14(2):123–133, 2000. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698590050045719.

Preston:1992:BR

[PS92] John M. Preston and Alan Soble. Book reviews. International Studies in the Philosophy of Science, 6(2):155–162, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698599208573423.

Pinnick:2001:CCP

[PS01] Cassandra L. Pinnick and Warren Schmaus. Changing conceptions of the philosophy of science. International Studies in the Philosophy of Science, 15(2):127–131, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120058997.

Politi:2023:BSN

[PS23] Vincenzo Politi and Yafeng Shan. Beyond structure: New frontiers of the philosophy of Thomas Kuhn. International Studies in the Philosophy of Science, 36(2):81–86, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2264158.

Psillos:1997:KR

[Psi97] Stathis Psillos. Kitcher on reference. International Studies in the Philosophy of Science, 11(3):259–272, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573570.

Psillos:2000:AEV

[Psi00] Stathis Psillos. Agnostic empiricism versus scientific realism: Belief in truth matters. International Studies in the Philosophy of Science, 14(1):57–75, 2000. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900111909.

Palacio-Vera:2020:PRP

[PV20] Alfonso Palacio-Vera. Popper’s ‘rationality principle’ and ‘epistemic’ rationality: an attempt at reconciliation. International Studies in the Philosophy of Science, 33(4):195–209, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1960108.

Petersen:2014:ENS

[PZ14] Graciana Petersen and Frank Zenker.

From Euler to Navier–Stokes: a spatial analysis of conceptual changes in nineteenth-century fluid dynamics. International Studies in the Philosophy of Science, 28(3):235–253, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.953341.

Psillos:2020:ACI

[PZ20] Stathis Psillos and Lisa Zorzato.

Against cognitive instrumentalism. International Studies in the Philosophy of Science, 33(4):247–257, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1961420.

Qinggang:2016:HCS

[Qin16] Hao Qinggang. A history of Chinese science and technology, volume 1. International Studies in the Philosophy of Science, 30(2):188–190, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265861.

Qiu:1988:MEE

[Qiu89] Ren-Zong Qiu. Models of explanation and explanation in medicine. International Studies in the Philosophy of Science, 3(2):199–212, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573336.

Reutlinger:2016:AVC

[RA16] Alexander Reutlinger and Holly Andersen. Abstract versus causal explanations? International Studies in the Philosophy of Science, 30(2):129–146, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2016.
1265867.

Radder:2002:RAO

[Rad02] Hans Radder. Review article: The origin and nature of modern science [Science and modernity: Toward an Integral Theory of Science, Srdan Lelas, Dordrecht, Kluwer Academic Publishers, 2000. xiii + 290 pp., EUR 111, US$129, £80 (hardback), ISBN 0-7923-6305-5; EUR 45, US$42, £28 (paperback), ISBN 1-4020-0247-5]. International Studies in the Philosophy of Science, 16(3):291–295, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013355.

Radder:2010:RSV

[Rad10] Hans Radder. Rethinking science and values. International Studies in the Philosophy of Science, 24(1):107–114, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590903467143.

Radder:2015:HIE

[Rad15] Hans Radder. How inclusive is European philosophy of science? International Studies in the Philosophy of Science, 29(2):149–165, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119416.

Raley:2007:FEC

[Ral07] Yvonne Raley. The facticity of explanation and its consequences. International Studies in the Philosophy of Science, 21(2):123–135, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701498035.

Ramsey:1993:WRL

[Ram93] Jeffry L. Ramsey. When reduction leads to construction: Design considerations in scientific methodology. International Studies in the Philosophy of Science, 7(3):241–253, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573468.

Ratzsch:1988:ESS

[Rat89] Del Ratzsch. Explanation, subjunctives and statistical theories. International Studies in the Philosophy of Science, 3(1):80–96, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573326.

Ratcliffe:2005:EPE

[Rat05] Matthew Ratcliffe. An epistemological problem for evolutionary psychology. International Studies in the Philosophy of Science, 19(1):47–63, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500051126.

Rayan:2009:NDW

[Ray09] Sobhi Rayan. Nominal definition in the writings of Ibn Taymiyya. International Studies in the Philosophy of Science, 23(2):123–141, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903006859.

Root-Bernstein:2002:AC

[RB02] Robert S. Root-Bernstein. Aesthetic cognition. International Studies in the Philosophy of Science, 16(1):61–77, 2002. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120118837.

Read:2020:GCM

[Rea20] James Read. Geometrical constructivism and modal relationalism: Further aspects of the dynamical/geometrical debate. International Studies in the Philosophy of Science, 33(1):23–41, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2020.1813530.

Recio:2017:AVU

[Rec17] Gonzalo L. Recio. Adjusting Venus:

The use of maximum elongations in the Almagest and Ptolemy’s Theory of Knowledge. International Studies in the Philosophy of Science, 31(2):113– 131, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2018.
1424762.

Redei:1992:WCN

[Réd92] Miklós Rédei. When can noncommutative statistical inference be Bayesian? International Studies in the Philosophy of Science, 6(2):129–132, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573421.

Reisch:2001:ATD

[Rei01] George A. Reisch. Against a third dogma of logical empiricism: Otto Neurath and “unpredictability in principle”. International Studies in the Philosophy of Science, 15(2):199–209, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120059068.

Renzi:2013:KES

[Ren13] Barbara G. Renzi. Kuhn’s evolutionary social epistemology. International Studies in the Philosophy of Science, 27(1):101–104, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.783978.

Resnik:1994:MCS

[Res94] D. Resnik. Methodological conservatism and social epistemology. International Studies in the Philosophy of Science, 8(3):247–264, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573499.

Reutlinger:2011:TNU

[Reu11] Alexander Reutlinger. A theory of non-universal laws. International Studies in the Philosophy of Science, 25(2):97–117, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574853.

Reutlinger:2013:CIN

[Reu13] Alexander Reutlinger. Can interventionists be neo-Russellians? Interventionism, the open systems argument, and the arrow of entropy. International Studies in the Philosophy of Science, 27(3):273–293, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.825497.

Reynolds:1999:WH

[Rey99] Andrew Reynolds. What is historicism? International Studies in the Philosophy of Science, 13(3):275–287, 1999. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573626.

Ruttkamp:2005:RRP

[RH05] Emma Ruttkamp and Johannes Heidema. Reviewing reduction in a preferential model-theoretic context. International Studies in the Philosophy of Science, 19(2):123–146, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500249431.

Rini:2012:EPI

[Rin12] Regina A. Rini. Experimental philosophy: An introduction. International Studies in the Philosophy of Science, 26(4):457–460, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.748500.

RiojaNieto:2017:NBP

[Rio17] Ana Rioja Nieto. Niels Bohr and the philosophy of physics: Twentyfirst-century perspectives. International Studies in the Philosophy of Science, 31(4):429–432, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565209.

Robinson:1992:HHI

[Rob92] Don Robinson. On Healey’s holistic interpretation of quantum mechanics. International Studies in the Philosophy of Science, 6(3):227–240, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573433.

Robinson:1995:RO

[Rob95] Daniel N. Robinson. Radical ontologies. International Studies in the Philosophy of Science, 9(3):215–223, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573520.

Rockmann:1998:GLH

[Roc98] Jutta Rockmann. Gravitational lensing and Hacking’s extragalactic irreality. International Studies in the Philosophy of Science, 12(2):151–164, 1998. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573589.

Rockmore:2010:RTR

[Roc10] Tom Rockmore. Reason, truth, and reality. International Studies in the Philosophy of Science, 24(4):449–451, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2010.
                                        543359.

Roe:2017:JDS

[Roe17] Sarah M. Roe. The journey from discovery to scientific change: Scientific communities, shared models, and specialised vocabulary. International Studies in the Philosophy of Science, 31(1):47–67, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1370932.

Rohrlich:2004:RDC

[Roh04] Fritz Rohrlich. Realism despite cognitive antireductionism. International Studies in the Philosophy of Science, 18(1):73–88, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590412331289260.

Rolin:2015:EIE

[Rol15] Kristina Rolin. Economics imperialism and epistemic cosmopolitanism. International Studies in the Philosophy of Science, 29(4):413–429, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1195145.

Romeyn:2005:ET

[Rom05] Jan-Willem Romeyn. Enantiomorphy and time. International Studies in the Philosophy of Science, 19(2):167–190, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500249498.

Roseberg:1988:HEM

[Rös89] Ulrich Röseberg. Historical explanations in modern physics? The lesson of quantum mechanics. International Studies in the Philosophy of Science, 3(1):68–79, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573325.

Rouse:2011:AWE

[Rou11] Joseph Rouse. Articulating the world: Experimental systems and conceptual understanding. International Studies in the Philosophy of Science, 25(3):243–254, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.605246.

Rowbottom:2008:BTC

[Row08] Darrell P. Rowbottom. The big test of corroboration. International Studies in the Philosophy of Science, 22(3):293–302, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802567357.

Rowbottom:2010:WSP

[Row10] Darrell P. Rowbottom. What scientific progress is not: Against Bird’s epistemic view. International Studies in the Philosophy of Science, 24(3):241–255, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522407.

Radner:1995:CNS

[RR95] Daisie Radner and Michael Radner. Cognition, natural selection and the intentional stance. International Studies in the Philosophy of Science, 9(2):109–119, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573511.

Ryghaug:2010:GWC

[RS10] Marianne Ryghaug and Tomas Moe Skjølsvold. The global warming of climate science: Climategate and the construction of scientific facts. International Studies in the Philosophy of Science, 24(3):287–307, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522411.

Rueger:2002:AAE

[Rue02] Alexander Rueger. Aesthetic appreciation of experiments: The case of 18th-century mimetic experiments. International Studies in the Philosophy of Science, 16(1):49–59, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120118828.

Ruetsche:2003:MSM

[Rue03] Laura Ruetsche. Modal semantics, modal dynamics and the problem of state preparation. International Studies in the Philosophy of Science, 17(1): 25–41, 2003. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698590305235.

Ruhmkorff:2013:GLP

[Ruh13] Samuel Ruhmkorff. Global and local pessimistic meta-inductions. International Studies in the Philosophy of Science, 27(4):409–428, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.868185.

Runehov:2011:RSC

[Run11] Anne L. C. Runehov. Religion and science in context: a guide to the debates. International Studies in the Philosophy of Science, 25(3):303–305, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2011.
                                        605254.

Ruphy:2005:WMA

[Rup05] Stéphanie Ruphy. Why metaphysical abstinence should prevail in the debate on reductionism. International Studies in the Philosophy of Science, 19(2):105–121, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500249415.

Russo:2010:CAS

[Rus10] Federica Russo. Are causal analysis and system analysis compatible approaches? International Studies in the Philosophy of Science, 24(1):67–90, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590903467127.

Rush:2013:AMS

[Rus13] Penelope Rush. The applicability of mathematics in science: Indispensability and ontology. International Studies in the Philosophy of Science, 27(2):219–222, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.813255.

Russo:2014:WIH

[Rus14] Federica Russo. What invariance is and how to test for it. International Studies in the Philosophy of Science, 28(2):157–183, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.932528.

Ruyant:2021:SIP

[Ruy21] Quentin Ruyant. Symmetries, indexicality and the perspectivist stance. International Studies in the Philosophy of Science, 34(1):21–39, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.2001723.

Ruzzene:2011:CSR

[Ruz11] Attilia Ruzzene. The case study as research method: a practical handbook. International Studies in the Philosophy of Science, 25(3):293–296, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.605251.

Russo:2007:ICH

[RW07] Federica Russo and Jon Williamson.

Interpreting causality in the health sciences. International Studies in the Philosophy of Science, 21(2):157–170, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590701498084.

Saatsi:2007:LHN

[Saa07] Juha T. Saatsi. Living in harmony: Nominalism and the explanationist argument for realism. International Studies in the Philosophy of Science, 21(1):19–33, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701305743.

Salvaggio:1993:STR

[Sal93] Salvino A. Salvaggio. A synthetic theory of rationality: Propositions and aporia. International Studies in the Philosophy of Science, 7(1):81–84, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573453.

Sankey:1996:NNC

[San96] Howard Sankey. Normative naturalism and the challenge of relativism: Laudan versus Worrall on the justification of methodological principles. International Studies in the Philosophy of Science, 10(1):37–51, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573528.

Sankey:1998:TI

[San98] Howard Sankey. Taxonomic incommensurability. International Studies in the Philosophy of Science, 12(1):7–16, 1998. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573578.

Sarkar:2000:EEU

[Sar00] Husain Sarkar. Empirical equivalence and underdetermination. International Studies in the Philosophy of Science, 14(2):187–197, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590050045755.

Saridakis:2016:IRM

[Sar16] Emmanuel Saridakis. Information, reality, and modern physics. International Studies in the Philosophy of Science, 30(4):327–341, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331980.

Schickore:2013:UMM

[SC13] Jutta Schickore and Klodian Coko. Using multiple means of determination. International Studies in the Philosophy of Science, 27(3):295–313, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.825498.

Scerri:1998:PNA

[Sce98] Eric R. Scerri. Popper’s naturalized approach to the reduction of chemistry. International Studies in the Philosophy of Science, 12(1):33–44, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573581. See

response [Nee99] and replies [Sce99, Sce00].

Scerri:1999:RN

[Sce99] Eric R. Scerri. Response to Needham.

International Studies in the Philosophy of Science, 13(2):185–192, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573617. See

[Sce98, Nee99, Sce00].

Scerri:2000:SRP

[Sce00] Eric R. Scerri. Second response to Paul Needham. International Studies in the Philosophy of Science, 14(3):307–315, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/026985900437818.  See

[Sce98, Nee99, Sce99].

Schneider:1998:GPH

[Sch98] Ulrich Johannes Schneider. GDR philosophy is history. International Studies in the Philosophy of Science, 12(3):213–215, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573596.

Schickore:2009:SJP

[Sch09] Jutta Schickore. Studying justificatory practice: An attempt to integrate the history and philosophy of science. International Studies in the Philosophy of Science, 23(1):85–107, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590902843401.

Schiaffonati:2011:PEE

[Sch11] Viola Schiaffonati. Philosophy and engineering: An emerging agenda. International Studies in the Philosophy of Science, 25(1):89–92, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.552422.

Schlosser:2012:FWM

[Sch12] Markus Schlosser. Free will and modern science. International Studies in the Philosophy of Science, 26(4):463– 466, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
748502.

Schyfter:2017:SSE

[Sch17] Pablo Schyfter. Science as social existence: Heidegger and the sociology of scientific knowledge. International Studies in the Philosophy of Science, 31(4):432–435, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565210.

Schliesser:2022:BBW

[Sch22] Eric Schliesser. Bad beliefs: Why they happen to good people. International Studies in the Philosophy of Science, 35(2):189–203, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2119492.

Sciortino:2016:SRH

[Sci16] Luca Sciortino. Styles of reasoning, human forms of life, and Relativism. International Studies in the Philosophy of Science, 30(2):165–184, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265868.

Seager:1986:SET

[Sea87] William E. Seager. Scientific explanation and the trial of Galileo. International Studies in the Philosophy of Science, 1(2):176–195, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573290.

Seager:1990:IP

[Sea90] William Seager. Instrumentalism in psychology. International Studies in the Philosophy of Science, 4(2):191–203, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573358.

Seager:1999:RN

[Sea99] William Seager. The reality of now. International Studies in the Philosophy of Science, 13(1):69–82, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573608.

Sedivy:1996:CNP

[Sed96] Sonia Sedivy. Conventional naturalism: a perceptualist account of pictorial representation. International Studies in the Philosophy of Science, 10(2):103–125, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573533.

Seifert:2016:EPC

[Sei16] Vanessa A. Seifert. Essays in the philosophy of chemistry. International Studies in the Philosophy of Science, 30(4):415–417, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331981.

Sepetyi:2025:PLW

[Sep25] Dmytro Sepetyi. Popper and Lakatos on what is distinctive about empirical science. International Studies in the Philosophy of Science, 38(1):39– 58, 2025. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2025.
2449809.

Sesardic:2011:MSB

[Ses11] Neven Sesardić. The mirage of a space between nature and nurture. International Studies in the Philosophy of Science, 25(4):417–420, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.623371.

Schank:1992:DSQ

[SH92] Roger C. Schank and Lucian P.

Hughes. Discovering scientific questions. International Studies in the Philosophy of Science, 6(1):57–59, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573414.

Shah:2007:IJA

[Sha07] Mehul Shah. Is it justifiable to abandon all search for a logic of discovery? International Studies in the Philosophy of Science, 21(3):253–269, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701589544.

Shahvisi:2016:THP

[Sha16] Arianne Shahvisi. Tainted: How philosophy of science can expose bad science. International Studies in the Philosophy of Science, 30(2):193–196, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2016.
                                        1265869.

Shaw:2020:DGS

[Sha20] Jamie Shaw. Duhem on good sense and theory pursuit: From virtue to social epistemology. International Studies in the Philosophy of Science, 33(2): 67–85, 2020. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2021.
                                        1888191.

Sheehan:2007:MSS

[She07] Helena Sheehan. Marxism and science studies: a sweep through the decades. International Studies in the Philosophy of Science, 21(2):197–210, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701498126.

Shech:2021:DSE

[She21] Elay Shech. Darwinian-selectionist explanation, radical theory change, and the observable-unobservable dichotomy. International Studies in the Philosophy of Science, 34(4):221–241, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.

2092824. See reply [Wra21].

Shogenji:2018:HPS

[Sho18] Tomoji Shogenji. Hume’s problem solved: The optimality of metainduction. International Studies in the Philosophy of Science, 32(3–4):229– 231, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2020.
1767892.

Sialaros:2019:GTB

[Sia19] Michalis Sialaros. Geometry: The third book of foundations. International Studies in the Philosophy of Science, 32(1):75–77, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1615667.

Sikic:1992:DAS

[Sik92] Zvonimir Sikić. The diagonal argument — a study of cases. International Studies in the Philosophy of Science, 6(3):191–203, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573430.

Sikic:1996:WN

[Šik96] Zvonimir Šikić. What are numbers?

International Studies in the Philosophy of Science, 10(2):159–171, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573536.

Simon:1992:SDPa

[Sim92a] Herbert A. Simon. Scientific discovery as problem solving. International Studies in the Philosophy of Science, 6(1):3–14, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573403. See

comments [Gil92, Hes92, NS92, Tra92] and response [Sim92b].

Simon:1992:SDPb

[Sim92b] Herbert A. Simon. Scientific discovery as problem solving: Reply to critics. International Studies in the Philosophy of Science, 6(1):69–88, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573417.

Simoes:2016:PCG

[Sim16] Ana Simões. Philosophical chemistry: Genealogy of a scientific field. International Studies in the Philosophy of Science, 30(3):311–314, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1316117.

Simons:2019:KIS

[Sim19] Massimiliano Simons. The Kuhnian image of science: Time for a decisive transformation? International Studies in the Philosophy of Science, 32(1): 78–80, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2019.
                                        1615668.

Sinaceur:2017:NN

[Sin17] Hourya Benis Sinaceur. Neurophilosophy of number. International Studies in the Philosophy of Science, 31(1): 1–25, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.
1370925.

Skow:2010:SAG

[Sko10] Bradford Skow. On a symmetry argument for the guidance equation in Bohmian mechanics. International Studies in the Philosophy of Science, 24(4):393–410, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.543350.

Shu-li:1991:ASD

[Sl91] Ji Shu-li. The antinomy of science and democracy in modern China. International Studies in the Philosophy of Science, 5(2):109–130, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573384.

Slater:2000:NNP

[Sla00] Ian Slater. Normative naturalism and Popperian views on reduction. International Studies in the Philosophy of Science, 14(3):325–326, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437836.

Slezak:1991:BBB

[Sle91] Peter Slezak. Bloor’s bluff: Behaviourism and the strong programme. International Studies in the Philosophy of Science, 5(3):241–256, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573397.

Sleutels:2009:BRP

[Sle09] Jan Sleutels. Beyond reduction: Philosophy of mind and post-reductionist philosophy of science. International Studies in the Philosophy of Science, 23(2):233–236, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903007279.

Slowik:2005:SOS

[Slo05] Edward Slowik. Spacetime, ontology, and structural realism. International Studies in the Philosophy of Science, 19(2):147–166, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500249456.

Sypel:1991:BR

[SM91] Roland Sypel and Radim Marada. Book reviews. International Studies in the Philosophy of Science, 5(2):187– 190, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        /doi/abs/10.1080/02698599108573389.

Sfendoni-Mentzou:1994:LNA

[SM94] Demetra Sfendoni-Mentzou. Laws of nature: Ante res or in rebus?. International Studies in the Philosophy of Science, 8(3):229–242, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573497. See re-

ply [Bro94].

Smith:1994:RMT

[Smi94] Edward E. Smith. Relating ‘a model theory’ to other research in induction. International Studies in the Philosophy of Science, 8(1):69–71, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573484.

Soe:2019:WI

[Søe19] Sille Obelitz Søe. What is information? International Studies in the Philosophy of Science, 32(1):73–75, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1615665.

Soler:2014:ARS

[Sol14] Léna Soler. Against robustness?

Strategies to support the reliability of scientific results. International Studies in the Philosophy of Science, 28(2):203–215, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.932530.

Spurrett:1999:FLC

[Spu99] David Spurrett. Fundamental laws and the completeness of physics. International Studies in the Philosophy of Science, 13(3):261–274, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573625.

Spurrett:2001:CLC

[Spu01] David Spurrett. Cartwright on laws and composition. International Studies in the Philosophy of Science, 15(3):253–268, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120080190.

Seselja:2014:CPV

[SS14] Dunja Seselja and Christian Straßer. Concerning Peter Vickers’s recent treatment of ‘paraconsistencitis’. International Studies in the Philosophy of Science, 28(3):325–340, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.953344.

Sevilla:2015:KPN

[SS15] Ana Sevilla and Elisa Sevilla. Knowledge production in non-European spaces of modernity: The Society of Jesus and the circulation of Darwinian ideas in Postcolonial Ecuador, 1860– 1890. International Studies in the Philosophy of Science, 29(3):233–250, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1179036.

Shaw:2022:FPP

[SS22] Jamie Shaw and Michael T. Stuart. Feyerabend and the philosophy of physics. International Studies in the Philosophy of Science, 35(1):1–4, 2022.

CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2193369.

Stuart:2024:FPP

[SS24] Michael T. Stuart and Jamie Shaw.

Feyerabend and the philosophy of physics, part II. International Studies in the Philosophy of Science, 37(4):155–159, 2024. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2024.2440672.

Sterrett:2006:MMM

[Ste06] Susan G. Sterrett. Models of machines and models of phenomena. International Studies in the Philosophy of Science, 20(1):69–80, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600641024.

Steel:2010:WIP

[Ste10] Daniel Steel. What if the principle of induction is normative? Formal learning theory and Hume’s problem. International Studies in the Philosophy of Science, 24(2):171–185, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.484544.

Steel:2011:CPR

[Ste11] Daniel Steel. On not changing the problem: a reply to Howson [MR2845407]. International Studies in the Philosophy of Science, 25(3):285– 291, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2011.

605250. See [How11, BTB16].

Schmidt:1996:BR

[STL^{+}96] Thomas Schmidt, Jens Timmermann, Brendan Larvor, Samir Okasha, Peter Goldie, Louis Caruana, Hasok Chang, Paul Castell, and Rom Harré. Book reviews. International Studies in the Philosophy of Science, 10(1):67–90, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573530.

Strand:2017:MRB

[Str17] Anders Strand. The multiple realization book. International Studies in the Philosophy of Science, 31(2):218– 221, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2018.
                                        1424767.

Suarez:2003:SRA

[Sua03] Mauricio Sua’rez. Scientific representation: against similarity and isomorphism. International Studies in the Philosophy of Science, 17(3):225–244, 2003. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859032000169442.

Suarez:2013:FCS

[Suá13] Mauricio Suárez. Fictions, conditionals, and stellar astrophysics. International Studies in the Philosophy of Science, 27(3):235–252, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.825499.

Svetlova:2014:MBA

[Sve14] Ekaterina Svetlova. Modelling beyond application: Epistemic and nonepistemic values in modern science. International Studies in the Philosophy of Science, 28(1):79–98, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915656.

Svorcova:2015:TTD

[Svo15] Jana Svorcová. Towards a theory of development. International Studies in the Philosophy of Science, 29(3):344– 347, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1179047.

Swan:2011:BML

[Swa11] Liz Stillwaggon Swan. The brain and the meaning of life. International Studies in the Philosophy of Science, 25(3):297–299, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.605252.

Szabo:2003:FSP

[Sza03] Laśzló E. Szabó. Formal systems as physical objects: a physicalist account of mathematical truth. International Studies in the Philosophy of Science, 17(2):117–125, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859031000160568.

Tamborini:2025:EBI

[Tam25] Marco Tamborini. The epistemology of bio-inspired disciplines: Unpacking modelling in biorobotics. International Studies in the Philosophy of Science, 38(1):21–37, 2025. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2024.2447979.

Teira:2013:PMC

[Tei13] David Teira. Philosophy of medicine: Causality, evidence and explanation. International Studies in the Philosophy of Science, 27(4):456–458, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.868186.

Tempczyk:1991:RDR

[Tem91] Michal Tempczyk. Random dynamics and the research programme of classical mechanics. International Studies in the Philosophy of Science, 5(3):227–239, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573396.

Tempczyk:1996:FGC

[Tem96] Michał Tempczyk. Fractal geometry — the case of a rapid career. International Studies in the Philosophy of Science, 10(1):53–65, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573529.

Tetens:1988:TCI

[Tet89] Holm Tetens. Teleology and the concept of information. International Studies in the Philosophy of Science, 3(2):230–237, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573338.

Thagard:2002:CCP

[Tha02] Paul Thagard. Curing cancer? Patrick Lee’s path to the reovirus treatment. International Studies in the Philosophy of Science, 16(1):79–93, 2002. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120118846.

Theodorou:2012:CEP

[The12] Panos Theodorou. Cognitive existentialism, phenomenology, and philosophy of science: Stimulating the dialogue. International Studies in the Philosophy of Science, 26(3):335–343, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
731734.

Theurer:2014:CBT

[The14] Kari L. Theurer. Complexity-based theories of emergence: Criticisms and constraints. International Studies in the Philosophy of Science, 28(3):277– 301, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
953342.

Thebault:2015:PTM

[Thé15] Karim Thébault. Physical theory: Method and interpretation. International Studies in the Philosophy of Science, 29(4):438–441, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1195148.

Thompson:1992:MBS

[Tho92] Paul Thompson. Mathematics in the biological sciences. International Studies in the Philosophy of Science, 6(3):241–248, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573434.

Thoren:2014:RUC

[Tho14] Henrik Thorén. Resilience as a unifying concept. International Studies in the Philosophy of Science, 28(3):303– 324, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2014.
                                        953343.

Thyssen:2022:WCE

[Thy22] Pieter Thyssen. What is a chemical element? A collection of essays by chemists, philosophers, historians, and educators. International Studies in the Philosophy of Science, 35(3–4):286– 289, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2023.
                                        2195306.

Tiles:1986:NRI

[Til87a] James Tiles. The normative role of images of science. International Studies in the Philosophy of Science, 1(2):196–208, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573291.

Tiles:1986:MMB

[Til87b] Mary Tiles. Mathesis and the masculine birth of time. International Studies in the Philosophy of Science, 1(1):16–35, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573280.

Tiles:1987:SDS

[Til88] Mary Tiles. Scientific dream space:

Symbolic forms and scientific theories. International Studies in the Philosophy of Science, 2(2):189–204, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573314.

Tiles:1990:ITH

[Til90a] J. E. Tiles. Information technology from Homer to DENDRAL. International Studies in the Philosophy of Science, 4(2):205–220, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573359.

Tiles:1990:HBT

[Til90b] Mary Tiles. Of heroes and butterflies:

Technological dreams and human realities. International Studies in the Philosophy of Science, 4(1):89–100, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573348.

Tiles:1998:ACD

[Til98] Mary Tiles. Adding a comparative dimension. International Studies in the Philosophy of Science, 12(2):109–110, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573586.

Tiles:1999:BAR

[Til99] Mary Tiles. Balancing acts: Rational agency and efficacious action. International Studies in the Philosophy of Science, 13(3):289–300, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573627.

Tishchenko:1988:SEB

[Tis89] V. J. Tishchenko. Social experiment and biological theory. International Studies in the Philosophy of Science, 3(2):238–247, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573339.

Talmont-Kaminski:2010:EE

[TK10] Konrad Talmont-Kaminski. Epistemology and emotions. International Studies in the Philosophy of Science, 24(2):229–233, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.484561.

Trpin:2022:I

[TO22] Borut Trpin and Barbara Osimani.

Introduction. International Studies in the Philosophy of Science, 35(3–4):209–210, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2218190.

Todd:2008:UTB

[Tod08] Cain S. Todd. Unmasking the truth beneath the beauty: Why the supposed aesthetic judgements made in science May not be aesthetic at all. International Studies in the Philosophy of Science, 22(1):61–79, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802280910.

Toon:2012:SSR

[Too12] Adam Toon. Similarity and scientific representation. International Studies in the Philosophy of Science, 26(3):241–257, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.731730.

Thoren:2023:PFM

[TP23] Henrik Thorén and Johannes Persson. Problem-feeding as a model for interdisciplinary research. International Studies in the Philosophy of Science, 36(1):39–59, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2152632.

Trautteur:1992:PSC

[Tra92] Giuseppe Trautteur. Problems with symbols. A commentary to Herbert Simon, “Scientific discovery as problem solving”. International Studies in the Philosophy of Science, 6(1):61–64, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573415. See

[Sim92a].

Trafimow:2020:TMP

[Tra20] David Trafimow. A taxonomy of major premises and implications for falsification and verification. International Studies in the Philosophy of Science, 33(4):211–229, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1964845.

Tresker:2021:TER

[Tre21] Steven Tresker. Treatment effectiveness and the Russo–Williamson thesis, EBM+, and Bradford Hill’s viewpoints. International Studies in the Philosophy of Science, 34(3):131–158, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2054396.

Trobok:2011:MR

[Tro11] Majda Trobok. Mathematics and reality. International Studies in the Philosophy of Science, 25(4):413–417, 2011.

CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.623370.

Tsou:2007:HLE

[Tso07] Jonathan Y. Tsou. Hacking on the looping effects of psychiatric classifications: What is an interactive and indifferent kind? International Studies in the Philosophy of Science, 21(3):329–344, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701589601.

Tuchanska:1990:CRR

[Tuc90] Barbara Tuchańska. Can relativism be reconciled with realism and causalism? International Studies in the Philosophy of Science, 4(3):285–294, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573368.

Tugby:2013:GTM

[Tug13] Matthew Tugby. Graph-theoretic models of dispositional structures. International Studies in the Philosophy of Science, 27(1):23–39, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.783979.

Tulkki:2016:OSN

[Tul16] Leena Tulkki. Objectivity in science: New perspectives from science and technology studies. International Studies in the Philosophy of Science, 30(3):314–317, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1316118.

Tympas:2021:AIA

[Tym21] Aristotle Tympas. Artifictional intelligence: Against humanity’s surrender to computers. International Studies in the Philosophy of Science, 34(1): 59–61, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2021.
                                        1980650.

Uebel:2001:CNE

[Ueb01] Thomas E. Uebel. Carnap and Neurath in exile: Can their disputes be resolved? International Studies in the Philosophy of Science, 15(2):211–220, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120059077.

Ugaglia:2015:AEE

[Uga15] Monica Ugaglia. Aristotle’s empiricism: Experience and mechanics in the 4th Century BC. International Studies in the Philosophy of Science, 29(1):99– 101, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1071553.

Vahid:1994:EBH

[Vah94] H. Vahid. Experience and belief: Haack on the problem of empirical basis. International Studies in the Philosophy of Science, 8(2):139–140, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573490.

Vaidman:1998:SEN

[Vai98] Lev Vaidman. On schizophrenic experiences of the neutron or why we should believe in the many-worlds interpretation of quantum theory. International Studies in the Philosophy of Science, 12(3):245–261, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573600.

Valentine:2012:HPP

[Val12] Elizabeth R. Valentine. History and philosophy of psychology. International Studies in the Philosophy of Science, 26(4):460–463, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.748501.

Vandevelde:1993:RR

[Van93] Toon Vandevelde. Rationality and reasonableness. International Studies in the Philosophy of Science, 7(1):85–86, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573454.

VanKerkhove:2010:PMP

[Van10] Bart Van Kerkhove. The philosophy of mathematical practice. International Studies in the Philosophy of Science, 24(1):118–122, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903467168.

Vaz:2016:PPP

[Vaz16] Sofia Guedes Vaz. Philosophy and the precautionary principle: Science, evidence, and environmental policy. International Studies in the Philosophy of Science, 30(1):95–98, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1195152.

vanBrakel:1987:PIM

[vBG88] J. van Brakel and J. P. M. Geurts. Pragmatic identity of meaning and metaphor. International Studies in the Philosophy of Science, 2(2):205–226, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573315.

vandenBerg:2014:KOE

[vdB14] Hein van den Berg. Kant’s organicism: Epigenesis and the development of critical philosophy. International Studies in the Philosophy of Science, 28(1):99–101, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.915657.

vanderSchaar:2014:BRB

[vdS14] Maria van der Schaar. Book review: The Bloomsbury Companion to Analytic Philosophy. International Studies in the Philosophy of Science, 28(4):437–439, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979672.

vanderTuin:2014:LWE

[vdT14] Iris van der Tuin. Life, war, Earth:

Deleuze and the sciences. International Studies in the Philosophy of Science, 28(2):226–229, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.932534.

Venturelli:2021:CCV

[Ven21] A. Nicolás Venturelli. Conceptual change in visual neuroscience: The receptive field concept. International Studies in the Philosophy of Science, 34(1):41–57, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.2002675.

Vetter:2011:LLS

[Vet11] Barbara Vetter. Laws and lawmakers: Science, metaphysics, and the laws of nature. International Studies in the Philosophy of Science, 25(1):83– 86, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
552420.

Vervoort:2015:MOD

[VG15] Louis Vervoort and Yves Gingras.

Macroscopic oil droplets mimicking quantum behaviour: How far can we push an analogy? International Studies in the Philosophy of Science, 29(3):271–294, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1179040.

Votsis:2007:BR

[VHO^{+}07] Ioannis Votsis, Katherine Hawley, Robert J. O’Hara, Lesley B. Cormack, and Diane Greco. Book reviews. International Studies in the Philosophy of Science, 21(1):103–117, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701306048.

Viale:1999:CCC

[Via99] Riccardo Viale. Causal cognition and causal realism. International Studies in the Philosophy of Science, 13(2):151–167, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573615.

Vicente:2006:CCP

[Vic06] Agustín Vicente. On the causal completeness of physics. International Studies in the Philosophy of Science, 20(2):149–171, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600814332.

Vicente:2013:WLE

[Vic13] Agustín Vicente. Where to look for emergent properties. International Studies in the Philosophy of Science, 27(2):137–156, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.813256.

Villena:2023:MMO

[Vil23] David Villena. Massive modularity:

an ontological hypothesis or an adaptationist discovery heuristic? International Studies in the Philosophy of Science, 36(4):317–334, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2277999.

Vistarini:2014:STS

[Vis14] Tiziana Vistarini. String theory and the scientific method. International Studies in the Philosophy of Science, 28(1):108–111, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915658.

Vitasovi:2016:ET

[Vit16] Daria Vitasovi. Experiencing time. International Studies in the Philosophy of Science, 30(4):409–412, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331982.

Vorontsov:1988:PSS

[Vor89] Nikolay N. Vorontsov. The problem of species and speciation. International Studies in the Philosophy of Science, 3(2):173–189, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573334.

vanStrien:2024:CQM

[vS24] Marij van Strien. The challenge of quantum mechanics to the rationality of science: Philosophers of science on Bohr. International Studies in the Philosophy of Science, 37(4):219–241, 2024. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2023.
2207404.

Wagner:2009:MVI

[Wag09] Roy Wagner. Mathematical variables as indigenous concepts. International Studies in the Philosophy of Science, 23(1):1–18, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590902843351.

Wagner:2013:HPI

[Wag13] Roy Wagner. A historically and philosophically informed approach to mathematical metaphors. International Studies in the Philosophy of Science, 27(2):109–135, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.813257.

Walker:2013:ESI

[Wal13] Rebecca L. Walker. The ethics of species: An introduction. International Studies in the Philosophy of Science, 27(2):225–228, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.813258.

Waskan:2008:KCI

[Was08] Jonathan Waskan. Knowledge of counterfactual interventions through cognitive models of mechanisms. International Studies in the Philosophy of Science, 22(3):259–275, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802567308.

Watkins:1992:BVK

[Wat92] John Watkins. Has BACON vindicated Kant? International Studies in the Philosophy of Science, 6(1):65–67, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573416.

Wayne:2015:CRE

[Way15] Andrew Wayne. Causal relations and explanatory strategies in physics. International Studies in the Philosophy of Science, 29(1):75–89, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1071552.

Walker:2006:AFT

[WĆ06] Mark A. Walker and Milan M.

Ćirković. Astrophysical fine tuning, naturalism, and the contemporary design argument. International Studies in the Philosophy of Science, 20(3):285–307, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600960945.

Weber:2009:HPC

[Web09] Erik Weber. How probabilistic causation can account for the use of mechanistic evidence. International Studies in the Philosophy of Science, 23(3):277–295, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903197757.

Weinert:1995:DQT

[Wei95] F. Weinert. The Duhem–Quine thesis revisited. International Studies in the Philosophy of Science, 9(2):147– 156, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        /doi/abs/10.1080/02698599508573514.

Weinstein:2013:STS

[Wei13] Steven Weinstein. Space, time, and stuff. International Studies in the Philosophy of Science, 27(1):98–101, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.783980.

Weinert:2016:TSG

[Wei16] Friedel Weinert. The time-symmetric Gold universe reconsidered. International Studies in the Philosophy of Science, 30(3):231–243, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1316119.

Wei:2024:NSM

[Wei24] Chuanke Wei. Natural selection, mechanism and phenomenon. International Studies in the Philosophy of Science, 37(1–2):37–50, 2024. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2024.2336269.

Whiteley:2015:BVS

[Whi15] Ella Whiteley. Beyond versus: The struggle to understand the interaction of nature and nurture. International Studies in the Philosophy of Science, 29(3):347–350, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1179048.

Whyte:1990:PPL

[Why90] J. T. Whyte. The psycho-physical laws of intentionality. International Studies in the Philosophy of Science, 4(3):295–304, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573369.

Whyte:2025:MSO

[Why25] Jennifer Whyte. Mathematical SETIbacks: Open texture in mathematics as a new challenge for messaging extra-terrestrial intelligence. International Studies in the Philosophy of Science, 38(1):59–77, 2025. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2024.2426194.

Wilkes:1991:MMC

[Wil91] K. V. Wilkes. Of mice and men: The comparative assumption in psychology. International Studies in the Philosophy of Science, 5(1):3–19, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573374.

Wilkes:1992:ESW

[Wil92] Kathy Wilkes. Editorial special war issue. International Studies in the Philosophy of Science, 6(3):171, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573426.

Wilson:1995:OMD

[Wil95] Fred Wilson. Once more to dissolve the ravens. International Studies in the Philosophy of Science, 9(2):135–146, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573513.

Williams:2011:FVD

[Wil11] James Williams. The force of the virtual: Deleuze, science, and philosophy. International Studies in the Philosophy of Science, 25(3):300–302, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.605253.

Williamson:2019:ECC

[Wil19] Jon Williamson. Establishing causal claims in medicine. International Studies in the Philosophy of Science, 32(1): 33–61, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2019.
                                        1630927.

Wilholt:2023:HRP

[Wil23] Torsten Wilholt. Harmful research and the paradox of credibility. International Studies in the Philosophy of Science, 36(3):193–209, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2273606.

Wong:1997:BR

[WMB97] James Wong, David Morris, and James Robert Brown. Book reviews. International Studies in the Philosophy of Science, 11(2):207–214, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573565.

Wolters:2015:GPC

[Wol15] Gereon Wolters. Globalized parochialism: Consequences of English as Lingua Franca in philosophy of science. International Studies in the Philosophy of Science, 29(2):189–200, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119420.

Wong:1997:MTP

[Won97] James Wong. The “making” of teenage pregnancy. International Studies in the Philosophy of Science, 11(3):273–288, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573571.

Woodward:2004:CCE

[Woo04] James Woodward. Counterfactuals and causal explanation. International Studies in the Philosophy of Science, 18(1):41–72, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590412331289251.

Woodcock:2007:BPN

[Woo07] Brian A. Woodcock. Bloch’s Paradox and the nonlocality of chance. International Studies in the Philosophy of Science, 21(2):137–156, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701498050.

Worrall:1990:RSS

[Wor90] John Worrall. Rationality, sociology and the symmetry thesis. International Studies in the Philosophy of Science, 4(3):305–319, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573370.

Webster:1993:BR

[WR93] Charles Webster and Don Robinson. Book reviews. International Studies in the Philosophy of Science, 7(3):267–272, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573470.

Wray:2008:AUG

[Wra08] K. Brad Wray. The argument from underconsideration as grounds for anti-realism: a defence. International Studies in the Philosophy of Science, 22(3):317–326, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802567399.

Wray:2015:PIF

[Wra15] K. Brad Wray. Pessimistic inductions: Four varieties. International Studies in the Philosophy of Science, 29(1):61–73, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1071551.

Wray:2021:IRR

[Wra21] K. Brad Wray. Induction, rationality, and the realism/anti-realism debate: a reply to Shech. International Studies in the Philosophy of Science, 34(4):243–247, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.

2113627. See [She21].

Wylie:1987:SAR

[Wyl88] Alison Wylie. ‘Simple’ analogy and the role of relevance assumptions: Implications of archaeological practice. International Studies in the Philosophy of Science, 2(2):134–150, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573311.

Wylie:2000:RON

[Wyl00] Alison Wylie. Rethinking objectivity: Nozick’s neglected third option.

International Studies in the Philosophy of Science, 14(1):5–9, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/026985900111864.

Yaghmaie:2017:HCP

[Yag17] Aboutorab Yaghmaie. How to characterise pure and applied science. International Studies in the Philosophy of Science, 31(2):133–149, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2018.1424763.

Yang:2008:MDP

[Yan08] Andrew S. Yang. Matters of demarcation: Philosophy, biology, and the evolving fraternity between disciplines. International Studies in the Philosophy of Science, 22(2):211–225, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590802496789.

Yan:2013:REV

[Yan13] Wong Muk Yan. Reliability and external validity of neurobiological experiments. International Studies in the Philosophy of Science, 27(4):429– 446, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2013.
                                        868188.

Ye:2010:NAE

[Ye10] Feng Ye. Naturalism and abstract entities. International Studies in the Philosophy of Science, 24(2):129–146, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
484539.

Yao:2024:SFA

[YH24] Siyu Yao and Amit Hagar. Searching for features with artificial neural networks in science: The problem of non-uniqueness. International Studies in the Philosophy of Science, 37(1–2): 51–67, 2024. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2024.
2346871.

Yaghmaie:2022:RSH

[YKMM22] Aboutorab Yaghmaie, Bijan Ahmadi Kakavandi, Saeed Masoumi, and Morteza Moniri. Representation and spacetime: The hole argument revisited. International Studies in the Philosophy of Science, 35(2):171–188, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2144457.

Yan:2023:SAI

[YTH23] Karen Yan, Meng-Li Tsai, and Tsung-Ren Huang. A scientometric approach to the integrated history and philosophy of science: Entrenched biomedical standardisation and citation-exemplar. International Studies in the Philosophy of Science, 36(2):143–165, 2023. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2243435.

Ye:2008:BR

[YTP08] Feng Ye, Emma Tobin, and Isabelle Peschard. Book reviews. International Studies in the Philosophy of Science, 22(2):227–236, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590802496805.

Zemplen:2013:HOG

[Zem13] Gábor Á. Zemplén. A history of optics from Greek antiquity to the nineteenth century. International Studies in the Philosophy of Science, 27(4):450–453, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868189.

Zenker:2016:CTS

[ZG16] Frank Zenker and Peter Gärdenfors. Continuity of theory structure: a conceptual spaces approach. International Studies in the Philosophy of Science, 30(4):343–360, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1331983.

Zhang:2015:BRL

[Zha15] Lu Zhang. Book review: Leibniz [Richard T. W. Arthur, Cambridge, Polity Press, 2014. xvi + 231 pp. ISBN 978-0-7456-5374-7, £55.00, EUR 68.80, US$69.95 (hardback); ISBN 978-0-7456-5375-4, £16.99, EUR 21.30, US$24.95 (paperback)]. International Studies in the Philosophy of Science, 29(2):215–218, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119423.

Zheng:1987:ISR

[Zhe88] Lan Zheng. Incommensurability and scientific rationality. International Studies in the Philosophy of Science, 2(2):227–236, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573316.

Zhu:2017:EWE

[Zhu17] Qin Zhu. Ethics within engineering: An introduction. International Studies in the Philosophy of Science, 31(3):327–330, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463698.

Zibakalam:1994:RDU

[Zib94] Saeid Zibakalam. Relativism due to underdetermination of theory by data. International Studies in the Philosophy of Science, 8(3):211–228, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573496.

Zinkernagel:2008:DTB

[Zin08] Henrik Zinkernagel. Did time have a beginning? International Studies in the Philosophy of Science, 22(3):237–258, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802567282.

Zwicky:2012:WI

[Zwi12] Jan Zwicky. What is ineffable? International Studies in the Philosophy of Science, 26(2):197–217, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.703480.